### v1

# Two point functions and quantum fields in the anti-de Sitter universe [^1] Dedicated to the memory of Henri Epstein

Abstract
We construct a manifestly covariant and coordinate-free plane-wave representation of scalar two-point functions in $d$-dimensional anti-de Sitter spacetime. The construction is based on a new class of holomorphic plane waves defined globally on the universal covering of the AdS via chiral cones in the complex null cone. Imposing AdS invariance, locality, positive definiteness and a spectral condition, we obtain integral representations as superpositions over relative homology cycles, reproducing the standard maximally analytic solutions in terms of Legendre functions of the second kind. In Poincaré coordinates, the two-point functions diagonalize into a Källén-Lehmann superposition of (d1)(d-1)-dimensional Minkowski correlators where the weight is a product of Bessel functions. This diagonalization clarifies the relation between Euclidean and Lorentzian AdS quantum field theory and allows Wick rotation of Euclidean Feynman diagrams to Lorentzian integrals supported on a single Poincaré patch while preserving full AdS covariance.
1 Introduction
The concept of a cyclic universe is deeply rooted in many cultures and long predates the notion of a linear, historical cosmos. In the Western tradition, Pythagoras, who is credited with introducing the term “cosmos”, was also the first to speak of a universe governed by recurring cycles. Porphyry, in his Life of Pythagoras, conveys the Pythagorean cosmological beliefs explicitly referring to the idea of cosmic cycles [1]: “He taught that […] after certain specified periods, the same events occur again.” In Eastern traditions, the idea of the universe as a cycle is often even more deeply ingrained. Hindu cosmology speaks of the Great Era, a vast cycle of creation, preservation, and destruction, while Buddhist thought highlights endless ages of arising and dissolution. Similarly, Chinese philosophy envisions the cosmos as continuously renewing itself through alternating cycles. Mircea Eliade, in his book Le mythe de l’éternel retour observed that virtually all ancient societies conceived of time as an endlessly recurring cycle, and rejected the notion of a singular, linear historical timeline [2]: “No event is unique, occurs once and for all… but it has occurred, occurs, and will occur perpetually; the same situations are reproduced that have already been produced in previous cycles and will be reproduced in subsequent cycles — ad infinitum.”
From a branch of philosophy and religion, cosmology evolved into a scientific discipline very recently, with the advent of Einstein’s general theory of relativity in 1915. In his initial cosmological model of 1917, Einstein assumed the universe to be static and unchanging, much like the skies of Ptolemy and Copernicus. To eliminate the notion of absolute space, he represented the universe as spherical, which led him to incorporate the cosmological constant—the key term to ensure staticity of a spherical dusty universe. Edwin Hubble’s 1929 observations later showed that the universe is actually expanding and consistent with Einstein’s original equations and therefore if there is no static universe away with the cosmological constant [3]. Then, in 1998, two teams, independently studying distant supernovae, found that the expansion of the universe is accelerating. This surprising discovery revived interest in the cosmological constant and suggested that it may constitute the dominant component of the universe’s energy content.
If the cosmological constant is truly nonzero, could it be negative? A negative cosmological constant entails fundamentally different physics: the vacuum energy density is negative and the gravitational effect attractive, in contrast to the repulsion caused by a positive cosmological constant. Recent research [4, 5] however indicates that a negative cosmological constant, when combined with other dark energy components, can be compatible with observational data, potentially allowing for a broader range of cosmic evolutions.
A universe devoid of matter and energy except for a negative cosmological constant has the anti-de Sitter (AdS) geometry. It features a constant negative curvature and exhibits closed timelike curves; while the light-cone ordering holds locally, it breaks down globally. In this respect, the AdS universe has a more evocative resemblance to the old cyclical cosmological worlds of Pythagoras rather than the modern relativistic expanding universes of Friedmann and Lemaître.
While the cyclical vision of time provided ancient peoples with a sense of continuity and meaning beyond the relentless progression of historical events, it is perceived as a scandal by physicists who are deeply committed to the causality principle and find such cyclicality at odds with the very idea of physical law: “many physicists are unhappy with the closed timelike curves and seek to assuage their feelings of guilt by claiming to pass to the universal covering spacetime. In this way they feel that they have exorcised the demon of “acausality”. However, therapeutic uttering these words may be, nothing is actually gained in this way.” (quoted from: G. Gibbons [6]).
The presence of closed timelike curves in AdS spacetime constitutes a genuine conceptual challenge in physical modeling. Calling it a scandal is perhaps an overstatement but, nevertheless, the word highlights a real and significant tension with causality. The challenge remains unresolved by passing to the universal covering spacetime: the geodesics originating from one point converge at the antipodal point after a half-period of “six great months” and again at the starting point after a full ”great era.” The universal cover removes the identification of points along the closed timelike curves but the periodic focusing of geodesics persists, providing no real gain in resolving the issues related to cyclic time in AdS (see Figure 2).
This conceptual challenge explains why even apparently simple questions—such as finding an explicitly AdS-covariant expression for the correlation functions of free quantum fields—have remained unresolved to date. In the standard approach to AdS quantum field theory, which mainly relies on imposing boundary conditions at spacelike infinity for modes decomposed in specific coordinate systems [7, 8, 9], the problem seems intractable. This method tends to obscure the global analyticity properties of both the AdS manifold and the correlation functions [11]. Neglecting these properties implies that the difficulties introduced by closed timelike curves could only be overcome with substantially greater effort.
This is one of the issues we tackle in the present work. We develop a new representation for the two-point functions of scalar quantum fields in a $d$-dimensional anti-de Sitter universe, based on a novel family of holomorphic plane waves that we introduce here and previously announced in [10]. These plane waves are almost always quasi-periodic functions on the AdS manifold and naturally reside on its covering space. The sole exception arises when the mass parameter takes integer values; in such cases, the plane waves become genuinely periodic.
Here is, deferring all detailed explanations, our formula for the two-point function of a massive scalar field:
$W^{(d)}_{\lambda}(z_{1},z_{2})=c_{d}(\lambda)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-d+1}d\mu_{\gamma}(\zeta).$ (1.1)
This integral representation is a superposition of plane waves; it does not separate space and time variables, nor does it depend on any particular coordinate system on the AdS manifold or any specific class of special functions. It is explicitly invariant under the anti-de Sitter group and clearly reveals the primitive analyticity domain of the two-point function [11]. Correlation functions for tensors and spinors can be straightforwardly derived by applying suitable differential operators to the right-hand side of this expression.
The significance of the new representation (1.1) is evident: it provides the closest possible analogue, within the AdS framework, to the plane-wave Fourier representation of the two-point function for a massive scalar field in (complex) Minkowski space $M_{d}^{(c)}$ with positive energy spectrum [12]:
${W}^{{M}_{d}}_{m}({{\rm z}}_{1},{{\rm z}}_{2})=\frac{1}{(2\pi)^{d}}\int e^{-i{\rm p}\cdot{\rm z}_{1}}e^{i{\rm p}\cdot{\rm z}_{2}}\theta({\rm p}^{0})\delta({\rm p}^{2}-m^{2})d{\rm p}.$ (1.2)
Manifestly covariant momentum space representations of quantum fields provide the physically intuitive ingredients of Feynman diagrammatics [13] and a great deal of the predictive power of quantum field theory depends on them. Furthermore, many crucial mathematical developments such as the PCT and the Spin-Statistics theorems [12] and the Euclidean formulation of quantum field theory [14] are based on the plane-wave Fourier representation of the correlation functions and the positivity of the energy spectrum. The importance of the Fourier representation (1.2) and of the corresponding representation of the Feynman propagator cannot be overestimated and we expect that (1.1) to occupy the same position in AdS QFT.
The second major topic addressed in this work pertains to the horospherical, or Poincaré, coordinate system, which covers exactly one half of the AdS manifold. We analyze the Fourier transform of the plane waves w.r.t. the Poincaré coordinates and show how this transform diagonalizes the two-point function (1.1). We present the final, remarkably elegant result here deferring detailed explanations to Sections 13 and 14:
$W^{(d)}_{\lambda}(z_{1}({\rm z}_{1},u),z_{2}({\rm z}_{2},u^{\prime})=$ (1.3)
$=\frac{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{2(2\pi)^{d-2}}\int e^{-i{\rm k}({\rm z}_{1}-{\rm z}_{2})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right)d{\rm k}$ (1.4)
The crucial role played by the global analyticity properties of the plane waves, which are inherited from our foundational axioms, will appear clearly in this calculation.
Moreover, this analysis yields an unexpected insight: Euclidean Feynman diagrams constructed on Lobachevski space can be mapped, via Wick rotation, onto diagrams entirely supported in the Poincaré patch. Multiple examples are given that substantiate the conjecture that any Feynman or Witten diagram computed by integration solely over a Poincaré patch of the real AdS manifold remains AdS covariant, even though the patch itself is not invariant.
The structure of the paper is as follows: in Section 2, we describe the geometric and group-theoretical foundations, detailing the real AdS spacetime as an embedded manifold, its complexification, and the universal covering of the real manifold.
In Section 3 we introduce the null cone in the complex ambient space and explore its connection to geodesic motion in the AdS universe following the approach developed in [15] for the geometrically simpler de Sitter case.
In Section 4 we describe the chiral tuboids originally introduced in [11]: these are complex analytic domains that play a pivotal role in the study of quantum fields on AdS spacetime. Within our framework, analyticity in these tuboids replaces the conventional approach, which relies on imposing boundary conditions at spacelike infinity on individual modes.
Section 5 investigates the analytic structure of two-point functions, which must meet the fundamental requirements of AdS covariance, locality, and positive definiteness—essential principles for any quantum field theory. However these general properties alone do not uniquely determine the solutions to the field equations of a given model; additional physical input is necessary to distinguish physically meaningful solutions from the vast set of mathematical possibilities. Here, as in Minkowski space, this role is played by a spectral condition, providing both a physical and a mathematical criterion for quantization. For two-point functions the spectral condition is equivalent to the maximal analyticity of the correlation functions in a certain complexification of the AdS manifold.
In Section 6, we construct plane wave solutions to the AdS-Klein-Gordon equation by exploiting the relationship between the pseudo-Laplacian in the ambient space and the AdS d’Alembertian. Plane waves are labeled by null vectors in the ambient space and by a mass parameter; in general they are well-defined only locally; the introduction of chiral cones in Section 7 provides a means to extend them into objects globally well-defined on the universal covering of the AdS manifold. Some mathematical aspects of our chiral cones are provided in Appendix A.
In Section 8 we return to the global analytic structure of the two-point functions for Klein-Gordon fields and recall their expression in terms of Legendre functions of the second kind whose argument is the scalar product of two complex AdS events, lifted to a complex domain of the Cartesian product of two copies of the covering manifold.
Section 9 and 10 focus on $AdS_{2}$ as a testing ground for the general case. We provide a detailed proof of the plane-wave expansion (1.1) and derive several novel integral representations of the two-point functions previously described in Section 8. The main challenge is the proper identification of the integration cycle $\gamma$ on the complex cone, a difficulty that reflects the intricate topological features of the AdS universe. Additionally, we address the bulk-to-boundary limit, which is crucial for holographic correspondences, and introduce the shadow transformation—a mathematical relation among dual waves of a given mass. These issues are resolved first in the two-dimensional context, which, while simpler, captures the critical aspects of the general case discussed in Section 11. Furthermore, we present in an example the value of our new formula (1.1) to derive nontrivial theorems concerning Legendre functions of the second kind.
In Section 12, we analyze the relevant integration cycles of the two-point function (1.1) within the particularly significant Poincaré coordinate system. In the subsequent Sections 13 and 14, we compute the horospherical-Fourier transform of the plane waves and show how this transform can be utilized to diagonalize the two-point function. The calculation fully exploits the holomorphic structure of the plane waves and the connection between the ambient AdS chiral tuboids and the Minkowskian tubes of the Poincaré leaves.
This approach allows for clarifying the issue of Wick rotations illustrating the relation between Lorentzian and Euclidean calculation in AdS. In the last Section 15 we derive a formula for the Feynman-Schwinger propagator corresponding to (1.4) and show how all banana Euclidean Feynman diagrams can be Wick-rotated into banana diagrams in the real manifold, where the integration is restricted solely to a Poincare patch. We also explore examples of diagrams with external legs where the integration is again restricted to a Poincaré patch and find that the result is AdS invariant.
The concluding remarks reflect on the broader significance of the results, their potential applications to perturbative quantum field theory in AdS, and the continuing evolution of both mathematical and physical understanding in this area.
2 AdS and its covering: geometrical setup
Let ${\bf R}^{d+1}_{2}$ denote [^2] The upper index in ${\bf R}^{d+1}_{2}$ indicates the dimension of the real vector space and the lower index indicates the number of positive directions. The same convention applies to ${\bf C}^{d+1}_{2}$ the vector space ${\bf R}^{d+1}$ equipped with the scalar product
$x\cdot y=x^{0}y^{0}+x^{d}y^{d}-\vec{x}\cdot\vec{y}=g_{\mu\nu}x^{\mu}x^{\nu}\ ,\ \ \ g_{\mu\nu}=diag(+1,-1,\ldots,-1,+1);$ (2.5)
the coordinates $x^{\mu}$ are relative to a chosen canonical basis $\{e_{0},e_{1},\ldots,e_{d}\}$ for ${\bf R}^{d+1}_{2}$:
$e_{\mu}\cdot e_{\nu}=g_{\mu\nu},\ \ x=x^{\mu}e_{\mu},\ \ \vec{x}=(x^{1},\ \dots,\ x^{d-1}).$ (2.6)
A vector $x$ is called timelike, spacelike or lightlike (aka null) according to whether $x\cdot x$ is positive, negative or null.
G=O(2,d1)G=O(2,d-1) is the orthogonal group of all real linear transformations of ${\bf R}^{d+1}_{2}$ which preserve the scalar product (2.5). $G$ has four components; the component of $G$ connected to the identity is denoted by $G_{0}=SO^{+}(2,d-1)$ and is interpreted as the relativity group of the AdS universe, which can be identified with the manifold
$AdS_{d}=\{x,\ x^{2}=x\cdot x=R^{2}\}.$ (2.7)
$G_{0}$ acts transitively on $AdS_{d}$; we chose
$e_{d}=(0,\ldots,0,R)$ (2.8)
to be the distinguished origin of $AdS_{d}$; in the folowing we will take for simplicity a unit radius $R=1$. The restriction to $AdS_{d}$ of the pseudo-Riemannian metric $g_{\mu\nu}dx^{\mu}dx^{\nu}$ is a Lorentzian metric with signature (+,,,)(+,\ -$\ldots,$-). The complex $d$-dimensional anti-de Sitter manifold $AdS_{d}^{(c)}$ can analogously be defined by its embedding in the complex space ${\bf C}^{d+1}_{2}$ endowed with the metric $g_{\mu\nu}$:
$AdS_{d}^{(c)}=\{z=x+iy\in{\bf C}^{d+1}_{2},\ z^{2}=1\}=\{z=x+iy,\ x^{2}-y^{2}=1,\ x\cdot y=0\}.$ (2.9)
The complexification $G_{0}^{(c)}$ of the real AdS group acts transitively $AdS_{d}^{(c)}$.
In spacetime dimension $d\geq 2$ the covering space of $AdS_{d}^{(c)}$ is $AdS_{d}^{(c)}$ itself. However, although the full space $AdS_{d}^{(c)}$ itself has a trivial covering, $G_{0}$ and $AdS_{d}$ admits nontrivial covering group $\widetilde{G}_{0}$ and covering space $\widetilde{}AdS_{d}$; also, the important tubular domains $\cal Z$$_{\pm}$ of $AdS_{d}^{(c)}$ defined below in Section 4 admit nontrivial coverings.
It is possible to single out in $AdS_{d}^{(c)}$ a Euclidean manifold $EAdS_{d}$; we choose the connected real submanifold defined by taking $z^{0}=ix^{0},\ x^{1},\ldots,x^{d}$ real, $x^{d}>0$:
$z^{\scriptscriptstyle{E}}=(ix^{0},\vec{x},x^{d}).$ (2.10)
It is the upper sheet of the two-sheeted hyperboloid with equation
$(x^{d})^{2}-(x^{0})^{2}-{\vec{x}}^{2}=1,\ \ \ x^{d}\geq 1$ (2.11)
equipped with the Riemannian metric induced by the scalar product (2.5). Global coordinates may be introduced on $AdS_{d}$ by the map
$(t,\,\vec{x})\mapsto\left(\sqrt{1+\vec{x}^{2}}\,\sin t,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,\cos t\right)$ (2.12)
where $t\in S_{1}={\bf R}/2\pi{\bf Z}$ and $\vec{x}\in{\bf R}^{d-1}$. Coordinates on $EAdS_{d}$ are obtained extending the above map to $S_{1}^{(c)}\times{\bf R}^{d-1}$ and taking $t=is$, namely
$(s,\,\vec{x})\mapsto(i\sqrt{1+\vec{x}^{2}}\,{\rm sh}\ s,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,{\rm ch}\ s).$ (2.13)
Lifting the coordinates on the covering ${\bf R}^{d}$ of $S^{1}\times{\bf R}^{d-1}$ provides a global coordinate system on $\widetilde{}AdS_{d}$. By further extending the coordinates to ${\bf C}\times{\bf R}^{d-1}$ provides a partial complexification of the covering manifold $\widetilde{}AdS_{d}$; this complexified manifold contains the same Euclidean space $EAdS_{d}$ as $AdS_{d}^{(c)}$.
3 The null cone and geodesical motion
Here we introduce the real null cone of the ambient space (minus its vertex at the origin)
${\cal C}_{d}=\{\xi\in{\bf R}^{d+1}_{2},\ \xi^{2}={(\xi^{0})^{2}}-{\vec{\xi}^{2}}+{(\xi^{d})^{2}}=0,\ \xi\not=0\}$ (3.14)
and its complexification
${\cal C}_{d}^{(c)}=\{\zeta=\xi+i\chi\in{\bf C}^{d+1}_{2},\ (\xi,\xi)-(\chi,\chi)=0,\ (\xi\cdot\chi)=0,\ \zeta\not=0\}.$ (3.15)
We single out the following basis manifold for the real cone ${\cal C}_{d}$:
$\gamma_{S}=\left\{(\xi^{0})^{2}+({\xi^{d}})^{2}=1=(\xi^{1})^{2}+\ldots+({\xi^{d-1}})^{2}\right\}.$ (3.16)
The manifold $\gamma_{S}$ is homeomorphic to $S_{1}\times S_{d-1}$; it may be parametrized by an angle $\alpha$ and a unit vector on a (d1)(d-1)-sphere as follows:
$\xi(\phi,\vec{n})=\left\{\begin{array}[]{lcl}\xi^{0}&=&\sin\phi\\ \xi^{i}&=&n^{i}\ \ \ \ \ ,\ \ \ \vec{n}^{2}=1\\ \xi^{d}&=&\cos\phi\end{array}\right$.; (3.17)
it should be also noted that both $\xi$ and $\xi$ belong to $\gamma_{S}$ and
$\xi(\phi+\pi,\vec{n})=-\xi(\phi,-\vec{n}).$ (3.18)
In the special two-dimensional $\gamma_{S}$ is the disjoint union of two disconnected components
$\gamma_{S(2)}^{\pm}=\left\{(\xi^{0})^{2}+({\xi^{d}})^{2}=1,\ \ \ \xi^{1}=\pm 1\right\}.$ (3.19)
Similarly ${\cal C}_{d}^{(c)}$ admits the basis manifold
$\gamma_{S}^{(c)}=\left\{(\zeta^{0})^{2}+({\zeta^{d}})^{2}=1=(\zeta^{1})^{2}+\ldots+({\zeta^{d-1}})^{2}\right\}$ (3.20)
which may be coordinatized by complexifying the parameters in Eq. (3.17).
3.1 Geodesics
The study of geodesics gives a clue for the construction and the interpretation of AdS plane waves suitable for quantum field theory. Since the AdS manifold is a real section of a complex sphere, geodesics of the AdS spacetime are conic sections of $AdS_{d}$ by 2-planes containing the origin of the ambient space ${\bf R}^{d+1}_{2}$ (see Figure 1; see also [15]).
For spacelike geodesics the 2-planes intersect the real null cone ${\cal C}_{d}$ along two generatrices.
Figure 1: Spacelike geodesics are parametrized by two real null vectors in the ambient space; they identify the asymptotic directions of the geodesics at spacelike infinity (the boundary).
Figure 2: Timelike geodesics are periodic curves. The planes containing these geodesics do not intersect the real null cone. Each timelike geodesic may be parametrized by a single complex null vector. All timelike geodesics originating from a given point reconverge at its antipodal point after half a period. This focusing phenomenon persists in the universal covering manifold as well.
Let $\xi$ and $\eta$ two null vectors identifying the generatrices, chosen in such a way that $(\xi,\eta)>0$ (changing $\eta$ with $\eta$ if necessary); a parametrization of the geodesic in terms of the proper length $s$ is then constructed as follows:
$x^{\mu}(s)=\frac{\xi^{\mu}e^{\alpha s}+\eta^{\mu}e^{-\alpha s}}{\sqrt{2\xi\cdot\eta}}$ (3.21)
where the constant $\alpha^{2}=-\dot{x}^{2}>0$. Real null vectors of the ambient space ${\bf R}^{d+1}_{2}$ may thus be interpreted as indicating asymptotic spacelike momentum directions on the AdS manifold.
On the other hand, planes containing timelike geodesics do not intersect the real cone ${\cal C}_{d}$ but they do intersect its complexification
${\cal C}_{d}^{(c)}=\{\zeta=\xi+i\chi\in{\bf C}^{d+1}_{2},\ (\xi,\xi)-(\chi,\chi)=0,\ (\xi\cdot\chi)=0,\ \zeta\not=0\}$ (3.22)
along two complex conjugated directions. Timelike geodesics on the real manifold $AdS_{d}$ can be thus parameterized by one complex null vector as follows:
$x^{\mu}(\tau)$ $=$ $\frac{\zeta^{\mu}e^{i\beta\tau}+{\zeta^{\mu}}^{*}e^{-i\beta\tau}}{\sqrt{2\zeta\cdot\zeta^{*}}},\ \ \ (\zeta\cdot\zeta)=0$ (3.23)
$v^{\mu}(\tau)$ $=$ $i\beta\frac{\zeta^{\mu}e^{i\beta\tau}-{\zeta^{\mu}}^{*}e^{-i\beta\tau}}{\sqrt{2\zeta\cdot\zeta^{*}}},\ \ \ (v\cdot v)=\beta^{2}.$ (3.24)
In terms of $\zeta$ we may construct a complete set of conserved quantities that characterise the geodesic motion:
$K^{\mu\nu}=\frac{m(x^{\mu}(\tau)v^{\nu}(\tau)-x^{\nu}(\tau)v^{\mu}(\tau))}{\sqrt{v\cdot v}}=\frac{i\,m(\zeta^{\mu}{\zeta^{\nu}}^{*}-\zeta^{\nu}{\zeta^{\mu}}^{*})}{\zeta\cdot\zeta^{*}}=\frac{i\,m(\zeta\wedge{\zeta}^{*})^{\mu\nu}}{\zeta\cdot\zeta^{*}}$ (3.25)
where $m$ is a mass parameter of the point particle in geodesical motion.
The parametrization (3.23) demonstrates that every temporal geodesic passing through a given event $x$ also passes through its antipodal event x-x. The proper time interval between $x$ and x-x along each of these geodesics is $\pi R$ (where we have set $R=1$ in the formulae), which corresponds to half of a ”Great Era,” [6].
On the other hand, there are events that are not connected by an arc of geodesics and therefore $AdS_{d}$ is not geodesically convex.
These characteristics also manifest in the universal covering space $\widetilde{}AdS_{d}$; In this covering manifold, geodesics continue to refocus at intervals of $\pi R$ i.e. half of the period. Thus, the intrinsic time periodicity; the time periodicity of the AdS spacetime does is not eliminated by passing to the cover, a point rightly emphasized by G. Gibbons [6].
4 The chiral tuboids $\cal Z$$_{+}$ and $\cal Z$$_{-}$
Let $\cal G$} be the Lie algebra of $G_{0}$ and let $M_{0d}$ the generator of rotations in the $(0,d)$-plane so that
$e^{tM_{0d}}=\left(\begin{array}[]{ccc}\cos t&\ldots&\sin t\\ \vdots&1&\vdots\\ -\sin t&\ldots&\cos t\end{array}\right).\$(4.26)
The following subsets of $\cal G$} plays a central role in our construction:
$C_{0}=\{M\in{\cal G}:\ M=g\,M_{0d}\,g^{-1}\ :\ g\in G_{0}\},$ (4.27)
$C_{+}=\bigcup_{\rho>0}\rho C_{0}.$ (4.28)
Let us now recall the definition [11] of the forward and backward chiral tuboids by the following definitions:
${\cal Z}_{\pm}=\{\exp(\tau M)\,x\ :\ M\in C_{0},\ \ x\in X_{d},\ \ \tau\in{\bf C},\ {\rm Im\,}\tau\gtrless 0\ \}.$ (4.29)
Their universal coverings $\widetilde{\cal Z}_{\pm}$ are given by the same formulae where $\exp(\tau M)$ is understood in the sense of the covering as an element of $\widetilde{G}_{0}^{(c)}$ and $x$ as an arbitrary element of $\widetilde{}AdS_{d}$. The chiral tuboids are manifestly invariant under the action of $G_{0}$. They can be equivalently described as follows [11]:
${\cal Z}_{\pm}=\{z=x+iy\in{\bf C}^{d+1}_{2}\ :\ y\cdot y>0,\ \epsilon(z)\gtrless 0\},$ (4.30)
where we denoted
$\epsilon(z)=y^{0}x^{d}-x^{0}y^{d}.$ (4.31)
The identity
$\epsilon(z)^{2}=(y\cdot y)(1+|\vec{x}|^{2}+|\vec{y}|^{2})+(y\cdot y)^{2}+|\vec{y}|^{2}+|\vec{x}|^{2}|\vec{y}|^{2}-(\vec{x}\cdot\vec{y})^{2}$ (4.32)
that is valid for any complex event $z\in AdS_{d}^{(c)}$, implies that $\epsilon(z)^{2}$ cannot vanish when $y\cdot y>0$; Eq. (4.30) then shows that $\cal Z$$_{+}$ and $\cal Z$$_{-}$ are disjoint open subsets of $AdS_{d}^{(c)}$ while Eq. (4.29) shows that they are connected. $\cal Z$$_{+}$ and $\cal Z$$_{-}$ may be generated simply by acting on the base point $e_{d}$ by a complex rotation followed by a Lorentz transformation:
${\cal Z}_{\pm}=\{g\,\exp(itM_{0d})\,e_{d}\ :\ g\in G_{0},\ \ \ t\gtrless 0\}.$ (4.33)
If $z$ belongs to either of the closed tuboids $\overline{{\cal Z}}_{\pm}$ then for any $M\in C_{0}$ and ${\rm Im\,}\tau\gtrless 0$ the point $\exp(\tau M)z$ belongs to either of their interiors $\cal Z$$_{\pm}$ .
The following geometrical properties are of central importance for AdS quantum field theory [11]:
(i) The Euclidean AdS manifold $EAdS_{d}$ is contained in ${\cal Z}_{+}\cup{\cal Z}_{-}\cup AdS_{d}$.
(ii) The image of either of the domains $\cal Z$$_{+}$ or $\cal Z$$_{-}$ under the coordinate map $z\mapsto z^{d}=e_{d}\cdot z$ is the cut-plane $\Delta={\bf C}\setminus[-1,1]$.
(iii) The image of the domain ${\cal Z}_{-}\times{\cal Z}_{+}$ (or ${\cal Z}_{+}\times{\cal Z}_{-}$) by the scalar product mapping $z_{1},z_{2}\mapsto z_{1}\cdot z_{2}$ is the cut-plane $\Delta$.
(iv) The map $z_{1},z_{2}\mapsto z_{1}\cdot z_{2}$ of ${\cal Z}_{-}\times{\cal Z}_{+}$ onto $\Delta$ can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$.
A more explicit characterization of the tuboids may be obtained by using the global complex coordinates of the complex AdS manifold that we write here as follows:
$z(t+is,\psi+i\phi,\vec{a}+i\vec{b})=\left\{\begin{array}[]{l}\mbox{ch\,}(\psi+i\phi)\sin(t+is),\cr\mbox{sh\,}(\psi+i\phi)\ (\vec{a}+i\vec{b}),\cr\mbox{ch\,}(\psi+i\phi)\cos(t+is),\end{array}\right.\ \ \ |\vec{a}|^{2}-|\vec{b}|^{2}=1,\ \ \vec{a}\cdot\vec{b}=0.$ (4.34)
The real covering manifold $\widetilde{}AdS_{d}$ is coordinatized by taking $\phi=0$, $s=0$, $\vec{b}=0$ and by unfolding the $t$ coordinate.
The chiral tuboids $\cal Z$$_{-}$ and $\cal Z$$_{+}$ are semi-tubes: they are domains in ${\bf C}^{d}$ invariant in one real direction:
${{\cal Z}_{+}}\ :\ \ \ \ \mbox{sh\,}s>\sqrt{\frac{|\vec{b}|^{2}\,\mbox{sh\,}^{2}(\psi)+\left(|\vec{b}|^{2}+1\right)\sin^{2}(\phi)}{\mbox{sh\,}^{2}(\psi)+\cos^{2}(\phi)}};$ (4.35)
${{\cal Z}_{-}}\ :\ \ \ \ \mbox{sh\,}s<-\sqrt{\frac{|\vec{b}|^{2}\,\mbox{sh\,}^{2}(\psi)+\left(|\vec{b}|^{2}+1\right)\sin^{2}(\phi)}{\mbox{sh\,}^{2}(\psi)+\cos^{2}(\phi)}}.$ (4.36)
bordered by the surfaces $s=s_{\pm}(\psi,\phi,|\vec{b}^{2}|)$. $\widetilde{\cal Z}_{\pm}$ are given by the same formulae where the real variable $t$ has been unfolded.
5 The analytic structure of two-point functions
The geometrical properties discussed in the previous section allow to fully characterize the two-point functions of scalar fields on either $\widetilde{}AdS_{d}$ or on the uncovered manifold $AdS_{d}$ for integer values of the mass parameters. We assume the following general properties [11]:
1. Covariance: $\cal W$} is invariant under the action of the group $\widetilde{G}_{0}$ (resp. $G_{0}$), i.e.
${\cal W}(gx_{1},gx_{2})={\cal W}(x_{1},x_{2})$ (5.37)
for all $g\in\widetilde{G}_{0}$ (resp. $G_{0}$).
2. Locality: if $x_{1}$ and $x_{2}$ are space-like separated then the commutator
${\cal C}({x_{1}},x_{2})={\cal W}({x_{1}},x_{2})-{\cal W}({x_{2}},x_{1})$ (5.38)
vanishes.
3. Positive Definiteness: For any smooth test function $f$ in a suitable test-function space ${\cal T}(\widetilde{}AdS_{d})$ the following positivity property holds:
${\cal W}(\overline{f}f)=\int{\cal W}(x_{1},x_{2})\overline{f(x)}f(y)d\sigma(x)d\sigma(y)\geq 0.$ (5.39)
[…]
${\cal W}(fg)=(\Omega,\ \varphi(f)\varphi(g)\Omega)\ \ for\ all\ \ f,g\in{\cal T}$ (5.40)
and that
$U(g)\,\varphi(f)\,U(g)^{-1}=\varphi(f_{\{g\}}),\ \ \ \ f_{\{g\}}(x)=f(g^{-1}x),\ \ \ \ U(g)\,\Omega=\Omega\$(5.41)
for all $g\in\widetilde{G}_{0}$ (resp. $G_{0}$).
[…]
4. (Strong) Spectral Condition: For every $M\in C_{+}$, every $\Psi\in{\cal H}$, and every $\cal C$$^{\infty}$ function $\widetilde{f}$ with compact support contained in (, 0)(-$\infty$,\ 0),
$\int_{{\bf R}}\left(\int\widetilde{f}(p)\,e^{-itp}\,dp\right)\,U(\exp tM)\,\Psi\,dt=0\ .$ (5.42)
Equivalently $\widehat{M}$ has its spectrum contained in $\bf R$$_{+}$.
The strong spectral condition actually broadens to cover general (interacting) QFT’s whose truncated $n$-point functions do not vanish [11]. When restricted to the case of two-point functions the condition tells us that the two-point function is not just any distribution on $\widetilde{}AdS_{d}\times\widetilde{}AdS_{d}$ but is realized as a boundary value of a certain analytic function: [11]:
$4^{\prime}$. Normal analyticity for two-point functions: $\mathcal{W}\left(x_{1},x_{2}\right)$ is the boundary value of a function $W\left(z_{1},z_{2}\right)$ holomorphic in the domain $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$
${\cal W}(x_{1},x_{2})=(\Omega,\ \varphi(x_{1})\varphi(x_{2})\,\Omega)=\lim_{\stackrel{{\scriptstyle\widetilde{\cal Z}_{-}\ni\ z_{1}\to x_{1}}}{{\scriptscriptstyle{\widetilde{\cal Z}_{+}\ni\ z_{2}\rightarrow x_{2}}}}}W(z_{1},z_{2}).$ (5.43)
AdS invariance then implies that to $W(z_{1},z_{2})$ there corresponds a function of a single complex variable that can be identified with the scalar product $z_{1}\cdot z_{2}$ when $z_{1}\in{\cal Z}_{-}$ and $z_{2}\in{\cal Z}_{+}$; the function $W(z_{1}\cdot z_{2})$ is called the reduced two-point function. Complex AdS invariance and normal analyticity then imply the following
$4^{\prime\prime}$. Maximal analyticity property: the reduced two-point function $W(z_{1}\cdot z_{2})$ extends to a function analytic in the covering $\widetilde{\Delta}$ of the cut-plane $\Delta=\{{\bf C}\setminus[-1,1]\}.$ For theories single-valued on $AdS_{d}$, $W(z_{1}\cdot z_{2})$ is analytic in $\Delta$.
Note that the two-point function of any field satisfying locality AdS invariance and the spectral condition enjoys maximal analyticity, as it happens in the Minkowski [12] and de Sitter cases [16, 17]. Importantly, this property holds even in the absence of the positive definiteness condition. However, if positive definiteness is satisfied, then the theory admits a quantum mechanical Hilbert space interpretation, allowing for a consistent quantum probabilistic interpretation.
6 Klein-Gordon equation, homogeneous functions and plane waves
Let us begin by recalling the relationship between the Laplace-Beltrami operator $\Box_{g}=g_{\mu\nu}\partial^{\mu}\partial^{\nu}$, defined with respect to the metric $g_{\mu\nu}$ of the ambient space ${\bf R}^{d+1}_{2}$, and the anti-de Sitter d’Alembert operator, which we will denote simply as $\Box$.
To this end let D{D} be the infinitesimal generator of dilations so that $({D}f)(x)=x^{\mu}\partial_{\mu}f(x)$ and let ${\cal M}_{\mu\nu}=(x_{\mu}\partial_{\nu}-x_{\nu}\partial_{\mu})$. The following relation holds on ${\bf R}^{d+1}_{2}$:
${\cal M}^{2}={\cal M}^{\mu\nu}{\cal M}_{\mu\nu}=2x^{2}\Box_{g}-2(d-1){D}-2{D}^{2}.$ (6.44)
If $f$ and $h$ are smooth functions on ${\bf R}^{d+1}_{2}$ and $h$ is $G_{0}$-invariant, then
${\cal M}_{\mu\nu}(h(x)f(x))=h(x){\cal M}_{\mu\nu}f(x)\ .$ (6.45)
This remains true if $h$ is a $G_{0}$-invariant distribution such as $\delta((x\cdot x)-1)$. In particular the operators ${\cal M}_{\mu\nu}$ and ${\cal M}^{2}$ commute with the restriction to invariant submanifolds of ${\bf R}^{d+1}_{2}$ and also with dilations. If $f$ is a smooth function or a distribution homogeneous of degree 0 on the open set $\Omega=\{x\in{\bf R}^{d+1}_{2}\ :\ (x,x)>0\}$ then
$\Box(f|_{AdS_{d}})=(\Box_{g}f)|_{AdS_{d}}=\frac{1}{2}{\cal M}^{2}f|_{AdS_{d}}$ (6.46)
because a homogeneous function of degree zero is dilation invariant. Since any smooth function or any distribution defined on $AdS_{d}$ has a unique homogeneous extension of degree zero in the open set $\Omega$, it follows that
$\Box\phi=\frac{1}{2}{\cal M}^{2}\phi$ (6.47)
for every smooth function or distribution $\phi$ defined on $AdS_{d}$.
Suppose now that $f$, defined in the ambient space ${\bf R}^{d+1}_{2}$ and homogeneous of degree $\lambda$, solves the massless equation $\Box_{g}f=0$. From Eq. (6.44) we get that
${\cal M}^{2}f(x)=-2\lambda(\lambda+d-1)f(x).$ (6.48)
Eqs. (6.47) and (6.48) then imply that the restriction of $f$ to the anti de Sitter manifold $AdS_{d}$ satisfies the Klein-Gordon equation
$(\Box+\lambda(\lambda+d-1))(f(x)|_{AdS_{d}})=0$ (6.49)
with complex squared mass
$m^{2}_{\lambda}=\lambda(\lambda+d-1).$ (6.50)
For real $\lambda$ and real $d$ the squared mass attains a minimum value which is negative:
$\lambda_{d}=-\frac{d-1}{2},\ \ \ \ m^{2}_{\lambda_{d}}=-\frac{(d-1)^{2}}{4}.$ (6.51)
It may be useful to introduce the complex parameter $\nu=\nu_{\lambda}$ by subtracting $\lambda_{d}$:
$\nu=\lambda+\frac{d-1}{2}$ (6.52)
so that $\nu=0$ corresponds precisely to the minimal value of the squared mass
$m_{\lambda}^{2}=\nu_{\lambda}^{2}-\frac{(d-1)^{2}}{4}.$ (6.53)
In the following we will use $\lambda$ and $\nu$ interchangeably.
Plane waves are the most important example of homogeneous solution of the massless equation in the ambient space:
$x,\zeta\mapsto(x\cdot\zeta)^{\lambda}$ (6.54)
where $\lambda$ is a complex number and $\zeta$ is a complex null vector of ${\cal C}_{d}^{(c)}$. Taking the restrictions to the anti de Sitter universe
$\phi_{\lambda}(x,\zeta)=(x\cdot\zeta)^{\lambda},\ \ \ x\in AdS_{d}$ (6.55)
satisfies the massive Klein-Gordon equation which we rewrite for clarity:
$(\Box+m_{\lambda}^{2})\phi_{\lambda}(x,\zeta)=0.$ (6.56)
The invariance of $m^{2}_{\lambda}$ under the involution
$\lambda\to 1-d-\lambda$ (6.57)
implies that, given the complex null vector $\zeta$, there is a second solution of Eq. (6.56) obtained by applying the involution to the exponent of the wave:
$\left(\Box+m_{\lambda}^{2}\right)\phi_{1-d-\lambda}(x,\zeta)=0.$ (6.58)
Now the main difficulty arises: although the complex parameter $\lambda$ in the above definition is unrestricted, the plane waves themselves are only locally well-defined on $AdS_{d}$. An exception occurs when $\lambda$ takes integer values; in that case the plane waves are globally well-defined. In the general case, however, it requires additional effort to construct single-valued plane waves on the covering space $\widetilde{}AdS_{d}$. This issue will be addressed in detail in the next section.
7 Chiral cones
The heuristic definition (6.54) is formally identical to the corresponding one employed in the de Sitter case [16, 17]; however, the analytic structure in the present context is significantly more intricate and requires a more elaborate analysis to be applicable for real non-integer and, more generally, for complex $\lambda$. With this in mind, we now introduce the chiral cones, defined as the following two $G_{0}$-invariant subsets of ${\cal C}_{d}^{(c)}$:
${\cal C}_{\pm}=\{\exp(\tau M)\,\xi\ :\ M\in C_{0},\ \ \xi\in{\cal C}_{d}\ \ \ \tau\in{\bf C},\ \ {\rm Im\,}\tau\gtrless 0\}.$ (7.59)
They may also be identified with the following subsets of ${\cal C}_{d}^{(c)}$
${\cal C}_{\pm}=\{\zeta=\xi+i\chi\in{\bf C}^{d+1}_{2}:\ \xi\cdot\xi=\chi\cdot\chi>0,\ \xi\cdot\chi=0,\ \epsilon(\zeta)\gtrless 0\}$ (7.60)
and may be generated simply by acting on the point $e_{d-1}+e_{d}$ as follows (see Appendix A):
${\cal C}_{\pm}=\{g\,\exp(itM_{0d})(e_{d-1}+e_{d})\ :\ g\in G_{0},\ \ \ t\gtrless 0\}.$ (7.61)
Here we summarize the properties that make use of plane waves on the AdS manifold and its covering possible:
  1. 1.
    Let
    $\overline{{\cal C}}_{+}=\{\zeta\in{\cal C}_{d}^{(c)}:\,\chi\cdot\chi\geq 0,\,\epsilon(\zeta)\geq 0,\,\zeta\not=0\},$ (7.62)
    $\overline{{\cal C}}_{-}=\{\zeta\in{\cal C}_{d}^{(c)}:\,\chi\cdot\chi\geq 0,\,\epsilon(\zeta)\leq 0,\,\zeta\not=0\}.$ (7.63)
    be the punctured closures of the tubes (including their real boundary points, minus the origin). It is obvious that
    $\exp(\tau M)(\overline{{\cal C}}_{\pm})\subset{\cal C}_{\pm},\ \ \ \makebox{for}\ \ M\in C_{0},\ \ {\rm Im\,}\tau\gtrless 0.$ (7.64)
  2. 2.
    The image of either $\cal C$$_{+}$ or $\cal C$$_{-}$ under the coordinate map $\zeta\mapsto\zeta^{d}=e_{d}\cdot\zeta$ is the punctured complex-plane ${\bf C}^{*}={\bf C}\setminus\{0\}$. This fact is intuitively clear as the chiral cones are the limit of the chiral tuboids when the radius $R$ tends to zero.
    An explicit check is also easy: using Eq. (A.273) for $\zeta\in{\cal C}_{+}$
    $e_{d}\cdot\zeta=x\cdot\exp(itM_{0d})(e_{d-1}+e_{d}),$ (7.65)
    where $x=g^{-1}e_{d}=\exp(sM_{0d})\,(0,\ \vec{x},\ \sqrt{1+\vec{x}^{2}})$, $s=\arctan(x^{0}/x^{d})$ and $t>0$ ; therefore
[…]
    Conversely, any $w\in{\bf C}^{*}$ can be written as $w=\lambda\cos(u+iv)$, with $v>0$ and $\lambda>0$, i.e. $w=e_{d}\cdot(\lambda\exp((u+iv)M_{0d})(e_{d-1}+e_{d}))$ is in the image of $\cal C$$_{+}$.
  3. 3.
    Let us consider now $z\in{\cal Z}_{-}$, $\zeta\in\overline{{\cal C}}_{+}$; there are $g\in G_{0}$, $M\in C_{0}$ and a complex $\tau$ with ${\rm Im\,}\tau>0$ such that
    $z\cdot\zeta=e_{d}\cdot\exp(\tau M)g^{-1}\zeta$ (7.67)
    Eqs. (7.64) and (7.66) then imply that the image of ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$ by the scalar product mapping $z,\zeta\mapsto z\cdot\zeta$ is also the punctured complex-plane $\bf C$$^{*}$ and the same is obviously true for ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$. In words: the scalar product $z\cdot\zeta$ cannot vanish when either $z\in{\cal Z}_{-}$ and $\zeta\in\overline{{\cal C}}_{+}$ or when $z\in{\cal Z}_{+}$ and $\zeta\in\overline{{\cal C}}_{-}$.
  4. 4.
    The map $z,\zeta\mapsto z\cdot\zeta$ of ${\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto $\bf C$$^{*}$ can be lifted to a map of $\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto the covering $\widetilde{\bf C}^{*}$ of the punctured plane $\bf C$$^{*}$. by abuse of language we denote the lifted scalar product by the same symbol.
  5. 5.
    Finally, let us consider $\chi\in{\cal C}_{-},\ \zeta\in\overline{{\cal C}}_{+}$, ; there are $g\in G_{0}$, $M\in C_{0}$ and a complex $\tau$ with ${\rm Im\,}\tau>0$ such that
    $\chi\cdot\zeta=(e_{d-1}+e_{d})\cdot\zeta^{\prime},\ \ \ \zeta^{\prime}=\exp(\tau M)g^{-1}\zeta\in{\cal C}_{+}$ (7.68)
    We are thus led to examine the scalar product
    $\xi\cdot\exp(itM_{0d})(e_{d-1}+e_{d})=i\ \mbox{sh\,}(t)\,\xi^{0}-\xi^{d-1}+\mbox{ch\,}(t)\xi^{d}\ \ \xi\in{\cal C}_{d},\ \ t>0$ (7.69)
    and this can never be zero.
    In conclusion the scalar product $\chi\cdot\zeta$ cannot vanish when either $\chi\in{\cal C}_{-}$ and $\zeta\in\overline{{\cal C}}_{+}$ or when $\chi\in\overline{{\cal C}}_{-}$ and $\zeta\in{\cal C}_{+}$ and.
7.1 Plane waves again
We are now in position to give a global definition of the anti-de Sitter plane waves. Given any complex number $\lambda$ the natural domains of definition for plane waves in the AdS universe are ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$ and ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$:
$\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}\ni\zeta,z\to\phi^{\pm}_{\lambda}(z,\zeta)=\left({z\cdot\zeta}\right)^{\lambda}=e^{\lambda\log\left({z\cdot\zeta}\right)}.$ (7.70)
A previous attempt [18] to introduce global waves was based on a local $i\epsilon$-prescription. Here, in their global domains of definition, the waves are univalued, holomorphic solutions of the Klein-Gordon equation with complex squared mass $m^{2}_{\lambda}=\lambda(\lambda+d-1).$
In the following we will also need to consider plane waves in subsets of ${\cal Z}_{+}\times\overline{{\cal C}}_{+}$ and ${\cal Z}_{-}\times\overline{{\cal C}}_{-}$ bordered; in that case it will be necessary not to trespass the hyperplanes where the argument of the waves vanishes. Hence, the admissible domains in these bordered subsets are restricted by these hyperplanes, which act as natural boundaries ensuring the holomorphicity and single-valuedness of the plane wave solutions. This condition secures the integrity of the analytic continuation of the waves and preserves the properties necessary for their role in the anti-de Sitter spacetime framework.
8 Klein-Gordon fields
Klein-Gordon fields display the simplest example of the analytic structure explained in Sect. 5. Their two-point functions are distributions on $\widetilde{}AdS_{d}\times\widetilde{}AdS_{d}$ satisfying the AdS Klein-Gordon equation w.r.t. both variables
$(\square+m^{2}_{\lambda}){\cal W}(x_{1},x_{2})=0,$ (8.71)
where the complex squared mass is parametrized by a complex parameter $\lambda$ as in Eq. (6.50). For any such $\lambda$ there are two invariant maximally analytic solutions of the above equation, satisfying all the properties specified in Section 5, apart from the positive definiteness which may hold or fail:
$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=$ $W^{(d)}_{\lambda}(z_{1}\cdot z_{2})={e^{-i\pi{d-2\over 2}}\over(2\pi)^{d\over 2}}((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}Q_{{d-2\over 2}+\lambda}^{d-2\over 2}(z_{1}\cdot z_{2})$ (8.72)
$=$ $\frac{(z_{1}\cdot z_{2})^{1-d-\lambda}\Gamma(d-1+\lambda)\,_{2}F_{1}\left(\frac{d-1+\lambda}{2},\frac{d+\lambda}{2};\frac{d+1+2\lambda}{2});{(z_{1}\cdot z_{2})^{-2}}\right)}{2^{d+\lambda}\pi^{\frac{d-1}{2}}\Gamma\left(\frac{d+1}{2}+\lambda\right)}$ (8.73)
$W^{(d)}_{1-d-\lambda}(z_{1},z_{2})$ $=$ $W^{(d)}_{1-d-\lambda}(z_{1}\cdot z_{2})={e^{-i\pi{d-2\over 2}}\over(2\pi)^{d\over 2}}((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}Q_{-{d\over 2}-\lambda}^{d-2\over 2}(z_{1}\cdot z_{2}).$ (8.76)
In the above formulae Q{Q} denotes the associated Legendre function of the second kind; it is a meromorphic function of the indices [19]; as function of the complex variable $z=z_{1}\cdot z_{2}$, it is analytic in the cut-plane
$\Delta_{1}=\{{\bf C}\setminus[-\infty,1]\}.$ (8.77)
As explained in Section 4, the scalar product can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$ and, thanks to the factor $((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}$, Eqs. (8.72) and (8.76) can and must be understood in this sense. In view of [19, Eq. (3.3.1.4)] one has also that
$W^{(d)}_{1-d-\lambda}(z)=W^{(d)}_{\lambda}(z)-\frac{\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)\Gamma(-\lambda)\Gamma(d+\lambda-1)}{(2\pi)^{d\over 2}}(z^{2}-1)^{-\frac{d-2}{4}}P_{\frac{d-2}{2}+\lambda}^{-\frac{d-2}{2}}(z).$ (8.78)
The second term at the rhs in this relation is regular on the cut $z\in[-1,1]$ and therefore does not contribute to the commutator; this means that the two solutions are equivalent at short distances. On the other hand, since it grows the faster the larger is $|\lambda+(d-2)/2|$ (see [19, Eqs. (3.9.2))], the two-solution have drastically different long range behaviors.
The question arises whether and when they are both acceptable or not [9]. The positive definiteness condition gives the answer. We will come back on this point later.
The Schwinger function (otherwise called the Euclidean propagator) is the restriction of the maximally analytic two-point function to the Euclidean Lobachevsky manifold extended to coincident points. Choosing the points in Eq. (8.72) as follows
$z_{1}=e_{d},\ \ z_{2}(u,\omega)=\left(i\omega^{0}\sqrt{u^{2}-1},\omega^{1}\sqrt{u^{2}-1},\ldots,\omega^{d-1}\sqrt{u^{2}-1},u\right),\ \ u>1,\ \ \omega^{2}=1$ (8.79)
so that $z_{1}\cdot z_{2}(u,\omega)=u>1$, we may write the Euclidean propagator as follows
$S^{(d)}_{\lambda}(u)=W^{(d)}_{\lambda}(z_{1},z_{2}(u,\omega))=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}(u^{2}-1)^{-\frac{d-2}{4}}Q^{\frac{d-2}{2}}_{\frac{d-2}{2}+\lambda}(u).$ (8.80)
9 Plane wave analysis of the two-point function: spacetime dimension $d=2$
With all ingredients in place, we begin by analyzing the plane-wave expansion of two-point functions within the two-dimensional spacetime $\widetilde{AdS_{2}}$. Though simpler than in the more general $d$-dimensional setting, this simpler case already illustrates many important features of plane-wave expansions while avoiding some of the geometrical complications encountered in higher dimensions.
The fundamental object is the two-point function
$\mathfrak{W}^{(2)}_{\lambda}(z_{1},z_{2})=c_{2}(\lambda)\int_{\gamma}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-1}d\mu_{\gamma}(\zeta),\ \ z_{1}\in{\widetilde{\cal Z}_{-}},z_{2}\in\widetilde{\cal Z}_{+}$. (9.81)
In this definition, the dependence on the integration cycle $\gamma$ is left implicit on the left-hand side. On the right-hand side, we have the integral of a closed differential form, where the cycle $\gamma$ is a curve on the complex cone $\mathbb{C}_{2}^{(c)}$ which generally depends on the points $z_{1}$ and $z_{2}$ that have been chosen. By construction, equation (9.81) defines a solution of the Klein–Gordon equation with respect to both variables and is analytic in $\mathcal{Z}_{-}\times\mathcal{Z}_{+}$ (provided the integral converges).
Our objective is to determine those choices of the integration cycle $\gamma$ that yield an AdS-invariant result; in such cases, the resulting expression must necessarily be a linear combination of (8.72) and (8.76). The selection of the integration cycle is crucial to enforce the AdS symmetry and obtain covariant quantizations of the Klein-Gordon field.
9.1 A closed cycle and the special case of half-integer mass parameter
Let us proceed by mimicking the de Sitter case [16, 17] and consider the integral (9.81) over a complete closed cycle $\gamma$ on the two-dimensional real cone ${\cal C}_{2}$ that does not depend on the two points $z_{1}$ and $z_{2}$; let us consider the integral over the upper component $\gamma^{+}_{S(2)}$ of the spherical basis of ${\cal C}_{2}$:
$\gamma_{S(2)}^{+}=\left\{\xi=(\sin\phi,1,\cos\phi),\ \ -\pi<\phi<\pi\right\},\ \ d\mu_{\gamma^{+}_{S(2)}}(\xi)=d\phi.$ (9.82)
Properties 3 and 4 of Sect. 7 guarantee that the integrand is always finite and therefore the integral converges. Unfortunately, in general, the integrand is not univalued on ${\cal C}_{2}$; after one revolution it gets multiplied by the global phase factor $\exp(-4\pi i\lambda)$. The conclusion is that Stokes theorem cannot be applied to deform the contour and the result is not AdS invariant.
Exceptions occur when $\lambda=l$ integer or $\lambda=l+\frac{1}{2}$ half-integer. When $\lambda=l$ the integrals over the two adjacent half-circles exactly cancel each other, causing the total integral to vanish identically. In contrast, when $\lambda=l+\frac{1}{2}$ the integrand has a period of $\pi$ and the contributions from the two halves combine coherently to yield, in a nontrivial manner, the invariant maximally analytic solution (8.72):
$\mathfrak{W}^{(2)}_{l+\frac{1}{2}}(z_{1},z_{2})=W^{(2)}_{l+\frac{1}{2}}(z_{1},z_{2})=\frac{1}{8\pi}\int_{\gamma_{S(2)}^{+}}(z_{1}\cdot\xi)^{l+\frac{1}{2}}(z_{2}\cdot\xi)^{-l-\frac{3}{2}}d\mu_{\gamma_{2+}}(\xi)=\frac{1}{2\pi}Q_{l+\frac{1}{2}}(z_{1}\cdot z_{2}).$ (9.83)
9.2 General case: relative homology
For general complex $\lambda$ integrating over a closed real cycle on the cone ${\cal C}_{2}$ such as $\gamma^{+}_{S(2)}$ gives a non-invariant result. Extending the integration to cycles on the covering space of the cone does not resolve this issue and is physically unsound, as it would lead to an overcounting of possible momentum directions. In fact, even a complete cycle of the type considered above already entails such an overcounting, which partly accounts for the vanishing of the integral (9.81) when $\lambda$ is an integer.
The solution for a general complex $\lambda$ is nevertheless given by Eq. (9.81), provided the integration cycle $\gamma$ is specified appropriately; here we generalize the results of [20] where integer values of $\lambda$ were studied: for ${\rm Re\,}\lambda>-1$ the cycle $\gamma$ should be open and have its endpoints on the plane orthogonal to $z_{1}$; more precisely, for ${\rm Re\,}\lambda>-1$ the cycle should belong to a relative homology class of $H_{1}({\cal C}_{-},\{\zeta:\,\zeta\cdot z_{1}=0\})$; for ${\rm Re\,}\lambda<0$ the cycle should belong to a relative homology class of $H_{1}({\cal C}_{+},\{\zeta:\,\zeta\cdot z_{2}=0\})$; the two possible choices actually give the same results.
Let us examine the first option. For ${\rm Re\,}\lambda\geq 0$ let us choose a relative cycle $\gamma(z_{1})$ in $H_{1}({\cal C}_{-},\{\zeta:\,\zeta\cdot z_{1}=0\})$ i.e. a path in $\cal C$$_{-}$ whose extremal points belong to the plane $\zeta\cdot z_{1}=0$:
$\mathfrak{W}^{(2)}_{\lambda}(z_{1},z_{2})=\frac{1}{4\pi}\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(\zeta\cdot z_{2})^{-\lambda-1}d\mu_{\gamma(z_{1})}(\zeta),\ \ z_{1}\in{{\cal Z}_{-}},z_{2}\in{\cal Z}_{+}.$ (9.84)
  • The value of the integral is independent of the chosen representative within the relative homology class. This follows from the fact that the integrand is a locally well-defined closed differential form, and that it vanishes on the relative boundary.
  • Convergence of the integral then allows to relax the condition to ${\rm Re\,}\lambda\geq-1$. The independence on the chosen representative in the relative homology class remains true.
  • AdS invariance follows by construction from the relation $\gamma(gz_{1})=g\gamma(z_{1})\subset{\cal C}_{-}$:
    $\mathfrak{W}^{(2)}_{\lambda}(gz_{1},gz_{2})=\frac{1}{4\pi}\int_{g\gamma(z_{1})}(z_{1}\cdot g^{-1}\zeta)^{\lambda}(g^{-1}\zeta\cdot z_{2})^{-\lambda-1}d\mu_{g\gamma(z_{1})}(\zeta)=$ (9.85)
    $=\frac{1}{4\pi}\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(\zeta\cdot z_{2})^{-\lambda-1}d\mu_{\gamma(z_{1})}(\zeta).$ (9.86)
$\mathfrak{W}_{\lambda}(z_{1},z_{2})$ is thus a function of one complex variable, namely of the scalar product $z_{1}\cdot z_{2}\in\Delta$; it is univalued only for integer $\lambda=l$. o evaluate this function, we choose the two points in the respective tuboids as follows:
$z_{1}=(\sin(-iv),0,\cos(-iv)),\ \ v\geq 0,$ $z_{2}=(\sin(iu),0,\cos(iu)),\ \ u>0.$ (9.87)
A representative of the relevant relative homology class is the following cycle in $\cal C$$_{-}$:
$\gamma(z_{1})=\left\{(\sin(\phi-iv),1,\cos(\phi-iv)),\ \ -\pi/2<\phi<\pi/2\right\}$ (9.88)
In the special case when $v=0$ the first point $z_{1}$ is on the real manifold and coincides with the base point $e_{2}$; the corresponding relative cycle is a subset of $\gamma^{+}_{S(2)}$ (see Figure (3))
$\gamma^{+}_{S(2)}(e_{2})=\left\{(\sin\phi,1,\cos\phi),\ \ -\pi/2<\phi<\pi/2\right\}=\left\{\xi\in\gamma^{+}_{S(2)}:e_{2}\cdot\xi=\cos\phi\geq 0\right\}.$ (9.89)
In all cases Eq. (9.84) becomes
$\mathfrak{W}^{(2)}_{\lambda}(z_{1},z_{2})=\frac{1}{4\pi}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{\lambda}(\cos(\phi-i(u+v)))^{-\lambda-1}d\phi$ (9.90)
and it is clear that we can set $v=0$ without loss of generality. The change of variables $\phi=\arctan(\mbox{sh\,}t)$ gives
$\mathfrak{W}^{(2)}_{\lambda}(z_{1},z_{2})=\frac{1}{4\pi}\int_{-\infty}^{\infty}{(\mbox{ch\,}u+i\ \mbox{sh\,}t\ \mbox{sh\,}u)^{-\lambda-1}}dt=\frac{1}{2\pi}Q_{\lambda}(\mbox{ch\,}u)$ (9.91)
i.e.
$\mathfrak{W}^{(2)}_{\lambda}(z_{1},z_{2})=W^{(2)}_{\lambda}(z_{1},z_{2})=\frac{1}{2\pi}Q_{\lambda}(z_{1}\cdot z_{2}).$ (9.92)
A similar formula holds true for the permuted two-point function $W_{\lambda}(x_{2},x_{1})$, but the primitive domain of analyticity is the opposite tuboid ${{\cal Z}_{+}}\times{{\cal Z}_{-}}$. Taking the difference of the boundary values gives the commutator:
$C^{(2)}_{\lambda}(x_{1},x_{2})=W^{(2)}_{\lambda}(x_{1},x_{2})-W^{(2)}_{\lambda}(x_{2},x_{1}).$ (9.93)
For a closer look into Eq. (9.84) we may change the integration variable in (9.90) by taking $z=ie^{i\phi}$:
$W^{(2)}_{\lambda}(\mbox{ch\,}u)=\frac{e^{(\lambda+1)u}}{2\pi}\int_{\gamma^{+}}\left(z^{2}-1\right)^{\lambda}\left(e^{2u}z^{2}-1\right)^{-\lambda-1}dz.$ (9.94)
The integration cycle (9.88) is mapped counterclockwise onto the upper unit semicircle in the complex $z$-plane: $\gamma^{+}=\{|z|=1,\ {\rm Im\,}z>0\}$. For $z=x+iy$ the null vector
$\zeta(z)=\left(-\frac{z^{2}+1}{2z},1,-\frac{i(z^{2}-1)}{2z}\right)$ (9.95)
is such that
$({\rm Im\,}\zeta)^{2}=\frac{\left(x^{2}+y^{2}-1\right)^{2}}{4\left(x^{2}+y^{2}\right)}\geq 0,\ \ \ \epsilon(\zeta)=\frac{\left(1-x^{2}-y^{2}\right)\left(1+x^{2}+y^{2}\right)}{4\left(x^{2}+y^{2}\right)}.$ (9.96)
This implies that $\zeta(z)$ belongs to $\cal C$$_{+}$ when $|z|<1$ and to $\cal C$$_{-}$ when $|z|>1$. This shows that $\zeta(z)$ lies in $\cal C$$_{+}$ when $|z|<1$, and in $\cal C$$_{-}$ when $|z|>1$. A deformation of the contour is allowed as long as the integration path remains within the closure of $\cal C$$_{-}$. Since the integral over the upper semicircle
$\gamma^{+}(r)=\{\,|z|=r,\,{\rm Im\,}z>0\,\}$
vanishes as $r\to\infty$, we can replace $\gamma^{+}$ in Eq. (9.94) by the two half-lines
$\{\,|x|>1,\,y=0\,\},$
and thus obtain
$W^{(2)}_{\lambda}(\mbox{ch\,}u)$ $=$ $\frac{e^{(\lambda+1)u}}{\pi}\int_{1}^{\infty}\left(x^{2}-1\right)^{\lambda}\left(e^{2u}x^{2}-1\right)^{-\lambda-1}dx=$ (9.97)
$=$ $\frac{\Gamma(\frac{1}{2})\Gamma(\lambda+1)\,}{2\pi\Gamma\left(\lambda+\frac{3}{2}\right)}e^{-(\lambda+1)u}{}_{2}F_{1}\left(\frac{1}{2},\lambda+1;\lambda+\frac{3}{2};e^{-2u}\right).$ (9.98)
Comparing this alternative representation of the two-point function with Eq. (9.92) we deduce a special case of a known (see [19, Eq. 45, page 136]) and useful identity which we reproduce here for the reader’s convenience:
$Q^{\mu}_{\nu}(\mbox{ch\,}u)=\frac{2^{\mu}e^{i\pi\mu}\sqrt{\pi}\Gamma(\mu+\nu+1)e^{-u(\mu+\nu+1)}(\mbox{sh\,}u)^{\mu}\,_{2}F_{1}\left(\mu+\frac{1}{2},\mu+\nu+1;\nu+\frac{3}{2};e^{-2u}\right)}{\Gamma\left(\nu+\frac{3}{2}\right)}.$ (9.99)
[…]
On the unit circle we have $\hat{z}=-ie^{-i\phi}$ so that the integration cycle is mapped counterclockwise onto the lower half $\gamma^{-}$ of the unit circle $|\hat{z}|=1$. By contour distortion taking place within $\cal C$$_{-}$ we may replace $\gamma^{-}$ with the compact interval [1,1][-1,1] thus obtaining
${\mathfrak{W}^{(2)}_{\lambda}}(\mbox{ch\,}u)=\frac{e^{-(\lambda+1)u}}{2\pi}\int_{-1}^{1}\left(1-x^{2}\right)^{\lambda}\left(1-e^{-2u}x^{2}\right)^{-\lambda-1}dx.$ (9.100)
For ${\rm Re\,}\lambda<0$ we may also consider relative cycles in $\cal C$$_{+}$ having their endpoints on the plane $\xi\cdot z_{2}=0$:
${\mathfrak{W}^{(2)}_{\lambda}}(z_{1},z_{2})=\frac{1}{4\pi}\int_{\gamma(z_{2})}(z_{1}\cdot\xi)^{\lambda}(\xi\cdot z_{2})^{-\lambda-1}d\mu_{\gamma(z_{2})}(\xi)$ (9.101)
A representative of the relevant homology class is the complex cycle
$\gamma(z_{2})=\left\{\xi=(\sin(\phi+iu),1,\cos(\phi+iu)),\ \ -\pi/2<\phi<\pi/2\right\}.$ (9.102)
The change of variable $z=ie^{i\phi}$ plus contour distortion give
${\mathfrak{W}^{\sharp(2)}_{\lambda}}(\mbox{ch\,}u)=\frac{e^{\lambda u}}{4\pi}\int_{-1}^{1}\left(1-e^{-2u}x^{2}\right)^{\lambda}\left(1-x^{2}\right)^{-\lambda-1}dx$ (9.103)
By comparing the latter expression with Eq. (9.100) we get
${\mathfrak{W}^{\sharp(2)}_{\lambda}}(z_{1},z_{2})=\frac{1}{2\pi}Q_{-\lambda-1}(z_{1},z_{2})$ (9.104)
which coincides with the previous one in the allowed range of values of $\lambda$.
9.3 Application to Legendre functions: a multiplication theorem for $Q_{\lambda}$
The key new feature of Eq. (9.84) is that it provides an integral representation of $Q_{\lambda}(z_{1}\cdot z_{2})$ in terms of two distinct points through their scalar product, opening up a wide range of potential applications. Equation (9.84) can be used to derive numerous new integral representations of Legendre functions with more complicated arguments (see, for instance, Eqs. (13.166) and (13.176)). Moreover, Eq. (9.84) can also be employed to generate new formulas for Legendre functions of the second kind.
To give an example let us choose $z_{1}=e_{d}$ and $z_{2}=(x\sin(iu),\sqrt{x^{2}-1},x\cos(iu))$; Eq. (9.84) gives
$Q_{\lambda}(x\,\mbox{ch\,}u)=\frac{1}{2}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{\lambda}\left(x\cos(\phi-iu)-\sqrt{x^{2}-1}\right)^{-\lambda-1}d\phi.$ (9.105)
By replacing the second factor in the integrand on the right-hand side with its binomial series, we are led to consider a series of integrals. These integrals can be evaluated using yet another new integral representation for associated Legendre functions of the second kind, which is derived in Eq. (11.129) below:
$\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{\lambda}(\cos(\phi-iu))^{-\lambda-1-n}d\phi=\frac{2^{1+\frac{n}{2}}\sqrt{\pi}\Gamma(1+\lambda)e^{-\frac{1}{2}in\pi}Q_{\frac{n}{2}+\lambda}^{\frac{n}{2}}(\mbox{ch\,}(u))}{\Gamma\left(\frac{1}{2}-\frac{n}{2}\right)\Gamma(1+n+\lambda)(\mbox{sh\,}u)^{\frac{n}{2}}}.$ (9.106)
What follows is a nontrivial — and possibly previously unknown — multiplication theorem for associated Legendre functions of the second kind:
$Q_{\lambda}(x\,\mbox{ch\,}u)=\frac{1}{x^{1+\lambda}}\sum_{n=0}^{\infty}\frac{1}{\Gamma(1+n)}\,\left(1-\frac{1}{x^{2}}\right)^{n}\frac{Q_{n+\lambda}^{n}(\mbox{ch\,}u)}{(2\,\mbox{sh\,}u)^{n}},\ \ x>1.$ (9.107)
10 Bulk-to-boundary limit and the shadow transformation
The plane-wave expansion (9.84) offers the possibility to study the boundary behaviour of the two-point function in a very simple way by sending $z_{1}$ to the boundary of $\widetilde{AdS_{2}}$ here identified with the covering of real cone ${\cal C}_{2}$. To this aim let us choose
$z_{1}(t,R)=\frac{\hat{z}_{1}(t,R)}{R}=\frac{1}{R}{\left(\sqrt{R^{2}+1}\sin t,1,\sqrt{R^{2}+1}\cos t\right)},\ \ z_{2}=(\sin(iu),0,\cos(iu))$ (10.108)
$\gamma(z_{1})=\{\xi=(\sin\phi,1,\cos\phi),\ \ t-t_{R}<\phi<t+t_{R}\},\ \ t_{R}=\arccos\left((R^{2}+1)^{-\frac{1}{2}}\right).$ (10.109)
Note that $z_{1}(t,R)\cdot z_{1}(t,R)=1$ while $\hat{z}_{1}(t,R)\cdot\hat{z}_{1}(t,R)=R^{2}$ and we are going to take the limit $R\to 0$. Suppose that $R>0$; when $R\to 0$ the integration is over the small interval [tR,t+R][t-R,t+R]; the plane-wave representation (9.84) of the two-point function immediately gives
$W^{(2)}_{\lambda}(z_{1}(t,R),z_{2})\simeq\frac{(\cos(t-iu))^{-\lambda-1}}{4\pi R^{\lambda}}\int_{-t_{R}}^{t_{R}}\left(\sqrt{R^{2}+1}\cos\phi-1\right)^{\lambda}d\phi$ (10.110)
$\simeq\frac{2^{-\lambda-2}\Gamma(\lambda+1)R^{\lambda+1}}{\sqrt{\pi}\Gamma\left(\lambda+\frac{3}{2}\right)}(\hat{z}(t,R)\cdot z_{2})^{-\lambda-1}.$ (10.111)
For $R<0$ the integration path tends instead to the whole circle $t-\pi,t+\pi]$ but the result does not change. For $\lambda>-1$ the rhs of (10.111) tends to zero but the following expression has a limit
$\lim_{R\to 0^{\pm}}R^{-\lambda-1}W^{(2)}_{\lambda}(z_{1}(t,R),z_{2})=\frac{2^{-\lambda-2}\Gamma(\lambda+1)}{\sqrt{\pi}\Gamma(\lambda+\frac{3}{2})}\left({\xi_{\pm}\cdot z}\right)^{-\lambda-1}$ (10.112)
where
$\xi_{\pm}=\pm(\sin t,1,\cos t)$ (10.113)
are the null vectors of ${\cal C}_{2}$ representing the relevant limit points of the boundary. The bulk-to-boundary two-point function is thus proportional to a plane wave.
10.1 The shadow transformation
If we now let the second point $z_{2}$ approach the boundary while keeping $z_{1}$ fixed, we obtain an integral formula relating the two waves conjugated by the involution (6.57). Such formulas are sometimes referred to in the literature as “shadow” transformations. Let us therefore choose
$z_{1}(v)=(\sin(-iv),0,\cos(-iv)),\ \ v>0,\ \ \ \ z_{2}(t,R)=\frac{\hat{z}_{2}(t+is,R)}{R},\ \ s>0$ (10.114)
and the cycle (9.88) in $\cal C$$_{-}$:
$W^{(2)}_{\lambda}(z_{1}(v),z_{2}(t,R))$ $=$ $\frac{{R}^{\lambda+1}}{4\pi}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{\lambda}\left({\sqrt{R^{2}+1}\cos(t+i(s+v)-\phi)-1}\right)^{-\lambda-1}d\phi$ (10.115)
$\simeq$ $\frac{R^{\lambda+1}\Gamma(\lambda+1)(\cos(t+i(s+v)))^{-\lambda-1}}{2^{\lambda+2}\sqrt{\pi}\Gamma\left(\lambda+\frac{3}{2}\right)}.$ (10.116)
By using AdS invariance and letting $R\to 0$ we deduce that for $z\in{\cal Z}_{-}$ and $\zeta\in{{\cal C}_{+}}$ there holds the following identity
$(z\cdot\zeta)^{-\lambda-1}=C_{2}(\lambda)\int_{\gamma(z)}(z\cdot\zeta^{\prime})^{\lambda}(\zeta^{\prime}\cdot\zeta)^{-\lambda-1}d\mu_{\gamma}(\zeta^{\prime})$ (10.117)
valid for ${\rm Re\,}\lambda>-1$, where $\gamma(z)$ is a cycle in $\cal C$$_{-}$ relative to the plane $\zeta^{\prime}\cdot z=0$ and
$C_{2}(\lambda)=\frac{2^{\lambda}\Gamma\left(\lambda+\frac{3}{2}\right)}{\Gamma\left(\frac{1}{2}\right)\Gamma(\lambda+1)}.$ (10.118)
This formula provides a concrete realization of the “shadow” transformation establishing an explicit integral relation between waves conjugated by the involution (6.57). A completely analogous formula holds for $z\in{\cal Z}_{+}$ and $\zeta\in{{\cal C}_{-}}$.
Putting everything together we finally arrive at the following more symmetric integral representation of the two point function: let $z_{1}\in{\cal Z}_{-}$, $z_{2}\in{\cal Z}_{+}$; $\gamma(z_{1})$ a cycle in $\cal C$$_{-}$ relative to the plane $z_{1}\cdot\zeta_{1}=0$; $\gamma(z_{2})$ a cycle in $\cal C$$_{+}$ relative to the plane $z_{2}\cdot\zeta_{2}=0$; then, for ${\rm Re\,}\lambda>-1$ there holds the following plane wave expansion of the two-point function:
$W^{(2)}_{\lambda}(z_{1},z_{2})=\frac{C_{2}(\lambda)}{4\pi}\int_{\gamma_{1}(z_{1})}\int_{\gamma_{2}(z_{2})}(z_{1}\cdot\zeta_{1})^{\lambda}(\zeta_{1}\cdot\zeta_{2})^{-\lambda-1}(\zeta_{2}\cdot z_{2})^{\lambda}d\mu_{\gamma(z_{1})}(\zeta_{1})d\mu_{\gamma(z_{2})}(\zeta_{2}).$ (10.119)
11 General spacetime dimension $d$
As is natural in the case of a Fourier-like plane-wave analysis, the results of the previous paragraph should be generalizable to arbitrary dimension $d$. We expect that the Fourier-like formula (9.84) should look the same in every dimension $d$; this means that, as anticipated in the Introduction, for $z_{1}\in\widetilde{\cal Z}_{-}$ and $z_{2}\in\widetilde{\cal Z}_{+}$ it should
$W^{(d)}_{\lambda}(z_{1},z_{2})=c_{d}(\lambda)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-d+1}d\mu_{\gamma}(\zeta)$ (11.120)
where $\gamma(z_{1})$ in $H_{d-1}({\cal C}_{-},\{\zeta:\,\zeta\cdot z_{1}=0\})$ is a cycle relative to the hyperplane $\zeta\cdot z_{1}=0$, the measure is the restriction to the chosen cycle of the invariant volume form on the cone and the constant $c_{d}(\lambda)$ is chosen by canonical normalization to reproduce Eq. (8.72).
Since the two-point function Eq. (11.120) is AdS invariant, we may compute it for the particular configuration
$z_{1}$ $=(\sin(-iv),\vec{0},\cos(-iv)),\quad v>0,\quad z_{1}\in\mathcal{Z}_{-},$ (11.121)
$z_{2}$ $=(\sin(iu),\vec{0},\cos(iu)),\quad\quad\ \,u>0,\quad z_{2}\in\mathcal{Z}_{+};$ (11.122)
the integration cycle in the relevant homology class is chosen as follows:
$\gamma(z_{1})=\left\{\zeta=(\sin(\phi-iv),\vec{n},\cos(\phi-iv)),\ \ -\frac{\pi}{2}<\phi<\frac{\pi}{2},\ \ \vec{n}\in S_{d-2}\right\}.$ (11.123)
In the special case where $v=0$, i.e. when $z_{1}=e_{d}$, the integration cycle can be taken real; in particular one can chose a relative cycle contained in $\gamma_{S}$ which, at variance with the two-dimensional cycle $\gamma_{S(2)}$, is a connected subset of ${\cal C}_{d}$:
$\gamma_{S}(e_{d})=\left\{\xi\in\gamma_{S}:e_{d}\cdot\xi=\cos\phi\geq 0\right\}.$ (11.124)
In the above configuration the integration over $S_{d-2}$ is trivial and we get
$W^{(d)}_{\lambda}(z_{1},z_{2})=$ $\frac{2\pi^{\frac{d-1}{2}}c_{d}(\lambda)}{\Gamma(\frac{d-1}{2})}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{\lambda}(\cos(\phi-iu))^{-\lambda-d+1}d\phi$ (11.125)
$=$ $-\frac{2^{d}i\pi^{\frac{d-1}{2}}c_{d}(\lambda)}{\Gamma(\frac{d-1}{2})}e^{u(\lambda+d-1)}\int_{\gamma^{+}}\left(-\frac{i\left(z^{2}-1\right)}{z}\right)^{\lambda}\left(-\frac{i\left(e^{2u}z^{2}-1\right)}{z}\right)^{-d-\lambda+1}\frac{dz}{z};$ (11.126)
in the second step we changed the integration variable to $z=ie^{i\phi}$; the integral is over the upper unit semicircle $\gamma^{+}$ in the complex $z$-plane counterclockwise. By contour deformation $\gamma^{+}$ may be replaced by the two half-lines $\{z=x+iy,\ |x|>1,y=0\}$. There is however a difference w.r.t. the two-dimensional case: the two integrals here differ by a phase factor; by adding their contributions we get
$\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{\lambda}(\cos(\phi-iu))^{-\lambda-d+1}d\phi=$ (11.127)
$=\frac{\pi\Gamma(\lambda+1)e^{-u(d+\lambda-1)}\,_{2}F_{1}\left(\frac{d-1}{2},d+\lambda-1;\frac{d+1}{2}+\lambda;e^{-2u}\right)}{2^{1-d}\Gamma\left(\frac{3}{2}-\frac{d}{2}\right)\Gamma\left(\frac{d+1}{2}+\lambda\right)}=$ (11.128)
$=\frac{2^{\frac{d}{2}}\sqrt{\pi}\Gamma(1+\lambda)}{\Gamma\left(\frac{3}{2}-\frac{d}{2}\right)\Gamma(d-1+\lambda)}\,(\mbox{sh\,}u)^{\frac{2-d}{2}}\,e^{-\frac{i\pi(d-2)}{2}}Q_{\frac{d-2}{2}+\lambda}^{\frac{d-2}{2}}(\mbox{ch\,}u)$ (11.129)
where, in the last step, we used Eq. (9.99). [^3] The above calculation provides another yet unknown (at least to me) integral representation of the Legendre function of the second kind that it is worth singling out: $e^{-i\pi\mu}Q^{\mu}_{\nu}(\mbox{ch\,}u)=\frac{2^{-\mu-1}\Gamma\left(\frac{1}{2}-\mu\right)\Gamma(\mu+\nu+1)}{\sqrt{\pi}\Gamma(-\mu+\nu+1)}(\mbox{sh\,}u)^{\mu}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}(\cos\phi)^{-\mu+\nu}(\cos(\phi-iu))^{-\mu-\nu-1}d\phi$ (11.130) By comparing Eqs. (8.72), (11.125) and (11.129) we finally get the value of the normalization constant
$c_{d}(\lambda)=-\frac{\pi^{1-d}\Gamma(\lambda+d-1)}{2^{d+1}\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)}.$ (11.131)
The above formula gives $c_{2}(\lambda)=1/{(8\pi)}$ whereas in Eq. (9.84) we had $1/({4\pi})$. Again, this is because in the two-dimensional case $\gamma_{2}$ has two disconnected components and we took the integral just on a cycle contained in one component $\gamma_{2+}$; the integral over the second component would produce an identical contribution. This peculiarity of the two-dimensional case disappears in the general case (see also Sect. 12.1 for another important choice of integration cycle).
When $d=2n+1$ is an odd integer the constant $c_{d}(\lambda)$ diverges but, at the same time, the integral vanishes; the corresponding two-point function may be obtained as a limit:
$W^{(2n+1)}_{\lambda}(z_{1},z_{2})=\frac{(-1)^{n+1}\Gamma(\lambda+2n)}{(2\pi)^{d}\Gamma(\lambda+1)}\int_{\gamma(z_{1})}(z_{1}\cdot\xi)^{\lambda}(\xi\cdot z_{2})^{-\lambda-2n}\log(\xi\cdot z_{2})d\mu_{\gamma}(\xi).\$(11.132)
As before we may use Eq. (11.120) to generate non trivial identities among Legendre functions. For example we may generalize the previous multiplication theorem (9.107) by choosing $z_{1}=e_{d}$ and
$z_{2}$ $=\left(x\sin(iu),\sqrt{x^{2}-1}\ \vec{\rm n},x\cos(iu)\right),\quad u>0,\ \ |\vec{n}|=1.$ (11.133)
Proceeding as for Eq. (9.107) we get
$Q_{\nu}^{\mu}(x\,\mbox{ch\,}u)\frac{(x^{2}\,\mbox{ch\,}^{2}u-1)^{\frac{\mu}{2}}}{x^{1+\mu+\nu}}\sum_{n=0}^{\infty}\frac{\left(\frac{1}{2}-\frac{1}{2\,x^{2}}\right)^{n}}{\Gamma(1+n)(\mbox{sh\,}u)^{\mu+n}}Q_{\nu+n}^{\mu+n}(\mbox{ch\,}(u))$ (11.134)
11.1 The shadow transformation
If in Eq. (11.120) we let $z_{2}$ move towards the boundary by keeping $z_{1}$ fixed, we discover the general ”shadow” transformation relating the two waves conjugated by the involution (6.57). To this aim we choose
$\hat{z}_{2}(t+iu,\vec{n}^{\prime},R)=\left(\sqrt{R^{2}+1}\sin(t+iu),\vec{n}^{\prime},\sqrt{R^{2}+1}\cos(t+iu)\right)\in AdS_{d}(R)$ (11.135)
$z_{2}(t+iu,R)=\frac{\hat{z}_{2}(t+iu,R)}{R},\ \ u>0$ (11.136)
and the cycle (11.123) in $\cal C$$_{-}$; we get
$W^{(d)}_{\lambda}(z_{1}(v),z_{2}(t+iu,R))=$ (11.137)
$=c_{d}(\lambda)R^{\lambda+d-1}\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}\int_{S^{d-2}}(\cos\phi)^{\lambda}(\cos(t+i(u+v)-\phi)-\vec{n}\cdot\vec{n}^{\prime})^{-\lambda-d+1}d\phi d\vec{n}$ (11.138)
$=\frac{R^{\lambda+d-1}(z_{1}\cdot\hat{z}_{2})^{1-d-\lambda}\Gamma(d-1+\lambda)\,}{2^{d+\lambda}\pi^{\frac{d-1}{2}}\Gamma\left(\frac{d+1}{2}+\lambda\right)}$ (11.139)
Letting $R\to 0$ and taking into account the AdS invariance of the above expression we deduce that
${(z\cdot\zeta^{\prime})^{1-d-\lambda}}=C_{d}(\lambda)\int_{\gamma(\zeta)}(z\cdot\zeta)^{\lambda}(\zeta\cdot\zeta^{\prime})^{-\lambda-d+1}d\mu_{\gamma}(\zeta)$ (11.140)
valid for $z\in\widetilde{\cal Z}_{-}$, $\zeta^{\prime}\in\overline{{\cal C}}_{+}$ and $\gamma(\zeta)\in H^{1}({\cal C}_{-},\{\zeta:\,z\cdot\zeta=0\})$. A similar formula holds for $z\in{\cal Z}_{+}$ and $\zeta^{\prime}\in{{\cal C}_{-}}$. In both cases the normalization constant takes the following value:
$C_{d}(\lambda)=-\frac{2^{\lambda-1}\Gamma\left(\frac{d+1}{2}+\lambda\right)}{\pi^{\frac{d-1}{2}}\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)}$ (11.141)
Note that again we have an extra 1/2 factor when we set $d=2$ w.r.t. Eq. (10.117) because here we are integrating on the two components over the disconnected sphere $S_{0}$.
Putting everything together we finally arrive at the following integral representation of the two point function: let $z_{1}\in{\cal Z}_{-}$, $z_{2}\in{\cal Z}_{+}$; $\gamma(z_{1})$ a cycle in $\cal C$$_{-}$ relative to the plane $z_{1}\cdot\zeta_{1}=0$; $\gamma(z_{2})$ a cycle in $\cal C$$_{+}$ relative to the plane $z_{2}\cdot\zeta_{2}=0$; then, for ${\rm Re\,}\lambda>-1$ there hold the following integral representation of the two-point function:
$W^{(d)}_{\lambda}(z_{1},z_{2})={c_{d}(\lambda)}{C_{d}(\lambda)}\int_{\gamma_{1}(z_{1})}\int_{\gamma_{2}(z_{2})}(z_{1}\cdot\zeta_{1})^{\lambda}(\zeta_{1}\cdot\zeta_{2})^{-\lambda-d+1}(\zeta_{2}\cdot z_{2})^{\lambda}d\mu_{\gamma(z_{1})}(\zeta_{1})d\mu_{\gamma(z_{2})}(\zeta_{2}).$ (11.142)
12 The Poincaré patch $AdS_{M}$
In this section and the ones that follow, we perform a decomposition of the plane waves (7.70) and diagonalize the two-point function (1.1) in the Poincaré coordinate system. This approach paves the way for novel possibilities to compute Feynman diagrams directly in the real AdS space, as opposed to working in its Euclidean counterpart as is usually done.
We denote by $AdS_{M}$ ($M$ stands for Minkowski) the region of the $d$-dimensional manifold $AdS_{d}$ (or, what is the same, of its covering $\widetilde{}AdS_{d}$) foliated by the family hyperplanes
$\Pi_{u}=\{x^{d-1}+x^{d}=1/u=e^{v}>0\}.$ (12.143)
The following system parametrizes the leaves by Lorentzian coordinates
[…]
where
${\rm x}^{2}={\rm x}^{2}_{M}={\rm x}\cdot{\rm x}=\eta_{\mu\nu}{\rm x}^{\mu}{\rm x}^{\nu},\ \ \ \ ds_{M}^{2}={(d{\rm x}^{0})}^{2}-{(d{\rm x}^{1})}^{2}-\cdots-{(d{\rm x}^{d-2})}^{2}$ (12.145)
are the Lorentzian scalar product and interval in a (d1)(d-1)-dimensional Minkowski space $M_{d-1}$.
In the Poincaré coordinates the AdS scalar product and interval are written as follows:
$(x-x^{\prime})_{AdS}^{2}=\frac{\left({\rm x}-{\rm x}^{\prime}\right)_{M}^{2}-({u^{2}-{u^{\prime}})^{2}}}{uu^{\prime}},\ \$(12.146)
${\mathrm{d}}s^{2}=\left.{\left(dx^{0}\right)}^{2}-{\left(dx^{1}\right)}^{2}-\ldots+{\left(dx^{d}\right)}^{2}\right|_{AdS}=\frac{1}{u^{2}}({{\mathrm{d}}s^{2}_{{M}}-{\mathrm{d}}u^{2}}).$ (12.147)
Eq. (12.146) implies that
$(x(u,{\rm x})-x(u,{\rm x}^{\prime}))_{AdS}^{2}=\frac{1}{u^{2}}({\rm x}-{\rm x}^{\prime})_{M}^{2}.$ (12.148)
This equation tells us that the Lorentzian geometry of the leaves is coherent with that of the AdS ambient space i.e. the causal ordering in any hyperplane $\Pi_{u}$ can be understood equivalently either in the Minkowskian sense of $\Pi_{u}$ or in the sense of the ambient AdS universe. The manifold $AdS_{M}$ is strongly causal; the price to pay is that it is geodesically incomplete.
The second crucial property is that the the complex leaves $\Pi_{u}^{(c)}$ contain copies of the Minkowsian tubes
$T_{\pm}=\{{\rm z}={\rm x}\pm i{\rm y}\in M_{d-1}^{(c)}:{\rm y}^{2}>0,\ {\rm y}^{0}\gtrless 0\}$ (12.149)
which are (respectively) contained in the ambient AdS tubes $\cal Z$$_{\pm}$.
This means the following: let us consider a partial complexification of $AdS_{M}$ where the Poincaré coordinates $\mathbf{x}^{\mu}$ in Eq. (12.144) are extended to complex values, while the coordinate $u$ remains real and positive:
$z({\rm z},u)=\left\{\begin{tabular}[]{lcl}$z^{\mu}$&=&$\frac{1}{u}{\rm z}^{\mu}$\cr$z^{d-1}$&=&$\frac{1-u^{2}}{2u}+\frac{1}{2u}{\rm z}^{2}$ \ ,\cr$z^{d}$&=&$\frac{1+u^{2}}{2u}-\frac{1}{2u}{\rm z}^{2}$\end{tabular}\right.\ \ {\rm z}^{2}=\eta_{\mu\nu}{\rm z}^{\mu}{\rm z}^{\nu},\ \ \ u>0.$ (12.150)
The leaves at fixed $u>0$ are copies of $M_{d-1}^{(c)}$. Suppose now that ${\rm z}={\rm x}+i{\rm y}$ and that ${\rm y}^{2}_{M}>0$, i.e. $\rm z$} belongs to either $T^{+}$ or $T^{-}$; it follows that
$(\mathrm{Im}\,z({\rm z},u))\cdot(\mathrm{Im}\,z({\rm z},u))=\frac{{\rm y}^{2}_{M}}{u^{2}}>0;$ (12.151)
Eq. (4.32) then implies that the Minkowskian tubes $T_{+}$ and $T_{-}$ of the leaves are contained into $\cal Z$$_{+}$ and, respectively $\cal Z$$_{-}$ (see [11] for more details). This property has also an important technical value; in particular it allows to compute the Fourier transforms of the waves w.r.t. the coordinates ${\rm x}^{\mu}$; we will do it in the next sections.
The last very important geometrical fact is that the above partial complexification already contains the whole Euclidean manifold $EAdS_{d}$.
In the following we will consider the submanifolds of the null cone defined by the condition
$\gamma_{P}^{\pm}=\left\{\xi\in{\cal C}_{d};\ \ \xi^{d-1}+\xi^{d}=\pm 1\right\}$ (12.152)
that we parametrize also by Lorentzian coordinates:
$\xi(\eta)=\left\{\begin{array}[]{lcc}\xi^{\mu}&=&{\eta^{\mu}}\\ \xi^{d-1}&=&\pm\frac{1}{2}(1+{\eta}^{2})\\ \xi^{d}&=&\pm\frac{1}{2}(1-{\eta}^{2})\end{array}\right.$. (12.153)
The invariant measure on $\gamma_{P}^{(\pm)}$ is $d\mu(\xi(\eta))=d\eta$ and the scalar product $x\cdot\xi$} take the form
$x(t,u)\cdot\xi(\eta)=\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u},\ \ \ \ \ \ \xi(\eta)\in\gamma_{P}^{(+)}$ (12.154)
$x(t,u)\cdot\xi(\chi)=-\frac{u}{2}+\frac{\left({{\rm x}}+{\chi}\right)^{2}}{2u},\ \ \ \ \xi(\chi)\in\gamma_{P}^{(-)}.$ (12.155)
12.1 Integration cycles
Let us specialize the first point in the two-point function (11.120) at the basis point $z_{1}=e_{d}=(0,\ldots,0,1)$. In spacetime dimension $d>2$ the cycle $\gamma_{S}(e_{d})$ (see Eq. (11.124)) is homeomorphic to a truncated cylinder with spherical section
$\gamma_{S}(e_{d})\cong[-\pi/2,\pi/2]\times S_{d-2}$ (12.156)
and its boundary is the disjoint union of two spheres
$\partial\gamma_{S}(e_{d})=\left\{(\xi^{0})=\pm 1,\ \ (\xi^{1})^{2}+\ldots+({\xi^{d-1}})^{2}=1\right\}.$ (12.157)
In Poincaré coordinates we would rather consider the following relative cycle contained in $\gamma_{P}^{+}$:
$\gamma_{P}^{+}(e_{d})=\left\{\xi\in\gamma^{+}_{P}:\ \ e_{d}\cdot\xi(\eta)=\frac{1}{2}(1-{\eta}^{2})\geq 0\right\}.$ (12.158)
The boundary of $\gamma_{P}^{+}(e_{d})$, namely its intersection with the plane $\xi^{d}=0$, is the disjoint union of two Lobatchevski spaces:
$\partial\gamma_{P}^{+}(e_{d})\rightarrow\xi(\eta)=\left\{\begin{array}[]{lcc}\xi^{0}&=&\pm\sqrt{1+\vec{\eta}^{2}}\\ \xi^{i}&=&{\eta^{i}}\\ \xi^{d-1}&=&1\\ \xi^{d}&=&0\end{array}\right.$. (12.159)
$\gamma_{P}^{+}$ is indeed mapped into $\gamma_{S}$ by the local rescaling
$\Theta(\xi(\eta)=\left\{\begin{array}[]{lcc}\xi^{\mu}&=&\frac{{2\eta^{\mu}}}{\sqrt{1+2{\eta^{0}}^{2}+2\vec{\eta}^{2}+(\eta^{2})^{2}}}\\ \xi^{d-1}&=&\frac{(1+{\eta}^{2})}{{\sqrt{1+2{\eta^{0}}^{2}+2\vec{\eta}^{2}+(\eta^{2})^{2}}}}\\ \xi^{d}&=&\frac{(1-{\eta}^{2})}{{\sqrt{1+2{\eta^{0}}^{2}+2\vec{\eta}^{2}+(\eta^{2})^{2}}}}\end{array}\right.$. (12.160)
The image of $\partial\gamma_{P}^{(+)}(e_{d})$ under the above map is the disjoint union of two open half-spheres:
$\Theta(\partial\gamma_{P}^{(+)}(e_{d}))=\left\{\pm 1,\frac{\vec{\eta}}{\sqrt{\vec{\eta}^{2}+1}},\frac{1}{\sqrt{\vec{\eta}^{2}+1}},0\right\};$ (12.161)
points of the equator are attained in the limit $\vec{\eta}\to\infty$ and correspond to points at infinity of the Lobachevsky space.
Points of $\Theta(\gamma_{P}^{+}(e_{d}))$ have positive (d1)(d-1)-coordinate; the image of the cycle $\gamma_{P}^{+}(e_{d})$ under the mapping $\Theta$ is therefore homeomorphic to a cylinder whose basis is a half-sphere; to build a cycle equivalent to $\gamma_{S}(e_{d})$ we must add to $\gamma_{P}^{+}(e_{d})$ a second cycle contained in $\gamma_{P}^{-}$, namely
$\gamma^{-}_{P}(e_{d})=\left\{\xi\in\gamma_{P}^{-}:\ \ e_{d}\cdot\xi(\chi)=-\frac{1}{2}(1-{\chi}^{2})\geq 0\right\}.$ (12.162)
It is immediately seen that
$\Theta[\partial\gamma^{-}_{P}(e_{d})]=\left\{\pm 1,\frac{\vec{\chi}}{\sqrt{\vec{\chi}^{2}+1}},-\frac{1}{\sqrt{\vec{\chi}_{2}^{2}+1}},0\right\}$ (12.163)
is the disjoint union of the two lacking halves of the spheres of the boundary of $\gamma_{S}(e_{d})$.
13 Waves and two-point functions in Poincaré coordinates: study of the two dimensional case
Figure 3: For a given real event $x$ all the spherical and parabolical integration cycles are shown; they have their endpoints on the plane $\xi\cdot x=0$ and belong to the same relative homology class. In the special case $d=2$ it is enough to integrate either one of the red or the blue t cycles which are disconnected.
The two-dimensional case is again simpler but represents a valuable training ground for solving the general case which is much harder.
In spacetime dimension $d=2$ the leaves $\Pi_{u}$ are time-like parabolae parametrized just by one time coordinate (i.e. $x^{0}=t/u$, see Eq. (12.144)). $\gamma^{+}_{P}$ is also parametrized by only one variable $\eta^{0}=\eta$ dual to the time $t$.
Let us chose the first point $z_{1}=x_{1}$ in Eq. (9.81) on the real AdS manifold (on the boundary of the backward tube); $x_{1}$ has real coordinates $(t,u)$; the second point $z_{2}$ is taken in the interior of the forward tube $\cal Z$$_{+}$ and is parametrized by the coordinates $(t^{\prime}+is^{\prime},u^{\prime})$ with $s^{\prime}$ positive. Our general formula requires that we integrate on a path homologous to the cycle
$\gamma_{P}^{+}(x_{1})=\left\{\xi\in\gamma^{+}_{P}:\ \ x_{1}(t,u)\cdot\xi(\eta)=\frac{u}{2}-\frac{\left({t}-{\eta}\right)^{2}}{2u}\geq 0\right\}.$ (13.164)
The result is
$W^{(2)}_{\lambda}(x_{1}(t,u),z_{2}(t^{\prime}+is^{\prime},u^{\prime}))=\frac{1}{2\pi}Q_{\lambda}\left(\frac{u^{2}+{u^{\prime}}^{2}-\left(t-t^{\prime}-is^{\prime}\right)^{2}}{2uu^{\prime}}\right)=$ (13.165)
$=\frac{1}{4\pi}\int_{t-u}^{t+u}\left(\frac{u^{2}-(t-\eta)^{2}}{2u}\right)^{\lambda}\left(\frac{{u^{\prime}}^{2}-(t^{\prime}+is^{\prime}-\eta)^{2}}{2u^{\prime}}\right)^{-\lambda-1}d\eta.$ (13.166)
This is yet another new integral representation of the Legendre function of the second kind $Q_{\lambda}$; this representation is naturally associated with the Poincaré coordinate system.
We now proceed to separate the variables in the above expression by taking the Fourier transform of the holomorphic plane waves (7.70) w.r.t. the time coordinate $t$:
$\int(x(t-i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\theta(\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{{\Gamma(-\lambda)}}\,e^{i\omega\eta}\,{\omega^{-\frac{1}{2}-\lambda}}\,J_{-\frac{1}{2}-\lambda}\left(\omega u\right),$ (13.167)
$\int(x(t+i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\theta(-\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{{\Gamma(-\lambda)}}\,e^{i\omega\eta}\,{|\omega|^{-\frac{1}{2}-\lambda}}\,J_{-\frac{1}{2}-\lambda}\left(|\omega|u\right),$ (13.168)
valid for for $1<\mathrm{Re}\,\lambda<0$; $\theta(\omega)$ is Heaviside’s step function.
The physically relevant feature of the above Fourier transforms is that planes waves holomorphic in the backward tube of the complex AdS manifold contain only positive frequencies w.r.t. the time coordinate of the Poincaré branes and planes waves holomorphic in the forward tube only contain negative frequencies. This support property is a consequence of the geometric inclusion of the tubes described in Eq. (12.151).
Inversion allows to rewrite the integral (13.166) as follows:
$\int\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}\left(\frac{u^{\prime}}{2}-\frac{\left({t^{\prime}+is^{\prime}}-{\eta}\right)^{2}}{2u}\right)^{-\lambda-1}d\eta=$ (13.169)
$=\frac{(2\pi u^{\prime})^{1/2}}{\Gamma(\lambda+1)}\int\int_{-t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{-i\omega(t^{\prime}+is^{\prime}-\eta)}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}(|\omega|u^{\prime})\theta(-\omega)}d\omega d\eta$ (13.170)
The integral over $\eta$ contains both positive and negative frequencies (no step function is present here); the result is
$\int_{t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{i\omega\eta}d\eta=\sqrt{2\pi u}|\omega|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)J_{\frac{1}{2}+\lambda}(|\omega|u),\ \ \ \mathrm{Re}\,\lambda>-1.$ (13.172)
Putting everything together we arrive at the final result:
$W^{(2)}_{\lambda}(x_{1}(t,u),z_{2}(t^{\prime}+is^{\prime},u^{\prime}))=\frac{1}{2}{(uu^{\prime})^{1/2}}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}J_{\frac{1}{2}+\lambda}(\omega u)J_{\frac{1}{2}+\lambda}(\omega u^{\prime})d\omega.$ (13.173)
This expression has been proven under the condition $1<\mathrm{Re}\,\lambda<0$ but is valid more generally in the sense of distributions. It provides a diagonalization of the two-point function which, for real $\lambda$, manifestly exhibits its positive-definiteness where it converges.
By comparing this expression with the one obtained by inserting the Fourier representation of both the first and the second factor in Eq. (13.166) we get as a bonus a time-honored formula of Sonine which we do not reproduce here [21].
13.1 The other cycle
In preparation for the general case, it is instructive to integrate also over the other non-compact parabolic cycle. (see Figure 3). Even though we expect a result identical to (13.174) the way to reach it is rather different as the integration cycle is now unbounded and touches infinity. The delicate point is the behavior of the waves at infinity and the way to circumvent such a difficulty will give crucial indications on how to deal with the problem in general spacetime dimension $d$.
The relevant cycle is
$\gamma_{P}^{-}(x_{1})=\left\{\xi\in\gamma_{P}^{-}:\ \ x_{1}\cdot\xi(\chi)=-\frac{u}{2}+\frac{\left({{\rm x}}+{\chi}\right)^{2}}{2u}\geq 0\right\}.$ (13.175)
This choice leads to an alternative representation of the two-point function:
$W^{(2)}_{\lambda}(x_{1},z_{2})=\frac{1}{4\pi}\int_{-\infty}^{-t-u}+\int_{-t+u}^{\infty}\left(\frac{(t+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{(t^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)^{-\lambda-1}d\chi$ (13.176)
where it is understood that
$\lim_{\chi\to\pm\infty}\text{Arg}\left(\frac{(t^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)=0.$ (13.177)
This condition has to be interpreted as follows: in formula (13.176) there are two disconnected semi-infinite intervals of integration; the phase of the second factor of the integrand has to be chosen with two different determinations. Eq. (13.176) actually means
$W^{(2)}_{\lambda}(z_{1},z_{2})=\frac{1}{4\pi}\int_{-\infty}^{-t-u}\left(\frac{(t+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{(t^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{-}^{-\lambda-1}d\chi$ (13.178)
$+\frac{1}{4\pi}\int_{-t+u}^{\infty}\left(\frac{(t+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{(t^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{+}^{-\lambda-1}d\chi$ (13.179)
with the two distinct determinations ($\pm$) are specified at plus and minus infinity by Eq. (13.177). The Fourier transforms of the waves according to the above determinations are given by
$\int\left(\frac{(t^{\prime}+is^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{\pm}^{-\lambda-1}e^{i\omega t^{\prime}}dt^{\prime}=$ (13.180)
$=\pm i\theta(-\omega)e^{-i\omega(\chi+is^{\prime})}|\omega|^{\frac{1}{2}+\lambda}\left(e^{\mp 2i\pi\lambda}-1\right)\sqrt{2\pi u^{\prime}}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)\Gamma(-\lambda).$ (13.181)
Inversion gives the Fourier representations of the plane waves:
$\left(\frac{(t^{\prime}+is^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{\pm}^{-\lambda-1}=$ (13.182)
$=\pm\frac{i\Gamma(-\lambda)\sqrt{u^{\prime}}\left(e^{\mp 2i\pi\lambda}-1\right)}{\sqrt{2\pi}}\int_{-\infty}^{0}e^{-i\omega(t^{\prime}+is^{\prime}+\chi)}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)}d\omega.$ (13.183)
Inserting these formulae into (13.179) we get
$W^{(2)}_{\lambda}(z_{1},z_{2})=$ (13.184)
$=$ $\frac{\Gamma(-\lambda)\sin(\pi\lambda)\sqrt{u^{\prime}}}{4\pi\sqrt{2\pi}}\int_{-\infty}^{0}e^{i\omega(t-t^{\prime}-is^{\prime})}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)}{}d\omega\int_{u}^{\infty}\cos(\omega\chi-\pi\lambda)\left(\frac{\eta^{2}-u^{2}}{2u}\right)^{\lambda}d\chi$ (13.186)
$=$ $\frac{1}{2}{(uu^{\prime})^{1/2}}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}J_{\frac{1}{2}+\lambda}(\omega u)J_{\frac{1}{2}+\lambda}(\omega u^{\prime})d\omega.$ (13.187)
As anticipated, we recover Eq. (13.174), though by means of a significantly more involved procedure. In the case of general dimensions, the two separate integration cycles contribute differently. Therefore, it is essential to account for both contributions to obtain the complete result.
14 Waves and two-point functions in Poincaré coordinates: general case
Let us now face the problem of diagonalizing the two-point function (1.1) in Poincaré coordinates in arbitrary spacetime dimension $d$. To this end let us choose $x_{1}=x_{1}({\rm x},u)$ real and $z_{2}=z_{2}({\rm z}^{\prime},u^{\prime})\in{\cal Z}_{+}$ with ${\rm z}^{\prime}\in T_{+}$ in the relevant complex Minkowskian brane.
The integration cycle consists of two disconnected parts, denoted as $\gamma_{P}(x_{1})={\gamma^{+}_{P}(x_{1})\,\cup\,\gamma^{-}_{P}(x_{1})}$ where
$\gamma^{+}_{P}(x_{1})=\left\{\xi(\eta)\in\gamma_{P}^{+}:\left(\eta^{0}-{\rm x}^{0}\right)^{2}\leq\left(\vec{\eta}-\vec{\rm x}\right)^{2}+u^{2}\right\},$ (14.188)
$\gamma^{-}_{P}(x_{1})=\left\{\xi(\chi)\in\gamma_{P}^{-}:\left(\chi^{0}+{\rm x}^{0}\right)^{2}\geq\left(\vec{\chi}+\vec{\rm x}\right)^{2}+u^{2}\right\}.$ (14.189)
As emphasized previously, unlike the two-dimensional case, it is essential to integrate over both disconnected cycles. The following expression is the one that requires evaluation.
$W^{(d)}_{\lambda}(x_{1}({\rm x},u),z_{2}({\rm z}^{\prime},u^{\prime}))=$ (14.190)
$=c_{d}(\lambda)\int d\vec{\eta}\int_{x^{0}-\sqrt{\left(\vec{\eta}-\vec{\rm x}\right)^{2}+u^{2}}}^{x^{0}+\sqrt{\left(\vec{\eta}-\vec{\rm x}\right)^{2}+u^{2}}}\left(\frac{u^{2}-({\rm x}-\eta)^{2}}{2u}\right)^{\lambda}\left(\frac{{u^{\prime}}^{2}-({\rm z}^{\prime}-\eta)^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}d\eta^{0}$ (14.192)
$+c_{d}(\lambda)\int d\vec{\chi}\int_{-\infty}^{-{\rm x}^{0}-\sqrt{\left(\vec{\chi}+\vec{\rm x}\right)^{2}+u^{2}}}\left(\frac{({\rm x}+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{({\rm z}^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}d\chi^{0}+$ (14.193)
$+c_{d}(\lambda)\int d\vec{\chi}\int^{\infty}_{-{\rm x}^{0}+\sqrt{\left(\vec{\chi}+\vec{\rm x}\right)^{2}+u^{2}}}\left(\frac{({\rm x}+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{({\rm z}^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}d\chi^{0}$ (14.194)
where the first line is the integral over $\gamma^{+}_{P}(x_{1})$} and the second two lines are the integral over $\gamma^{-}_{P}(x_{1})$}. The challenging task of evaluating the three integrals above will be rewarded by an exceptionally simple final result. Additionally, the calculation will yield new integral formulae involving Bessel functions.
Let us start with the integral over $\gamma^{+}_{P}(x_{1})$}; in this case the relevant Fourier transform of the waves is as follows:
$\widetilde{\phi^{\pm}_{\lambda}}({\rm k})=\int(x(t\pm i\epsilon,\vec{\rm x},u)\cdot\xi(\eta))^{\lambda}e^{i{\rm k}{\rm x}}\,d{\rm x}\,=e^{i{\rm k}\eta}\int\left[\frac{u}{2}-\frac{({\rm y}^{0}\pm i\epsilon)^{2}-\vec{\rm y}^{2}}{2u}\right]^{\lambda}e^{i{\rm k}{\rm y}}d{\rm y}$ (14.195)
where ${\rm y}={\rm x}-\eta$ (and we left the dependence on $u$ of the lhs implicit). Let us chose ${\rm k}=({\rm k}^{0},\vec{0})$ and perform the radial integration first:
$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)=e^{i{\rm k}^{0}\eta^{0}}\frac{2\pi^{\frac{d-2}{2}}}{\Gamma(\frac{d-2}{2})}\int_{-\infty}^{\infty}d{\rm y}^{0}\int_{0}^{\infty}\left[\frac{u}{2}-\frac{({\rm y}^{0}+i\epsilon)^{2}-r^{2}}{2u}\right]^{\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}r^{d-3}\,dr=$ (14.196)
$=\frac{(2u)^{-\lambda}e^{i{\rm k}^{0}\eta^{0}}\pi^{\frac{d-2}{2}}\Gamma\left(\frac{2-d}{2}-\lambda\right)}{\Gamma(-\lambda)}\int_{-\infty}^{\infty}\left(u^{2}-{({\rm y}^{0}+i\epsilon)^{2}}\right)^{\frac{d-2}{2}+\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}d{\rm y}^{0}$ (14.197)
valid for $\mathrm{Re}\,(2\lambda+d)<2$ and $\mathrm{Re}\,d>2$ (these restrictions will be removed at the end when the integral will be considered in the sense of distributions).
When ${\rm k}^{0}>0$ the contour of integration may be pushed to infinity and the integral is shown to be zero. The geometric and physical reason behind this possibility is once more the already mentioned immersion of the Minkowskian tubes of the branes in the AdS chiral tuboids.
When ${\rm k}^{0}<0$ the above contour distortion is not possible and we get
$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)=\frac{(2\pi)^{\frac{d+1}{2}}}{\Gamma\left(-\lambda\right)}\theta(-{\rm k}^{0}){|{\rm k}^{0}|}^{-\frac{d-1}{2}-\lambda}e^{i{\rm k}^{0}\eta^{0}}u^{\frac{d-1}{2}}J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u).$ (14.198)
For what follows it is also useful to take in Eq. (14.197) the ${\rm y}^{0}$ integration first:
$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)=\frac{4\sqrt{2}|{\rm k}^{0}|^{-\frac{1}{2}-\lambda}e^{i\eta{\rm k}^{0}}\pi^{\frac{d+1}{2}}u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)\Gamma(-\lambda)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right).$ (14.199)
By comparing Eqs. (14.198) and (14.199) we deduce the identity
$\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)=2^{\frac{d-4}{2}}\Gamma\left(\frac{d-2}{2}\right)|{\rm k}^{0}|^{1-\frac{d}{2}}u^{\frac{d-1}{2}+\lambda}J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u)$ (14.200)
valid when the integral converges. This identity is special instance of a known integral formula by Watson [21] and will be used later.
By restoring Lorentz invariance we obtain the Fourier integral representation of the wave (which we write in a way ready to be used in our calculation):
$(x({\rm z}^{\prime},u^{\prime})\cdot\xi(\eta))^{-\lambda-d+1}=\left(\frac{{u^{\prime}}^{2}-({\rm z}^{\prime}-\vec{\eta})^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}=$ (14.201)
$=\frac{(2\pi)^{\frac{3-d}{2}}{u^{\prime}}^{\frac{d-1}{2}}}{\Gamma(\lambda+d-1)}\int e^{-i{\rm k}({\rm z}^{\prime}-\eta)}\theta(-{\rm k}^{0})\theta({\rm k}^{2})\left(\sqrt{{\rm k}^{2}}\right)^{\frac{d-1}{2}+\lambda}J_{\frac{d-1}{2}+\lambda}(\sqrt{{\rm k}^{2}}\,u^{\prime})d{\rm k},$ (14.202)
where the factor $\theta(-{\rm k}^{0})$ guarantees the integral convergence as long as ${\rm z}^{\prime}\in T_{+}$.
We are now ready to evaluate the first line (14.194) (i.e. the integral over the cycle $\gamma^{(+)}_{P}(x_{1})$). Without loss of generality we may simplify the calculation by choosing ${\rm x}=0$ i.e. $x_{1}=x_{1}(0,u).$ Proceeding as in the two-dimensional case (see Eq. (13.172)) we are led to examine the integral
$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}dr\int_{-\sqrt{r^{2}+u^{2}}}^{\sqrt{r^{2}+u^{2}}}e^{i{\rm k}^{0}\eta^{0}}\left(\frac{u^{2}+r^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}d\eta^{0}=$ (14.204)
$=\frac{2\sqrt{2}\pi^{\frac{d-1}{2}}|{\rm k}^{0}|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)dr$ (14.205)
where we set $|\vec{\eta}|=r$ and integrated over the angles; we have assumed $\mathrm{Re}\,\lambda>-1$.
Though similar to (14.199) the integral on the rhs of (14.206) seems to be unknown. Following the example of Eq. (14.199), we may evaluate it by integrating first the variable $r$ on the lhs of (14.206):
$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}dr\int_{-\sqrt{r^{2}+u^{2}}}^{\sqrt{r^{2}+u^{2}}}e^{i{\rm k}^{0}\eta^{0}}\left(\frac{u^{2}+r^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}d\eta^{0}=$ (14.207)
$=-\frac{(2\pi u)^{\frac{d-1}{2}}|{\rm k}^{0}|^{-\frac{d-1}{2}-\lambda}\Gamma(\lambda+1)}{\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)}\left(\cos(\pi\lambda)J_{\frac{d-1}{2}+\lambda}(u|{\rm k}^{0}|)-\sin\left(\frac{\pi d}{2}\right)J_{-\frac{d-1}{2}-\lambda}(u|{\rm k}^{0}|)\right).$ (14.208)
Comparison gives the following result (valid where the integral converges):
$\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)dr=$ (14.210)
$=\frac{2^{\frac{d-4}{2}}\pi\,|{\rm k}^{0}|^{-\frac{d-2}{2}}u^{\frac{d-1}{2}+\lambda}}{\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)\Gamma\left(2-\frac{d}{2}\right)}\left(\frac{\cos(\pi\lambda)}{\sin\left(\frac{d\pi}{2}\right)}J_{\frac{d-1}{2}+\lambda}(|{\rm k}^{0}|u)-J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u)\right).$ (14.212)
Restoring Lorentz invariance in Eq. (14.209) gives
$\int\int_{-\sqrt{u^{2}+\vec{\eta}^{2}}}^{\sqrt{u^{2}+\vec{\eta}^{2}}}\left(\frac{u^{2}+\vec{\eta}^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}e^{i{\rm k}\eta}d\eta^{0}d\eta=-\frac{(2\pi u)^{\frac{d-1}{2}}\left(\sqrt{{\rm k}^{2}}\right)^{-\frac{d-1}{2}-\lambda}\Gamma(\lambda+1)}{\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)}\times$ (14.213)
$\times\theta({\rm k}^{2})\left(\cos(\pi\lambda)J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)-\sin\left(\frac{\pi d}{2}\right)J_{-\frac{d-1}{2}-\lambda}\left(u\sqrt{{\rm k}^{2}}\right)\right).$ (14.214)
By inserting Eqs. (14.203) and (14.214) into Eq. (14.194) we get the first part of the result computed in the special case when $x_{1}=x_{1}(0,u)$:
$\int\int_{-\sqrt{u^{2}+\vec{\eta}^{2}}}^{\sqrt{u^{2}+\vec{\eta}^{2}}}\left(\frac{u^{2}+\vec{\eta}^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}\left(\frac{{u^{\prime}}^{2}({\rm z}^{\prime}-\eta)^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}d\eta^{0}d\eta=$ (14.215)
$=-\frac{2\pi({uu^{\prime}})^{\frac{d-1}{2}}\Gamma(\lambda+1)}{\Gamma(\lambda+d-1)\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)}\int e^{-i{\rm k}{\rm z}^{\prime}}\theta(-{\rm k}^{0})\theta({\rm k}^{2})\frac{}{}\times$ (14.216)
$\times\left(\cos(\pi\lambda)J_{\frac{d-1}{2}+\lambda}(u\sqrt{{\rm k}^{2}})-\sin\left(\frac{\pi d}{2}\right)J_{-\frac{d-1}{2}-\lambda}(u\sqrt{{\rm k}^{2}})\right)J_{\frac{d-1}{2}+\lambda}(u\sqrt{{\rm k}^{2}})d{\rm k}.$ (14.217)
Let us now focus on the other cycle $\gamma^{(-)}_{P}(x_{1})$. Here the situation is more complicated because the cycle goes through infinity and, as in Eq. (13.179), we need the Fourier transforms of the waves whose phase is vanishing at infinity. The analogous of Eq. (13.177) is here
$\lim_{\chi^{0}\to\pm\infty}\text{Arg}\left(\frac{({\rm x}^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)=0.$ (14.218)
Proceeding as before we compute the relevant Fourier transforms of the waves with the correct phases at infinity:
$\left(\frac{({\rm x}^{\prime}+\chi)^{2}+{u^{\prime}}^{2}}{2u^{\prime}}\right)_{\pm}^{-\lambda-d+1}=\pm{i(2\pi)^{\frac{1-d}{2}}\left(e^{\mp 2i\pi(d+\lambda)}-1\right)}{\Gamma(2-\lambda-d)}\times$ (14.219)
$\times{u^{\prime}}^{\frac{d-1}{2}}\int e^{-i{\rm k}({\rm x}^{\prime}+\chi)}\theta(-{\rm k}^{0})\theta({\rm k}^{2})\left(\sqrt{{\rm k}^{2}}\right)^{\frac{d-1}{2}+\lambda}J_{\frac{d-1}{2}+\lambda}(\sqrt{{\rm k}^{2}}\,u^{\prime})d{\rm k}.$ (14.220)
In the last step we evaluate the integral
$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}dr\int_{\sqrt{r^{2}+u^{2}}}^{\infty}e^{-i{\rm k}^{0}\chi^{0}}\left(\frac{{\chi^{0}}^{2}-u^{2}-r^{2}}{2u}\right)^{\lambda}d\chi^{0}=$ (14.221)
$=-\frac{\sqrt{2}\,\pi^{\frac{d-1}{2}}\Gamma(\lambda+1)u^{-\lambda}|{\rm k}^{0}|^{-\lambda-\frac{1}{2}}}{\Gamma\left(\frac{d}{2}-1\right)}\times$ (14.222)
$\int_{0}^{\infty}r^{d-3}dr\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}\left(\frac{J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)}{\cos(\pi\lambda)}+(\mbox{tg\,}(\pi\lambda)-i)J_{-\lambda-\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)\right)=$ (14.223)
$=\frac{(2\pi u)^{\frac{d-1}{2}}\Gamma(\lambda+1)|{\rm k}^{0}|^{-\frac{d-1}{2}-\lambda}}{2}\left(\frac{J_{\frac{d-1}{2}+\lambda}(u|{\rm k}^{0}|)}{\cos\left(\frac{\pi d}{2}+\pi\lambda\right)}-\left(\mbox{tg\,}\left(\frac{\pi d}{2}+\pi\lambda\right)-i\right)J_{-\frac{d-1}{2}-\lambda}(u|{\rm k}^{0}|)\right)$ (14.224)
where we used both Eqs. (14.199) and (14.212); as before we have assumed $\mathrm{Re}\,\lambda>-1$ and set ${\rm k}=({\rm k}^{0},0)$ (with ${\rm k}^{0}<0$). Similarly
$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}dr\int^{-\sqrt{r^{2}+u^{2}}}_{-\infty}e^{-i{\rm k}^{0}\chi^{0}}\left(\frac{{\chi^{0}}^{2}-u^{2}-r^{2}}{2u}\right)_{-}^{\lambda}d\chi^{0}=$ (14.226)
$=\frac{(2\pi u)^{\frac{d-1}{2}}\Gamma(\lambda+1)|{\rm k}^{0}|^{-\frac{d-1}{2}-\lambda}}{2}\left(\frac{J_{\frac{d-1}{2}+\lambda}(u|{\rm k}^{0}|)}{\cos\left(\frac{\pi d}{2}+\pi\lambda\right)}-\left(\mbox{tg\,}\left(\frac{\pi d}{2}+\pi\lambda\right)+i\right)J_{-\frac{d-1}{2}+\lambda}(u|{\rm k}^{0}|)\right).$ (14.227)
Taking into account Eq.(14.220) we get
$\int_{\gamma^{-}_{P}(x_{1})}\left(\frac{{\chi^{0}}^{2}-u^{2}-\vec{\chi}^{2}}{2u}\right)\left(\frac{({\rm z}^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{-}^{-\lambda-d+1}d\chi+$ (14.229)
$+$ $\int_{\gamma^{+}_{P}(x_{1})}\left(\frac{{\chi^{0}}^{2}-u^{2}-\vec{\chi}^{2}}{2u}\right)\left(\frac{({\rm z}^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{+}^{-\lambda-d+1}d\chi=$ (14.230)
$=$ $\frac{2\,\Gamma(\lambda+1)\sin(\pi(d+\lambda))\Gamma(-d-\lambda+2)\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{\cos\left(\frac{\pi d}{2}+\pi\lambda\right)}\int e^{-i{\rm k}{\rm z}}\theta(-{\rm k}^{0})\theta({\rm k}^{2})\times$ (14.232)
$\times J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right)\left(\cos(\pi(d+\lambda))J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)+\sin\left(\frac{\pi d}{2}\right)J_{-\frac{d-1}{2}-\lambda}\left(u\sqrt{{\rm k}^{2}}\right)\right).$
By summing Eqs. (14.217) and (LABEL:yuppy) and changing $\rm k$} to $\rm k$} we obtain the diagonalization we were looking for:
$\int_{\gamma_{P}(x_{1})}(x_{1}\cdot\xi)^{\lambda}(\xi\cdot z_{2})^{-\lambda-d+1}d\mu_{\gamma}(\xi)=$ (14.234)
$=$ $-\frac{4\pi\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{\Gamma(d+\lambda-1)}\int e^{-i{\rm k}({\rm x}-x^{\prime})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}(u\sqrt{k^{2}})J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{k^{2}}\right)$ (14.235)
Finally we restore the correct normalization of the two-point function by multiplying both sides of the above formula by $c_{d}(\lambda)$ and get
$W^{(d)}_{\lambda}(z_{1}({\rm z}_{1},u),z_{2}({\rm z}_{2},u^{\prime}))=$ (14.237)
$=\frac{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{2(2\pi)^{d-2}}\int e^{-i{\rm k}({\rm z}_{1}-{\rm z}_{2})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right)d{\rm k}.$ (14.238)
In the final step we write
$\theta({\rm k}^{2})=\int_{0}^{\infty}\delta({\rm k}^{2}-m^{2})dm^{2}$ (14.239)
and insert Eqs. (1.2) and (14.239) into Eq. (14.238) to get
$W^{(d)}_{\lambda}(z_{1}({\rm z}_{1},u),z_{2}({\rm z}_{2},u^{\prime}))=\frac{1}{2}\left(uu^{\prime}\right)^{\frac{d-1}{2}}\int W^{M_{d-1}}_{m}({\rm z}_{1},{\rm z}_{2})J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2}.$ (14.240)
That this expression solves the Klein-Gordon equation
$u^{2}\Box_{M}\phi-u^{d}{\partial_{u}}(u^{2-d}\partial_{u}\phi)+m^{2}_{\lambda}\phi=0$ (14.241)
may be verified by using the relation
$u^{d}{\partial_{u}}(u^{2-d}\partial_{u}\left(u^{\frac{d-1}{2}}J_{\frac{d-1}{2}+\lambda}\left(mu\right)\right)=u^{\frac{d-1}{2}}\left(\lambda(d+\lambda-1)-m^{2}u^{2}\right)J_{\frac{d-1}{2}+\lambda}(mu).$ (14.242)
The remarkable formula (14.240) shows that the maximally analytic AdS two-point function can be represented as a Källén-Lehmann superposition of maximally analytic two-point functions defined in a Minkowski space with one dimension less, explicitly linking the analytic and geometric structures of AdS and Minkowski spacetimes. Eq. (14.240) was anticipated in [23, 24] by employing Hilbert space methods suitable for generic warped manifolds; the value of the derivation in [23, 24] was also to clarify the Breitenlohner and Friedmann phenomenon [9] in terms of the self-adjointness of a one-dimensional transverse Schrodinger operator; however, a crucial part of the proof was based on a known Hankel transform of the Legendre function $Q$. The present derivation is fully self-contained and specifically tailored to the AdS geometry.
Eq. (14.240) may be used to show the positive-definiteness of the two-point function. Let us briefly sketch how by introducing a sesquilinear form in the space ${\cal T}(AdS_{M})$ of test function which are smooth and vanish outside $AdS_{M}$ as indicated in in Eq. (5.39):
$\langle f,f\rangle=\int W_{\lambda}^{(d)}(x_{1},x_{2})\overline{f(x_{1})}f(x_{2})\sqrt{g(x_{1})}\sqrt{g(x_{2})}dx_{1}dx_{2}=$ (14.243)
$=\int d{\rm k}\frac{\theta({\rm k}^{0})\theta({\rm k}^{2})}{2(2\pi)^{d-2}}\int\overline{\hat{f}(k,u)}\hat{f}(k,u^{\prime})\left(uu^{\prime}\right)^{-\frac{d+1}{2}}J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right){du}\ {du^{\prime}}$ (14.244)
where
$\hat{f}({\rm k},u)=\int e^{i{\rm k}{\rm x}}{f(x({\rm x},u))}d{\rm x}$ (14.246)
is the Fourier transform the test function w.r.t the Poincaré coordinate $\rm x$}. When $\lambda$ is real the above expression becomes formally positive
$\langle f,f\rangle=\frac{1}{2(2\pi)^{d-2}}\int\theta({\rm k}^{0})\theta({\rm k}^{2})\left|\hat{F}_{\frac{d-1}{2}+\lambda}({\rm k})\right|^{2}d{\rm k}.$ (14.247)
where
$\widehat{F}_{\frac{d-1}{2}+\lambda}({\rm k})=\int_{0}^{\infty}{u^{-\frac{d+1}{2}}}{\hat{f}({\rm k},u)}J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right){du}$ (14.248)
A thorough study of the Hilbert space structure will presented elsewhere.
15 Green functions, Schwinger functions and Wick rotations
Eq. (14.240) may be used to promptly build a new explicit integral of the Schwinger function (8.80) extended to non coincident Euclidean points of $EAdS_{d}$:
$S^{(d)}_{\lambda}(z_{\scriptscriptstyle{E}},z^{\prime}_{\scriptscriptstyle{E}})=\frac{1}{2}\left(uu^{\prime}\right)^{\frac{d-1}{2}}\int S^{E_{d-1}}_{m}({\rm x}_{\scriptscriptstyle{E}},{\rm x}^{\prime}_{\scriptscriptstyle{E}})J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2},$ (15.249)
where
$S^{E_{d-1}}_{m}({\rm x}_{\scriptscriptstyle{E}})=\frac{1}{(2\pi)^{d-1}}\int\frac{e^{-i{\rm k}_{\scriptscriptstyle{E}}{\rm x}_{\scriptscriptstyle{E}}}}{{{\rm k}_{\scriptscriptstyle{E}}^{2}+m^{2}}}d{\rm k}_{\scriptscriptstyle{E}}$ (15.250)
is the Schwinger function of a (d1)(d-1)-dimensional massive Euclidean scalar field and ${\rm k}_{\scriptscriptstyle{E}}{\rm x}_{\scriptscriptstyle{E}}={\rm k}^{0}_{\scriptscriptstyle{E}}{\rm x}^{0}_{\scriptscriptstyle{E}}+\ldots+{\rm k}^{d-1}_{\scriptscriptstyle{E}}{\rm x}^{d-1}_{\scriptscriptstyle{E}}$ is the Euclidean scalar product.
There exists a commonly used integral representation of the AdS Schwinger functions called “split representation” [25, 26, 27, 28]; it may be obtained by first representing the Legendre function of the second kind $Q$ as a Mehler-Fock transform of the corresponding Legendre function of the first kind $P$ [19, 29] and then expanding $P$ in plane waves on the Lobachevsky manifold [17, 30]. The advantage of the representation (15.249) is that it has admits a counterpart on the reals that may be obtained by Wick rotation while the split representation cannot be Wick-rotated. Here is the corresponding representation of the Feynman propagator of the Klein-Gordon operator $\Box+m^{2}_{\lambda}$ in the real manifold $AdS_{M}$:
$G^{(d)}_{\lambda}(x_{1}({\rm x}_{1},u),x_{2}({\rm x}_{2},u^{\prime}))=\frac{1}{2}\left(uu^{\prime}\right)^{\frac{d-1}{2}}\int G^{M_{d-1}}_{m}({\rm x}_{1}-{\rm x}_{2})J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2}$ (15.251)
where
$G^{M_{d-1}}_{m}({\rm x})=-\frac{1}{(2\pi)^{d-1}}\int\frac{e^{-i{\rm k}{\rm x}}}{{{\rm k}^{2}-m^{2}+i\epsilon}}d{\rm k}$ (15.252)
is the Feynman propagator of the Klein-Gordon operator $\Box_{M}+m^{2}$ in a (d1)(d-1)-dimensional Minkowski spacetime $M_{d-1}$. The verification that this expression is indeed a propagator is easy but also instructive:
$(\Box+m^{2}_{\lambda})G^{(d)}_{\lambda}(x_{1},x_{2})=u^{2}\Box_{M}G^{(d)}_{\lambda}(x_{1},x_{2})-u^{d}{\partial_{u}}(u^{2-d}\partial_{u}G^{(d)}_{\lambda}(x_{1},x_{2}))+m^{2}_{\lambda}G^{(d)}_{\lambda}(x_{1},x_{2})=$ (15.253)
$=m^{2}_{\lambda}G^{(d)}_{\lambda}(x_{1},x_{2})+\frac{1}{2}u^{2}{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}\delta({\rm x}_{1}-{\rm x}_{2})\int_{0}^{\infty}J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2}+$ (15.255)
$-\frac{1}{2}m^{2}u^{2}{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}\int_{0}^{\infty}G^{M_{d-1}}_{m}({\rm x}_{1}-{\rm x}_{2})J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2}+$ (15.256)
$-\frac{1}{2}{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}\int_{0}^{\infty}G^{M_{d-1}}_{m}({\rm x}_{1}-{\rm x}_{2})\left(\lambda(\lambda+d-1)-m^{2}u^{2}\right)J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2}$ (15.257)
$=u^{d}\delta({\rm x}_{1}-{\rm x}_{2})\delta(u,u^{\prime})$ (15.259)
where we used Eq. (14.242) and the inversion theorem of the Hankel transform [22].
The relation of AdS Euclidean diagrams with real space diagrams [18, 26, 31] remains however poorly understood also because calculations based on the split representation cannot be Wick-rotated. Eqs. (15.249) and (15.251) may shed some new light on this issue. In particular they immediately implies that banana diagrams [32] and more generally diagrams with no external lines computed in the Euclidean manifold $EAdS_{d}$ can be Wick-rotated and coincide with the same diagrams calculated in $AdS_{M}$.
What is less obvious is whether a more general diagram computed by integrating only in a Poincaré patch $AdS_{M}$ produces an AdS invariant result. We want to suggest that this might be generally true by computing just the simplest one-line diagram i.e. the integral of the propagator (15.251) over $AdS_{M}$ w.r.t. one of the two points letting the other point fixed. The calculation in flat space is just trivial:
$\int_{M_{d-1}}G^{M_{d-1}}_{m}({\rm x}_{1}-{\rm x}_{2})d{\rm x}_{1}=-\frac{1}{(2\pi)^{d-1}}\int\frac{e^{-i{\rm k}({\rm x}_{1}-{\rm x}_{2})}}{{{\rm k}^{2}-m^{2}+i\epsilon}}d{\rm x}_{1}d{\rm k}=\frac{1}{m^{2}}.$ (15.260)
Using this result we can perform the integral of (15.251) over $AdS_{M}$:
$\int_{AdS_{M}}G^{d}_{\lambda}(x_{1},x_{2})\sqrt{g(x_{1})}\,dx_{1}=\int_{0}^{\infty}\int_{0}^{\infty}{\left(\frac{u^{\prime}}{u}\right)^{\frac{d-1}{2}}}\frac{J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)}{mu}{du}dm=$ (15.261)
$=\frac{2^{-\frac{d+1}{2}}\Gamma\left(\frac{\lambda}{2}\right)}{\Gamma\left(\frac{1}{2}(d+\lambda+1)\right)}\int_{0}^{\infty}{{(mu^{\prime})}^{\frac{d-1}{2}}J_{\frac{d-1}{2}+\lambda}(mu^{\prime})}\frac{dm}{m}=\frac{1}{\lambda(\lambda+d-1)}=\frac{1}{m^{2}_{\lambda}}.$ (15.262)
This is, mutatis mutandis, the same result as in flat space; the independence of the integral on $x_{2}$ is remarkable and makes the result AdS invariant even though $AdS_{M}$ is not invariant.
The corresponding Euclidean calculation, on the contrary, is expected to give an invariant result; using Eq. (8.80) we get [32]
$\int_{EAdS}S^{d}_{\lambda}(z^{\scriptscriptstyle E}_{1},z^{\scriptscriptstyle E}_{2})\sqrt{g(z^{\scriptscriptstyle E}_{1})}\,dz^{\scriptscriptstyle E}_{1}=\frac{2^{\frac{2-d}{2}}}{\Gamma\left({\frac{d}{2}}\right)}\int_{1}^{\infty}{}{e^{-i\pi\frac{d-2}{2}}}Q^{\frac{d-2}{2}}_{\frac{d-2}{2}+\lambda}(u)(u^{2}-1)^{\frac{d-2}{4}}du=\frac{1}{m_{\lambda}^{2}}$ (15.263)
i.e. the same result as in the above Lorentzian calculation restricted to $AdS_{M}$ (see [32] for details). The study of more complicated examples and in particular of Witten diagrams will be presented in a forthcoming paper.
16 Concluding remarks
AdS quantum field theory admits a formulation that is structurally rich as its Minkowskian Wightman counterpart provided one fully exploits it on the intrinsic complex geometry of AdS rather than working in specific coordinates. By introducing holomorphic AdS plane waves and identifying suitable relative homology cycles supported on the complex null cone, the scalar two-point function is expressed in a manifestly covariant, coordinate-free integral representation that implements covariance, locality, maximal analyticity in a unified way, while addressing some of the conceptual difficulties associated with closed timelike curves. Its diagonalisation in Poincaré coordinates, encoded in a Källén–Lehmann-type superposition over lower-dimensional Minkowski two-point functions with Bessel kernels, not only yields explicit Schwinger and Feynman propagators but also helps to clarify the status of the Poincaré patch and the precise sense in which Euclidean Feynman diagrams on Lobachevski space can be Wick-rotated to diagrams supported entirely in a single AdS patch without loss of AdS covariance.
Altogether, the framework provides a robust and versatile starting point for perturbative AdS QFT and offers a new setting in which to revisit and refine holographic constructions that rely on AdS correlation functions and their boundary limits.
Appendix A Algebraic characterization of the chiral cones
In order to characterize the chiral cones $\cal C$$_{\pm}$ by algebraic inequalities we write $M=gM_{0d}g^{-1}$; given a vector $\zeta$ belonging to, say, $\cal C$$_{+}$ there is a $\xi\in{\cal C}_{d}$ such that
$\zeta=e^{(s+it)M}\,\xi=g\ e^{(s+it)M_{0d}}g^{-1}\xi=g\ e^{(s+\alpha+it)M_{0d}}\xi^{\prime}=g^{\prime}\zeta^{\prime},$ (A.264)
where
[…]
Inspection shows that
[…]
By $G_{0}$-invariance Eq. (A.264) implies that also $\mathrm{Im}\,\zeta\cdot\mathrm{Im}\,\zeta>0$. Furthermore, for any null complex vector $\zeta=\xi+i\chi\in{\cal C}_{d}^{(c)}$ there holds the identity
$\epsilon(\zeta)^{2}=(\chi^{0}\xi^{d}-\chi^{d}\xi^{0})^{2}=(\chi\cdot\chi)(|\vec{\xi}|^{2}+|\vec{\chi}|^{2})+(\chi\cdot\chi)^{2}+|\vec{\xi}|^{2}|\vec{\chi}|^{2}-(\vec{\xi}\cdot\vec{\chi})^{2}.$ (A.267)
Eq. (A.267) shows that $\epsilon(\zeta)^{2}$ is strictly positive when $\chi\cdot\chi$ is strictly positive; since $G_{0}$ is connected, Eq. (A.264) in turn implies that there is a continuous path connecting $\zeta^{\prime}$ and $\zeta$ and therefore also $\epsilon(\zeta)>0$. Summarizing, the open cone
${\cal C}=\{\zeta=\xi+i\chi\in{\cal C}_{d}^{(c)}:\ \xi\cdot\xi=\chi\cdot\chi>0,\ \xi\cdot\chi=0\}$ (A.268)
is the disjoint union of two $G_{0}$-invariant subsets ${\cal C}^{\prime}_{\pm}=\{\zeta=\xi+i\chi\in{\cal C},\ \epsilon(\zeta)\gtrless 0\}$ and ${\cal C}_{\pm}\subset{\cal C}^{\prime}_{\pm}$.
To show the opposite inclusion, let us suppose that $\zeta=\xi+i\chi\in{\cal C}^{\prime}_{+}$; there is a $t>0$ such that $\xi\cdot\xi=\chi\cdot\chi=\mbox{sh\,}^{2}t$. A transformation $g_{0}\in G_{0}$ can be found such that $\xi=g_{0}\,\mbox{sh\,}(t)\,e_{d}$ and $\chi=g_{0}\,\mbox{sh\,}(t)\,e_{0}$. By posing
$g(t)=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&1_{d-2}&0&0\\ 0&0&\coth(t)&1/\mbox{sh\,}t\\ 0&0&1/\mbox{sh\,}t&\coth(t)\\ \end{array}\right)$ (A.269)
(note that $g(t)g(t^{\prime})\not=g(t+t^{\prime})$) we get that
$g(t)(i\,\mbox{sh\,}(t)\,e_{0}+e_{d-1}+\mbox{ch\,}(t)\,e_{d})=e^{itM_{0d}}(e_{d-1}+e_{d})$ (A.270)
and therefore
$\zeta=\xi+i\chi=g_{0}g(t)^{-1}e^{itM_{0d}}(e_{d-1}+e_{d})\in{\cal C}_{+}.$ (A.271)
We may thus identify the chiral cones with the following subsets of ${\cal C}_{d}^{(c)}$:
${\cal C}_{\pm}=\{\zeta=\xi+i\chi\in{\bf C}^{d+1}_{2}:\ \xi\cdot\xi=\chi\cdot\chi>0,\ \xi\cdot\chi=0,\ \epsilon(\zeta)\gtrless 0\}.$ (A.272)
The above construction also shows that the chiral cones may be generated simply by acting on the point $e_{d-1}+e_{d}$ as follows:
${\cal C}_{\pm}=\{g\,\exp(itM_{0d})(e_{d-1}+e_{d})\ :\ g\in G_{0},\ \ \ t\gtrless 0\}.$ (A.273)
By using the complex spherical coordinates of the cone ${\cal C}_{d}^{(c)}$
$\zeta(t+is,\lambda,\vec{a}+i\vec{b})=\left\{\begin{array}[]{l}\lambda\sin(t+is),\cr\lambda\ (\vec{a}+i\vec{b}),\cr\lambda\cos(t+is),\end{array}\right.\ \ \ |\vec{a}|^{2}-|\vec{b}|^{2}=1,\ \ \vec{a}\cdot\vec{b}=0$ (A.274)
we see that also $\cal C$$_{-}$ and $\cal C$$_{+}$ are semi-tubes in ${\bf C}^{d}$ bordered by surfaces $s=s_{\pm}(|\vec{b}^{2}|)$ :
${{\cal C}_{+}}\ :\ \ \ \ s>+{\rm arcsh}(|\vec{b}|)$ (A.275)
${{\cal C}_{-}}\ :\ \ \ \ s<-{\rm arcsh}(|\vec{b}|)$ (A.276)
$\widetilde{\cal C}_{\pm}$ are given by the same formulae where the real variable $t$ has been unfolded.
Acknowledgments. I want to express my gratitude to Sergio Cacciatori for numerous suggestions. I would also like to thank the Institut des Hautes Etudes Scientifiques (Bures sur Yvette) and the Physics Department of Fudan University for their warm hospitality and support while writing this paper.

### v2

@@
 Here is, deferring all detailed explanations, our formula for the two-point function of a massive scalar field:
-$W^{(d)}_{\lambda}(z_{1},z_{2})=c_{d}(\lambda)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-d+1}d\mu_{\gamma}(\zeta).$ (1.1)
+$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=$ $c_{d}(\lambda)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-d+1}d\mu_{\gamma}(\zeta)=$ (1.1)
+$=$ $c_{d}(\nu)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{-\frac{d-1}{2}+\nu}(z_{2}\cdot\zeta)^{-\frac{d-1}{2}-\nu}d\mu_{\gamma}(\zeta).$ (1.2)
 This integral representation is a superposition of plane waves; it does not separate space and time variables, nor does it depend on any particular coordinate system on the AdS manifold or any specific class of special functions.
@@
 Correlation functions for tensors and spinors can be straightforwardly derived by applying suitable differential operators to the right-hand side of this expression.
-The significance of the new representation (1.1) is evident: it provides the closest possible analogue, within the AdS framework, to the plane-wave Fourier representation of the two-point function for a massive scalar field in (complex) Minkowski space $M_{d}^{(c)}$ with positive energy spectrum [12]:
+The significance of the new representation (1.1) is evident: it provides the closest possible analogue, within the AdS framework, to the plane-wave Fourier representation of the two-point function for a massive scalar field in (complex) Minkowski space $M_{d}^{(c)}$ with positive energy spectrum [12] which we quote here
 ${W}^{{M}_{d}}_{m}({{\rm z}}_{1},{{\rm z}}_{2})=\frac{1}{(2\pi)^{d}}\int e^{-i{\rm p}\cdot{\rm z}_{1}}e^{i{\rm p}\cdot{\rm z}_{2}}\theta({\rm p}^{0})\delta({\rm p}^{2}-m^{2})d{\rm p}.$ (1.3)
+and to the plane wave expansion for the de Sitter quantum fields we have constructed long ago [13, 14].
 Manifestly covariant momentum space representations of quantum fields provide the physically intuitive ingredients of Feynman diagrammatics [15] and a great deal of the predictive power of quantum field theory depends on them.
 Furthermore, many crucial mathematical developments such as the PCT and the Spin-Statistics theorems [12] and the Euclidean formulation of quantum field theory [16] are based on the plane-wave Fourier representation of the correlation functions and the positivity of the energy spectrum.
-The importance of the Fourier representation (1.2) and of the corresponding representation of the Feynman propagator cannot be overestimated and we expect that (1.1) to occupy the same position in AdS QFT.
+The Fourier representation (1.3) and the related representation of the Feynman propagator are of paramount importance.
+We expect (1.1) to be just as fundamental in the framework of AdS QFT.
 The second major topic addressed in this work pertains to the horospherical, or Poincaré, coordinate system, which covers exactly one half of the AdS manifold.
@@
 the Poincaré coordinates and show how this transform diagonalizes the two-point function (1.1).
-We present the final, remarkably elegant result here deferring detailed explanations to Sections 13 and 14:
+We present here the final result, which exhibits remarkable elegance, while deferring all detailed derivations to Sect.
+13 and 14:
 $W^{(d)}_{\lambda}(z_{1}({\rm z}_{1},u),z_{2}({\rm z}_{2},u^{\prime})=$ (1.4)
-$=\frac{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{2(2\pi)^{d-2}}\int e^{-i{\rm k}({\rm z}_{1}-{\rm z}_{2})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right)d{\rm k}$ (1.4)
-The crucial role played by the global analyticity properties of the plane waves, which are inherited from our foundational axioms, will appear clearly in this calculation.
-Moreover, this analysis yields an unexpected insight: Euclidean Feynman diagrams constructed on Lobachevski space can be mapped, via Wick rotation, onto diagrams entirely supported in the Poincaré patch.
-Multiple examples are given that substantiate the conjecture that any Feynman or Witten diagram computed by integration solely over a Poincaré patch of the real AdS manifold remains AdS covariant, even though the patch itself is not invariant.
+$=\frac{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{2(2\pi)^{d-2}}\int e^{-i{\rm k}({\rm z}_{1}-{\rm z}_{2})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right)d{\rm k}.$ (1.5)
+Some historical remarks on this integral representation may help to put it in perspective.
+The Euclidean counterpart of Eq.
+(1.5), which represents the Schwinger propagator, appears in Bateman’s tables of integral transforms [17, Eq.
+(12), p.
+64].
+There, a Legendre function of the second kind [18], with a rather intricate argument, is identified as the Hankel transform of a product of two Bessel functions.
+A particular case of Eq.
+(1.5), corresponding to spacetime dimension $d=2$, is discussed in Watson’s classic treatise [19], together with indications of the proof.
+A similar integral—Watson notes—had already been discovered by Kirchhoff as early as 1853.
+Watson also credits other particular cases to [20, 21, 22, 23] and refers to all of them as Lipschitz–Hankel integrals.
+Formula (1.5) may therefore be called a Lipschitz–Hankel formula [24], or simply a Hankel formula, as we shall do throughout the rest of this paper.
+In the AdS literature, the Euclidean Hankel formula for the Schwinger propagator was reported in an appendix [27, Eq.
+(A.3)], where it appears as an intermediate step toward reconstructing a more conventional representation of the Feynman propagator, always in terms of Bessel functions, previously introduced by [28].
+The geometrical interpretation in anti-de Sitter (AdS) space of the Hankel formula (1.5), together with its relation to the maximally analytic Wightman functions of AdS quantum field theory [10]—which also include the Euclidean Schwinger propagator—was established shortly thereafter in [25, 26].
+This result was obtained by employing complexified Poincaré coordinates and exploiting the general analyticity properties of Wightman functions that satisfy the AdS spectral condition [10].
+From a technical perspective, the Hankel formula was derived in [25] by solving a Schrödinger-type equation in the spacelike coordinate transverse to the Poincaré foliation.
+A key ingredient in this derivation is Bateman’s formula, whose geometrical interpretation was clarified in that work.
+This approach also provides insight into the so-called Breitenlohner–Freedman phenomenon [9].
+In fact, the operator appearing in the transverse Schrödinger problem is not essentially self-adjoint, and its possible self-adjoint extensions correspond to different admissible quantizations in the regime where two such schemes coexist, thereby offering an alternative interpretation of the Breitenlohner–Freedman phenomenon.
+The Lorentzian Hankel formula for the Feynman propagator, which had been somewhat overlooked for a period of time, was eventually revived and brought to general attention in [29, 30].
+These studies not only reintroduced the formula but also demonstrated its significant utility in calculations of propagators, correlation functions, and scattering amplitudes in the AdS background.
+Since then, the formula has become part of the standard AdS toolbox.
+Here, the Hankel formula and its untapped potential role in AdS QFT are brought into sharper focus.
+In particular, its connection to the complex structure of the AdS manifold, the geometry of the complex cone, and the plane-wave solutions of the Klein–Gordon operator is fully elucidated.
+This perspective highlights the crucial role played by the global analyticity properties of the plane waves, which are inherited from the AdS spectral condition [10], and makes their significance evident in the context of the new derivation.
+Moreover, our analysis uncovers an unexpected and intriguing insight.
+It demonstrates that Euclidean Feynman diagrams constructed in Lobachevsky space can, through a standard Wick rotation, be mapped onto diagrams entirely supported within the Poincaré patch of AdS.
+Several concrete examples are provided that support the conjecture that any Feynman or Witten diagram, computed by integration solely over a Poincaré patch of the real AdS manifold, continues to respect full AdS covariance, even though the Poincaré patch itself is not invariant under the full isometry group.
 The structure of the paper is as follows: in Section 2, we describe the geometric and group-theoretical foundations, detailing the real AdS spacetime as an embedded manifold, its complexification, and the universal covering of the real manifold.
@@
 The concluding remarks reflect on the broader significance of the results, their potential applications to perturbative quantum field theory in AdS, and the continuing evolution of both mathematical and physical understanding in this area.
+Finally in Appendix A we discuss some technical details and in Appendix B we revisit the so called ”split representation” of the Euclidean propagator
 2 AdS and its covering: geometrical setup
@@
 $AdS_{d}=\{x,\ x^{2}=x\cdot x=R^{2}\}.$ (2.8)
+The restriction to $AdS_{d}$ of the pseudo-Riemannian metric $g_{\mu\nu}dx^{\mu}dx^{\nu}$ is a Lorentzian metric with signature (+,,,)(+,\ -$\ldots,$-).
 $G_{0}$ acts transitively on $AdS_{d}$; we chose
 $e_{d}=(0,\ldots,0,R)$ (2.9)
-to be the distinguished origin of $AdS_{d}$; in the folowing we will take for simplicity a unit radius $R=1$.
-The restriction to $AdS_{d}$ of the pseudo-Riemannian metric $g_{\mu\nu}dx^{\mu}dx^{\nu}$ is a Lorentzian metric with signature (+,,,)(+,\ -$\ldots,$-).
+to be the distinguished origin of $AdS_{d}$; in the following we will take for simplicity a unit radius $R=1$.
 The complex $d$-dimensional anti-de Sitter manifold $AdS_{d}^{(c)}$ can analogously be defined by its embedding in the complex space ${\bf C}^{d+1}_{2}$ endowed with the metric $g_{\mu\nu}$:
-$AdS_{d}^{(c)}=\{z=x+iy\in{\bf C}^{d+1}_{2},\ z^{2}=1\}=\{z=x+iy,\ x^{2}-y^{2}=1,\ x\cdot y=0\}.$ (2.9)
+$AdS_{d}^{(c)}=\{z=x+iy\in{\bf C}^{d+1}_{2},\ z^{2}=z\cdot z=1\}=\{z=x+iy,\ x^{2}-y^{2}=1,\ x\cdot y=0\}.$ (2.10)
 The complexification $G_{0}^{(c)}$ of the real AdS group acts transitively $AdS_{d}^{(c)}$.
@@
 However, although the full space $AdS_{d}^{(c)}$ itself has a trivial covering, $G_{0}$ and $AdS_{d}$ admits nontrivial covering group $\widetilde{G}_{0}$ and covering space $\widetilde{}AdS_{d}$; also, the important tubular domains $\cal Z$$_{\pm}$ of $AdS_{d}^{(c)}$ defined below in Section 4 admit nontrivial coverings.
-It is possible to single out in $AdS_{d}^{(c)}$ a Euclidean manifold $EAdS_{d}$; we choose the connected real submanifold defined by taking $z^{0}=ix^{0},\ x^{1},\ldots,x^{d}$ real, $x^{d}>0$:
-$z^{\scriptscriptstyle{E}}=(ix^{0},\vec{x},x^{d}).$ (2.10)
-It is the upper sheet of the two-sheeted hyperboloid with equation
-$(x^{d})^{2}-(x^{0})^{2}-{\vec{x}}^{2}=1,\ \ \ x^{d}\geq 1$ (2.11)
-equipped with the Riemannian metric induced by the scalar product (2.5). Global coordinates may be introduced on $AdS_{d}$ by the map
-$(t,\,\vec{x})\mapsto\left(\sqrt{1+\vec{x}^{2}}\,\sin t,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,\cos t\right)$ (2.12)
+It is possible to single out in $AdS_{d}^{(c)}$ a Euclidean manifold $EAdS_{d}$; we choose the connected real submanifold
+$EAdS_{d}=\{z^{\scriptscriptstyle{E}}=(ix^{0},\vec{x},x^{d})\in AdS_{d}^{(c)},\ \ x^{d}\geq 1\}$ (2.11)
+equipped with the Riemannian metric induced by the scalar product (2.6). It is the upper sheet of the two-sheeted hyperboloid with equation
+${x^{d}}^{2}-{x^{0}}^{2}-{|\vec{x}|}^{2}=1$ (2.12)
+i.e. it is a copy of the Lobachevsky space $H_{d}$.
+Global coordinates may be introduced on $AdS_{d}$ by the map
+$(t,\,\vec{x})\rightarrow\left(\sqrt{1+\vec{x}^{2}}\,\sin t,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,\cos t\right)$ (2.13)
 where $t\in S_{1}={\bf R}/2\pi{\bf Z}$ and $\vec{x}\in{\bf R}^{d-1}$. Coordinates on $EAdS_{d}$ are obtained extending the above map to $S_{1}^{(c)}\times{\bf R}^{d-1}$ and taking $t=is$, namely
-$(s,\,\vec{x})\mapsto(i\sqrt{1+\vec{x}^{2}}\,{\rm sh}\ s,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,{\rm ch}\ s).$ (2.13)
+$(s,\,\vec{x})\rightarrow(i\sqrt{1+\vec{x}^{2}}\,{\rm sh}\ s,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,{\rm ch}\ s).$ (2.14)
 Lifting the coordinates on the covering ${\bf R}^{d}$ of $S^{1}\times{\bf R}^{d-1}$ provides a global coordinate system on $\widetilde{}AdS_{d}$.
@@
 (i) The Euclidean AdS manifold $EAdS_{d}$ is contained in ${\cal Z}_{+}\cup{\cal Z}_{-}\cup AdS_{d}$.
-(ii) The image of either of the domains $\cal Z$$_{+}$ or $\cal Z$$_{-}$ under the coordinate map $z\mapsto z^{d}=e_{d}\cdot z$ is the cut-plane $\Delta={\bf C}\setminus[-1,1]$.
-(iii) The image of the domain ${\cal Z}_{-}\times{\cal Z}_{+}$ (or ${\cal Z}_{+}\times{\cal Z}_{-}$) by the scalar product mapping $z_{1},z_{2}\mapsto z_{1}\cdot z_{2}$ is the cut-plane $\Delta$.
-(iv) The map $z_{1},z_{2}\mapsto z_{1}\cdot z_{2}$ of ${\cal Z}_{-}\times{\cal Z}_{+}$ onto $\Delta$ can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$.
+(ii) The image of either of the domains $\cal Z$$_{+}$ or $\cal Z$$_{-}$ under the coordinate map $z\rightarrow z^{d}=e_{d}\cdot z$ is the cut-plane $\Delta={\bf C}\setminus[-1,1]$.
+(iii) The image of the domain ${\cal Z}_{-}\times{\cal Z}_{+}$ (or ${\cal Z}_{+}\times{\cal Z}_{-}$) by the scalar product mapping $z_{1},z_{2}\rightarrow z_{1}\cdot z_{2}$ is the cut-plane $\Delta$.
+(iv) The map $z_{1},z_{2}\rightarrow z_{1}\cdot z_{2}$ of ${\cal Z}_{-}\times{\cal Z}_{+}$ onto $\Delta$ can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$.
 A more explicit characterization of the tuboids may be obtained by using the global complex coordinates of the complex AdS manifold that we write here as follows:
@@
 $4^{\prime\prime}$.
-Maximal analyticity property: the reduced two-point function $W(z_{1}\cdot z_{2})$ extends to a function analytic in the covering $\widetilde{\Delta}$ of the cut-plane $\Delta=\{{\bf C}\setminus[-1,1]\}.$ For theories single-valued on $AdS_{d}$, $W(z_{1}\cdot z_{2})$ is analytic in $\Delta$.
+Maximal analyticity property: the reduced two-point function $W(z_{1}\cdot z_{2})$ extends to a function analytic in the covering $\widetilde{\Delta}_{1}$ of the cut-plane $\Delta_{1}=\{{\bf C}\setminus[-1,1]\}.$ For theories single-valued on $AdS_{d}$, $W(z_{1}\cdot z_{2})$ is analytic in $\Delta_{1}$.
 Note that the two-point function of any field satisfying locality AdS invariance and the spectral condition enjoys maximal analyticity, as it happens in the Minkowski [12] and de Sitter cases [13, 14].
@@
 In particular the operators ${\cal M}_{\mu\nu}$ and ${\cal M}^{2}$ commute with the restriction to invariant submanifolds of ${\bf R}^{d+1}_{2}$ and also with dilations.
-If $f$ is a smooth function or a distribution homogeneous of degree 0 on the open set $\Omega=\{x\in{\bf R}^{d+1}_{2}\ :\ (x,x)>0\}$ then
+If $f$ is a smooth function or a distribution homogeneous of degree zero on the open set $\Omega=\{x\in{\bf R}^{d+1}_{2}\ :\ (x,x)>0\}$ then
 $\Box(f|_{AdS_{d}})=(\Box_{g}f)|_{AdS_{d}}=\frac{1}{2}{\cal M}^{2}f|_{AdS_{d}}$ (6.47)
@@
 ${\cal M}^{2}f(x)=-2\lambda(\lambda+d-1)f(x).$ (6.49)
-Eqs. (6.47) and (6.48) then imply that the restriction of $f$ to the anti de Sitter manifold $AdS_{d}$ satisfies the Klein-Gordon equation
-$(\Box+\lambda(\lambda+d-1))(f(x)|_{AdS_{d}})=0$ (6.49)
+Eqs. (6.48) and (6.49) then imply that the restriction of $f$ to the anti-de Sitter manifold $AdS_{d}$ satisfies the Klein-Gordon equation
+$(\Box+m^{2}_{\lambda})(f(x)|_{AdS_{d}})=0$ (6.50)
 with complex squared mass
 $m^{2}_{\lambda}=\lambda(\lambda+d-1).$ (6.51)
-For real $\lambda$ and real $d$ the squared mass attains a minimum value which is negative:
-$\lambda_{d}=-\frac{d-1}{2},\ \ \ \ m^{2}_{\lambda_{d}}=-\frac{(d-1)^{2}}{4}.$ (6.51)
-It may be useful to introduce the complex parameter $\nu=\nu_{\lambda}$ by subtracting $\lambda_{d}$:
-$\nu=\lambda+\frac{d-1}{2}$ (6.52)
-so that $\nu=0$ corresponds precisely to the minimal value of the squared mass
-$m_{\lambda}^{2}=\nu_{\lambda}^{2}-\frac{(d-1)^{2}}{4}.$ (6.53)
-In the following we will use $\lambda$ and $\nu$ interchangeably.
-Plane waves are the most important example of homogeneous solution of the massless equation in the ambient space:
-$x,\zeta\mapsto(x\cdot\zeta)^{\lambda}$ (6.54)
-where $\lambda$ is a complex number and $\zeta$ is a complex null vector of ${\cal C}_{d}^{(c)}$. Taking the restrictions to the anti de Sitter universe
-$\phi_{\lambda}(x,\zeta)=(x\cdot\zeta)^{\lambda},\ \ \ x\in AdS_{d}$ (6.55)
-satisfies the massive Klein-Gordon equation which we rewrite for clarity:
-$(\Box+m_{\lambda}^{2})\phi_{\lambda}(x,\zeta)=0.$ (6.56)
-The invariance of $m^{2}_{\lambda}$ under the involution
-$\lambda\to 1-d-\lambda$ (6.57)
-implies that, given the complex null vector $\zeta$, there is a second solution of Eq. (6.56) obtained by applying the involution to the exponent of the wave:
+Plane waves provide the most important example of homogeneous solutions of the massless equation in the ambient space:
+$x,\zeta\rightarrow(x\cdot\zeta)^{\lambda},$ (6.52)
+where $\lambda$ is a complex parameter and $\zeta$ is a complex null vector in ${\cal C}_{d}^{(c)}$.
+The restriction $\phi_{\lambda}(x,\zeta)$ of (6.52) to the anti–de Sitter manifold satisfies the massive Klein–Gordon equation (6.50):
+$\phi_{\lambda}(x,\zeta)=(x\cdot\zeta)^{\lambda},\qquad x\in AdS_{d}.$ (6.53)
+The invariance of $m_{\lambda}^{2}$ under the involution
+$\lambda\to\bar{\lambda}=1-d-\lambda$ (6.54)
+implies that, for any given complex null vector $\zeta$, there exists a second solution of Eq. (6.50), obtained by applying this involution to the exponent of the wave:
 $\left(\Box+m_{\lambda}^{2}\right)\phi_{1-d-\lambda}(x,\zeta)=0.$ (6.55)
-Now the main difficulty arises: although the complex parameter $\lambda$ in the above definition is unrestricted, the plane waves themselves are only locally well-defined on $AdS_{d}$.
-An exception occurs when $\lambda$ takes integer values; in that case the plane waves are globally well-defined.
-In the general case, however, it requires additional effort to construct single-valued plane waves on the covering space $\widetilde{}AdS_{d}$.
-This issue will be addressed in detail in the next section.
+To establish a connection with notations commonly used in the AdS–CFT literature, we introduce three additional complex parameters: $\nu$, $\Delta_{-}$, and $\Delta_{+}$.
+These parameters are related to the homogeneity degree $\lambda$ through
+$\lambda=-\frac{d-1}{2}+\nu=-\Delta_{-}.$ (6.56)
+Under the involution (6.54)
+$\bar{\lambda}=1-d-\lambda=-\frac{d-1}{2}-\nu=-\Delta_{+}$ (6.57)
+the complex parameter $\nu$ simply changes sign,
+$\nu\rightarrow-\nu;$ (6.58)
+the involution does not coincide with complex conjugation unless $\nu$ is purely imaginary, which corresponds to a tachyonic wave devoid of physical interpretation in AdS.
+We may further clarify the relation between the homogeneity degree $\lambda$ and the scaling dimnsion $\Delta_{-}$ by using the Poincaré coordinates (12.155) and (12.164) to parametrize the plane wave:
+$(x(t,u)\cdot\xi(\eta))^{\lambda}=(x(t,u)\cdot\xi(\eta))^{-\frac{d-1}{2}+\nu}=\left(\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u}\right)^{-\Delta_{-}}\simeq\frac{u^{\Delta_{-}}}{(\left({{\rm x}}-{\eta}\right)^{2})^{\Delta_{-}}}$ (6.59)
+$(x(t,u)\cdot\xi(\eta))^{\bar{\lambda}}=(x(t,u)\cdot\xi(\eta))^{-\frac{d-1}{2}-\nu}=\left(\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u}\right)^{-\Delta_{+}}\simeq\frac{u^{\Delta_{+}}}{(\left({{\rm x}}-{\eta}\right)^{2})^{\Delta_{+}}}$ (6.60)
+(unless $\rm x$} and $\eta$ are colinear). The squared mass can then be written as
+$m_{\lambda}^{2}=\nu^{2}-\frac{(d-1)^{2}}{4}=-\Delta_{+}\Delta_{-}.$ (6.61)
+Let us assume that the spacetime dimension parameter $d$ is real. For real $\lambda$ (equivalently, for real $\nu$), the squared mass (6.61) reaches its minimum negative value at $\nu=0$:
+$m^{2}_{\min}=-\frac{(d-1)^{2}}{4}.$ (6.62)
+This minimum corresponds to the homogeneity degree
+$\lambda_{d}=-\frac{d-1}{2}.$ (6.63)
+In the following, we will use $\lambda$ and $\nu$ interchangeably.
+The main challenge now emerges: although the complex parameter $\lambda$ in the above definition can take any value, the plane waves themselves are only locally well-defined on AdSd.
+An exception occurs when $\lambda$ is an integer; in this case, the plane waves are globally well-defined.
+For non-integer $\lambda$, however, constructing single-valued plane waves requires additional effort on the covering space $\widetilde{\text{AdS}}_{d}$.
+This issue will be discussed in detail in the following section.
 7 Chiral cones
@@
   2. 2.
-    The image of either $\cal C$$_{+}$ or $\cal C$$_{-}$ under the coordinate map $\zeta\mapsto\zeta^{d}=e_{d}\cdot\zeta$ is the punctured complex-plane ${\bf C}^{*}={\bf C}\setminus\{0\}$.
+    The image of either $\cal C$$_{+}$ or $\cal C$$_{-}$ under the coordinate map $\zeta\rightarrow\zeta^{d}=e_{d}\cdot\zeta$ is the punctured complex-plane ${\bf C}^{*}={\bf C}\setminus\{0\}$.
 This fact is intuitively clear as the chiral cones are the limit of the chiral tuboids when the radius $R$ tends to zero.
@@
     Eqs.
-(7.64) and (7.66) then imply that the image of ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$ by the scalar product mapping $z,\zeta\mapsto z\cdot\zeta$ is also the punctured complex-plane $\bf C$$^{*}$ and the same is obviously true for ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$.
+(7.69) and (7.71) then imply that the image of ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$ by the scalar product mapping $z,\zeta\rightarrow z\cdot\zeta$ is also the punctured complex-plane $\bf C$$^{*}$ and the same is obviously true for ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$.
 In words: the scalar product $z\cdot\zeta$ cannot vanish when either $z\in{\cal Z}_{-}$ and $\zeta\in\overline{{\cal C}}_{+}$ or when $z\in{\cal Z}_{+}$ and $\zeta\in\overline{{\cal C}}_{-}$.
   4. 4.
-    The map $z,\zeta\mapsto z\cdot\zeta$ of ${\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto $\bf C$$^{*}$ can be lifted to a map of $\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto the covering $\widetilde{\bf C}^{*}$ of the punctured plane $\bf C$$^{*}$.
+    The map $z,\zeta\rightarrow z\cdot\zeta$ of ${\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto $\bf C$$^{*}$ can be lifted to a map of $\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto the covering $\widetilde{\bf C}^{*}$ of the punctured plane $\bf C$$^{*}$.
 by abuse of language we denote the lifted scalar product by the same symbol.
@@
 We are now in position to give a global definition of the anti-de Sitter plane waves.
+The same problem has been solved long ago in the de Sitter case [13, 14] (see Appendix B).
 Given any complex number $\lambda$ the natural domains of definition for plane waves in the AdS universe are ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$ and ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$:
 $\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}\ni\zeta,z\to\phi^{\pm}_{\lambda}(z,\zeta)=\left({z\cdot\zeta}\right)^{\lambda}=e^{\lambda\log\left({z\cdot\zeta}\right)}.$ (7.75)
-A previous attempt [18] to introduce global waves was based on a local $i\epsilon$-prescription.
+A previous attempt [32] to introduce global waves in AdS was based on a local $i\epsilon$-prescription.
 Here, in their global domains of definition, the waves are univalued, holomorphic solutions of the Klein-Gordon equation with complex squared mass $m^{2}_{\lambda}=\lambda(\lambda+d-1).$
@@
 8 Klein-Gordon fields
-Klein-Gordon fields display the simplest example of the analytic structure explained in Sect.
-5.
-Their two-point functions are distributions on $\widetilde{}AdS_{d}\times\widetilde{}AdS_{d}$ satisfying the AdS Klein-Gordon equation w.r.t.
-both variables
-$(\square+m^{2}_{\lambda}){\cal W}(x_{1},x_{2})=0,$ (8.71)
-where the complex squared mass is parametrized by a complex parameter $\lambda$ as in Eq.
-(6.50).
-For any such $\lambda$ there are two invariant maximally analytic solutions of the above equation, satisfying all the properties specified in Section 5, apart from the positive definiteness which may hold or fail:
-$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=$ $W^{(d)}_{\lambda}(z_{1}\cdot z_{2})={e^{-i\pi{d-2\over 2}}\over(2\pi)^{d\over 2}}((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}Q_{{d-2\over 2}+\lambda}^{d-2\over 2}(z_{1}\cdot z_{2})$ (8.72)
-$=$ $\frac{(z_{1}\cdot z_{2})^{1-d-\lambda}\Gamma(d-1+\lambda)\,_{2}F_{1}\left(\frac{d-1+\lambda}{2},\frac{d+\lambda}{2};\frac{d+1+2\lambda}{2});{(z_{1}\cdot z_{2})^{-2}}\right)}{2^{d+\lambda}\pi^{\frac{d-1}{2}}\Gamma\left(\frac{d+1}{2}+\lambda\right)}$ (8.73)
-$W^{(d)}_{1-d-\lambda}(z_{1},z_{2})$ $=$ $W^{(d)}_{1-d-\lambda}(z_{1}\cdot z_{2})={e^{-i\pi{d-2\over 2}}\over(2\pi)^{d\over 2}}((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}Q_{-{d\over 2}-\lambda}^{d-2\over 2}(z_{1}\cdot z_{2}).$ (8.76)
-In the above formulae Q{Q} denotes the associated Legendre function of the second kind; it is a meromorphic function of the indices [19]; as function of the complex variable $z=z_{1}\cdot z_{2}$, it is analytic in the cut-plane
-$\Delta_{1}=\{{\bf C}\setminus[-\infty,1]\}.$ (8.77)
-As explained in Section 4, the scalar product can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$ and, thanks to the factor $((z_{1}\cdot z_{2})^{2}-1)^{-\frac{d-2}{4}}$, Eqs.
-(8.72) and (8.76) can and must be understood in this sense.
-In view of [19, Eq.
-(3.3.1.4)] one has also that
-$W^{(d)}_{1-d-\lambda}(z)=W^{(d)}_{\lambda}(z)-\frac{\cos\left(\frac{1}{2}\pi(d+2\lambda)\right)\Gamma(-\lambda)\Gamma(d+\lambda-1)}{(2\pi)^{d\over 2}}(z^{2}-1)^{-\frac{d-2}{4}}P_{\frac{d-2}{2}+\lambda}^{-\frac{d-2}{2}}(z).$ (8.78)
-The second term at the rhs in this relation is regular on the cut $z\in[-1,1]$ and therefore does not contribute to the commutator; this means that the two solutions are equivalent at short distances.
-On the other hand, since it grows the faster the larger is $|\lambda+(d-2)/2|$ (see [19, Eqs.
-(3.9.2))], the two-solution have drastically different long range behaviors.
-The question arises whether and when they are both acceptable or not [9]. The positive definiteness condition gives the answer. We will come back on this point later.
-The Schwinger function (otherwise called the Euclidean propagator) is the restriction of the maximally analytic two-point function to the Euclidean Lobachevsky manifold extended to coincident points.
+Klein–Gordon fields provide the simplest example of the analytic structure described in section 5.
+Their two-point functions are distributions on $\widetilde{AdS_{d}}\times\widetilde{AdS_{d}}$ satisfying the AdS Klein–Gordon equation with respect to both variables:
+$\left(\square_{x_{1}}+m_{\lambda}^{2}\right)\mathcal{W}(x_{1},x_{2})=\left(\square_{x_{2}}+m_{\lambda}^{2}\right)\mathcal{W}(x_{1},x_{2})=0,$ (8.76)
+where the (complex) squared mass is parametrized by the complex number $\lambda$ as in eq. (6.51).
+For any such $\lambda$ there exist two invariant maximally analytic solutions satisfying all the properties listed in section 5, except possibly positive definiteness (unitarity).
+They are function of the complex variable $z=z_{1}\cdot z_{2}$ and are mapped into each other by the involution (6.54):
+$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=W^{(d)}_{\lambda}(z)=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{\frac{d-2}{2}+\lambda}(z)$ (8.77)
+$=\frac{z^{1-d-\lambda}\Gamma(d-1+\lambda)\,{}_{2}F_{1}\!\left(\frac{d-1+\lambda}{2},\frac{d+\lambda}{2};\frac{d+1+2\lambda}{2};z^{-2}\right)}{2^{d+\lambda}\pi^{(d-1)/2}\Gamma\!\left(\frac{d+1}{2}+\lambda\right)}$ (8.78)
+$W^{(d)}_{1-d-\lambda}(z_{1},z_{2})$ $=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{-\frac{d}{2}-\lambda}(z).$ (8.79)
+By using the parameter $\nu$ introduced in eq. (6.56), the above solutions can also be written as
+$W^{(d)}_{\nu(\lambda)}(z_{1},z_{2})$ $=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(z)$ (8.80)
+$=\frac{\Gamma\!\left(\frac{d-1}{2}+\nu\right)(2z)^{-\frac{d-1}{2}-\nu}\,{}_{2}F_{1}\!\left(\frac{d-1}{4}+\frac{\nu}{2},\frac{d+1}{4}+\frac{\nu}{2};\nu+1;\frac{1}{z^{2}}\right)}{2\pi^{\frac{d-1}{2}}\Gamma(\nu+1)},$ (8.81)
+$W^{(d)}_{-\nu(\lambda)}(z_{1},z_{2})$ $=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}\big(z^{2}-1\big)^{-\frac{d-2}{4}}Q^{\,\frac{d-2}{2}}_{-\frac{1}{2}-\nu}(z).$ (8.82)
+As expected, the involution acts simply by changing the sign of the parameter $\nu$.
+In the above formulae, and in what follows, $P^{\alpha}_{\beta}$ and $Q^{\alpha}_{\beta}$ denote the associated Legendre functions, which are meromorphic in their indices [18].
+As functions of the complex variable $z$, they are analytic in the cut plane
+$\Delta_{2}=\mathbf{C}\setminus[-\infty,1].$ (8.83)
+As explained in Section 4, the AdS scalar product $z_{1}\!\cdot\!z_{2}$ lifts to a map
+$\widetilde{\mathcal{Z}}_{-}\times\widetilde{\mathcal{Z}}_{+}\to\widetilde{\Delta}_{1},$
+(the covering of the cut plane $\Delta_{1}=\mathbb{C}\setminus[-1,1]).$ Equations (8.77)–(8.82) are to be understood in this sense.
+The Legendre function of the second kind $Q^{\,\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(z)$ (or equivalently $Q^{-\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(z)$), multiplied by the factor $(z-1)^{-\frac{d-2}{4}}(z+1)^{-\frac{d-2}{4}}$, are holomorphic in $\widetilde{\Delta}_{1}$, a domain naturally adapted to the anti-de Sitter geometry.
+Points on the cut [1,1][-1,1] of the fundamental sheet $\Delta_{1}$ correspond to pairs of timelike separated events on $AdS_{d}$}.
+In contrast, the following Legendre functions of the first kind [^3] In this case the sign of the upper index is uniquely fixed; these functions are also known as Gegenbauer functions [18]., multiplied by the same prefactor
+$(z-1)^{-\frac{d-2}{4}}(z+1)^{-\frac{d-2}{4}}P^{-\frac{d-2}{2}}_{-\frac{1}{2}+i\kappa}(z),\qquad\kappa\in\mathbb{C},$ (8.84)
+possess analyticity properties that are naturally adapted to the Lorentzian geometry of the de Sitter manifold [13, 14] (see Appendix B) as they are analytic in the cut plane
+$\Delta=\mathbf{C}\setminus(-\infty,-1].$ (8.85)
+The prefactor exactly compensates the singularity of the Legendre function and the product i regular there.
+Points lying on the cut $\infty,-1]$ correspond to pairs of timelike separated real events on the de Sitter manifold $dS_{d}$.
+The symmetry property of the Legendre functions of the first kind,
+$P^{\mu}_{-\frac{1}{2}-i\kappa}(z)=P^{\mu}_{-\frac{1}{2}+i\kappa}(z),\qquad\kappa\in\mathbb{C},\;z\in\Delta.$ (8.86)
+does not hold for the Legendre functions of the second kind.
+As a consequence, in the AdS case two distinct two-point functions can arise for a given mass, although they are not independent.
+Indeed, using [18, Eq.
+(3.3.1.4)], one obtains the identity
+$Q_{-\frac{1}{2}-\nu}^{\mu}(z)=Q_{-\frac{1}{2}+\nu}^{\mu}(z)+e^{i\pi\mu}\sin(\pi\nu)\Gamma\!\left(\tfrac{1}{2}+\mu-\nu\right)\Gamma\!\left(\tfrac{1}{2}+\mu+\nu\right)P_{-\frac{1}{2}+\nu}^{-\mu}(z).$ (8.87)
+This relation implies that
+$W^{(d)}_{-\nu}(z)=W^{(d)}_{\nu}(z)+\frac{\sin(\pi\nu)\Gamma\left(\frac{d-1}{2}+\nu\right)\Gamma\left(\frac{d-1}{2}-\nu\right)}{(2\pi)^{\frac{d}{2}}}(z^{2}-1)^{-\frac{d-2}{4}}P_{-\frac{1}{2}+\nu}^{-\frac{d-2}{2}}(z).$ (8.88)
+The second term on the right-hand side is regular on the AdS cut $z\in[-1,1]$ and therefore does not contribute to the commutator.
+Thus the two solutions coincide at short distances.
+However, since this term grows increasingly rapidly as $|\nu|$ becomes large (see [18, Eqs.
+(3.9.2)]), the two solutions exhibit markedly different long-distance behavior.
+The question of whether both solutions are physically admissible has been solved long ago: the answer is provided by the requirement of positive definiteness (unitarity) [9].
+An alternative viewpoint invoke the possible self-adjoint extensions of a certain Schrödinger operator [25, 26].
+The Schwinger function is the restriction of the maximally analytic two-point function to non-coincident Euclidean points:
+$S(z_{1}^{\scriptscriptstyle{E}},{z_{2}^{\scriptscriptstyle{E}})=\left.W(z_{1},z_{2})\right|_{EAdS_{d}\times EAdS_{d}\setminus\delta}}$ (8.89)
+where $\delta$ is the diagonal.
+The Euclidean propagator coincides with the two-point Schwinger function away from coincident points and is its distributional extension to the diagonal.
 Choosing the points in Eq.
@@
 Comparing this alternative representation of the two-point function with Eq.
-(9.92) we deduce a special case of a known (see [19, Eq.
+(9.103) we deduce a special case of a known (see [18, Eq.
 45, page 136]) and useful identity which we reproduce here for the reader’s convenience:
@@
 12 The Poincaré patch $AdS_{M}$
-In this section and the ones that follow, we perform a decomposition of the plane waves (7.70) and diagonalize the two-point function (1.1) in the Poincaré coordinate system.
-This approach paves the way for novel possibilities to compute Feynman diagrams directly in the real AdS space, as opposed to working in its Euclidean counterpart as is usually done.
-We denote by $AdS_{M}$ ($M$ stands for Minkowski) the region of the $d$-dimensional manifold $AdS_{d}$ (or, what is the same, of its covering $\widetilde{}AdS_{d}$) foliated by the family hyperplanes
+In this and the subsequent sections, we decompose the plane waves (7.75) and diagonalize the two-point function (1.1) using the Poincaré coordinate system.
+This framework opens up new opportunities to compute Feynman diagrams directly in Lorentzian AdS space, rather than relying on the conventional Euclidean approach.
+We denote by $AdS_{M}$ (with $M$ standing for Minkowski) the region of the $d$-dimensional manifold $AdS_{d}$} - or equivalently, of its universal covering space $\widetilde{AdS_{d}} -$ that is foliated by the family of hyperplanes
 $\Pi_{u}=\{x^{d-1}+x^{d}=1/u=e^{v}>0\}.$ (12.154)
-The following system parametrizes the leaves by Lorentzian coordinates
+The following system parametrizes the leaves by Lorentzian (Poincaré) inertial coordinates
 […]
@@
 ${\rm x}^{2}={\rm x}^{2}_{M}={\rm x}\cdot{\rm x}=\eta_{\mu\nu}{\rm x}^{\mu}{\rm x}^{\nu},\ \ \ \ ds_{M}^{2}={(d{\rm x}^{0})}^{2}-{(d{\rm x}^{1})}^{2}-\cdots-{(d{\rm x}^{d-2})}^{2}$ (12.156)
-are the Lorentzian scalar product and interval in a (d1)(d-1)-dimensional Minkowski space $M_{d-1}$.
-In the Poincaré coordinates the AdS scalar product and interval are written as follows:
+are the scalar product and the interval in a (d1)(d-1)-dimensional Minkowski space $M_{d-1}$. In Poincaré coordinates, the AdS scalar product and interval are expressed as follows:
 $(x-x^{\prime})_{AdS}^{2}=\frac{\left({\rm x}-{\rm x}^{\prime}\right)_{M}^{2}-({u^{2}-{u^{\prime}})^{2}}}{uu^{\prime}},\ \$(12.157)
@@
 $(x(u,{\rm x})-x(u,{\rm x}^{\prime}))_{AdS}^{2}=\frac{1}{u^{2}}({\rm x}-{\rm x}^{\prime})_{M}^{2}.$ (12.159)
-This equation tells us that the Lorentzian geometry of the leaves is coherent with that of the AdS ambient space i.e.
-the causal ordering in any hyperplane $\Pi_{u}$ can be understood equivalently either in the Minkowskian sense of $\Pi_{u}$ or in the sense of the ambient AdS universe.
-The manifold $AdS_{M}$ is strongly causal; the price to pay is that it is geodesically incomplete.
+This equation shows that the Lorentzian geometry of the leaves is consistent with that of the ambient AdS space; that is, the causal ordering within any hyperplane $\Pi_{u}$ can be understood either in the Minkowskian sense of $\Pi_{u}$ or in the context of the ambient AdS spacetime.
+The manifold $AdS_{M}$ is strongly causal, though this comes at the cost of geodesic incompleteness.
 The second crucial property is that the the complex leaves $\Pi_{u}^{(c)}$ contain copies of the Minkowsian tubes
@@
 which are (respectively) contained in the ambient AdS tubes $\cal Z$$_{\pm}$.
-This means the following: let us consider a partial complexification of $AdS_{M}$ where the Poincaré coordinates $\mathbf{x}^{\mu}$ in Eq.
+This can be stated as follows: consider a partial complexification of $AdS_{M}$ in which the Poincaré coordinates $\mathbf{x}^{\mu}$ in Eq.
 (12.155) are extended to complex values, while the coordinate $u$ remains real and positive:
 $z({\rm z},u)=\left\{\begin{tabular}[]{lcl}$z^{\mu}$&=&$\frac{1}{u}{\rm z}^{\mu}$\cr$z^{d-1}$&=&$\frac{1-u^{2}}{2u}+\frac{1}{2u}{\rm z}^{2}$ \ ,\cr$z^{d}$&=&$\frac{1+u^{2}}{2u}-\frac{1}{2u}{\rm z}^{2}$\end{tabular}\right.\ \ {\rm z}^{2}=\eta_{\mu\nu}{\rm z}^{\mu}{\rm z}^{\nu},\ \ \ u>0.$ (12.161)
-The leaves at fixed $u>0$ are copies of $M_{d-1}^{(c)}$. Suppose now that ${\rm z}={\rm x}+i{\rm y}$ and that ${\rm y}^{2}_{M}>0$, i.e. $\rm z$} belongs to either $T^{+}$ or $T^{-}$; it follows that
+The leaves at fixed $u>0$ are copies of $M_{d-1}^{(c)}$.
+Suppose now that ${\rm z}={\rm x}+i{\rm y}$ and that ${\rm y}^{2}_{M}>0$, i.e.
+$\rm z$} belongs to either $T^{+}$ or $T^{-}$; it follows immediately that
 $(\mathrm{Im}\,z({\rm z},u))\cdot(\mathrm{Im}\,z({\rm z},u))=\frac{{\rm y}^{2}_{M}}{u^{2}}>0;$ (12.162)
-Eq.
-(4.32) then implies that the Minkowskian tubes $T_{+}$ and $T_{-}$ of the leaves are contained into $\cal Z$$_{+}$ and, respectively $\cal Z$$_{-}$ (see [11] for more details).
-This property has also an important technical value; in particular it allows to compute the Fourier transforms of the waves w.r.t.
-the coordinates ${\rm x}^{\mu}$; we will do it in the next sections.
-The last very important geometrical fact is that the above partial complexification already contains the whole Euclidean manifold $EAdS_{d}$.
+equation (4.33) then implies that the Minkowskian tubes $T_{+}$ and $T_{-}$ of the leaves are contained within $\mathcal{Z}_{+}$ and, respectively, $\mathcal{Z}_{-}$ (see [10] for further details).
+This property also has important technical significance: in particular, it allows the computation of Fourier transforms of the waves with respect to the coordinates $\mathbf{x}^{\mu}$, which we will carry out in the following sections.
+Finally, a crucial geometrical observation is that the above partial complexification already encompasses the entire Euclidean manifold $EAdS_{d}$.
 In the following we will consider the submanifolds of the null cone defined by the condition
@@
 $\xi(\eta)=\left\{\begin{array}[]{lcc}\xi^{\mu}&=&{\eta^{\mu}}\\ \xi^{d-1}&=&\pm\frac{1}{2}(1+{\eta}^{2})\\ \xi^{d}&=&\pm\frac{1}{2}(1-{\eta}^{2})\end{array}\right.$. (12.164)
-The invariant measure on $\gamma_{P}^{(\pm)}$ is $d\mu(\xi(\eta))=d\eta$ and the scalar product $x\cdot\xi$} take the form
-$x(t,u)\cdot\xi(\eta)=\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u},\ \ \ \ \ \ \xi(\eta)\in\gamma_{P}^{(+)}$ (12.154)
+The invariant measure on $\gamma_{P}^{(\pm)}$ is $d\mu(\xi(\eta))=d\eta$ and the scalar product $x\cdot\xi$} takes the forms
+$x(t,u)\cdot\xi(\eta)=\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u},\ \ \ \ \ \ \xi(\eta)\in\gamma_{P}^{(+)},$ (12.165)
 $x(t,u)\cdot\xi(\chi)=-\frac{u}{2}+\frac{\left({{\rm x}}+{\chi}\right)^{2}}{2u},\ \ \ \ \xi(\chi)\in\gamma_{P}^{(-)}.$ (12.166)
@@
 In the special case $d=2$ it is enough to integrate either one of the red or the blue t cycles which are disconnected.
-The two-dimensional case is again simpler but represents a valuable training ground for solving the general case which is much harder.
-In spacetime dimension $d=2$ the leaves $\Pi_{u}$ are time-like parabolae parametrized just by one time coordinate (i.e.
-$x^{0}=t/u$, see Eq.
-(12.144)).
-$\gamma^{+}_{P}$ is also parametrized by only one variable $\eta^{0}=\eta$ dual to the time $t$.
+The two-dimensional case is simpler but still provides a valuable training ground for tackling the more general and challenging higher-dimensional case.
+In spacetime dimension $d=2$, the leaves $\Pi_{u}$ are time-like parabolae, each parametrized by a single time coordinate ${\rm x}^{0}=t$, i.e.,
+$x^{0}=\frac{t}{u}\,$, (13.175)
+(see Eq. 12.155). Similarly, $\gamma^{+}_{P}$ is parametrized by a single dual variable $\eta^{0}=\eta$, which is conjugate to the time $t$.
 Let us chose the first point $z_{1}=x_{1}$ in Eq.
@@
 $=\frac{1}{4\pi}\int_{t-u}^{t+u}\left(\frac{u^{2}-(t-\eta)^{2}}{2u}\right)^{\lambda}\left(\frac{{u^{\prime}}^{2}-(t^{\prime}+is^{\prime}-\eta)^{2}}{2u^{\prime}}\right)^{-\lambda-1}d\eta.$ (13.178)
-This is yet another new integral representation of the Legendre function of the second kind $Q_{\lambda}$; this representation is naturally associated with the Poincaré coordinate system.
-We now proceed to separate the variables in the above expression by taking the Fourier transform of the holomorphic plane waves (7.70) w.r.t. the time coordinate $t$:
-$\int(x(t-i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\theta(\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{{\Gamma(-\lambda)}}\,e^{i\omega\eta}\,{\omega^{-\frac{1}{2}-\lambda}}\,J_{-\frac{1}{2}-\lambda}\left(\omega u\right),$ (13.167)
-$\int(x(t+i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\theta(-\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{{\Gamma(-\lambda)}}\,e^{i\omega\eta}\,{|\omega|^{-\frac{1}{2}-\lambda}}\,J_{-\frac{1}{2}-\lambda}\left(|\omega|u\right),$ (13.168)
-valid for for $1<\mathrm{Re}\,\lambda<0$; $\theta(\omega)$ is Heaviside’s step function.
-The physically relevant feature of the above Fourier transforms is that planes waves holomorphic in the backward tube of the complex AdS manifold contain only positive frequencies w.r.t.
-the time coordinate of the Poincaré branes and planes waves holomorphic in the forward tube only contain negative frequencies.
-This support property is a consequence of the geometric inclusion of the tubes described in Eq.
+This provides yet another integral representation of the Legendre function of the second kind, $Q_{\lambda}$, which is naturally associated with the Poincaré coordinate system.
+We now proceed to separate variables in the above expression by taking the Fourier transform of the holomorphic plane waves (7.75) with respect to the time coordinate $t$. The result is
+$\int(x(t-i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\,\theta(\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{\Gamma(-\lambda)}\,e^{i\omega\eta}\,\omega^{-\frac{1}{2}-\lambda}\,J_{-\frac{1}{2}-\lambda}(\omega u),$ (13.179)
+$\int(x(t+i\epsilon,u)\cdot\xi(\eta))^{\lambda}e^{i\omega t}\,dt\,=\,\theta(-\omega)\,\frac{(2\pi)^{\frac{3}{2}}u^{\frac{1}{2}}}{\Gamma(-\lambda)}\,e^{i\omega\eta}\,|\omega|^{-\frac{1}{2}-\lambda}\,J_{-\frac{1}{2}-\lambda}(|\omega|u),$ (13.180)
+valid for $1<\mathrm{Re}\,\lambda<0$, where $\theta(\omega)$ is Heaviside’s step function.
+A physically relevant feature of the above Fourier transforms is that plane waves holomorphic in the backward tube of the complex AdS manifold contain only positive frequencies with respect to the time coordinate of the Poincaré branes, whereas plane waves holomorphic in the forward tube contain only negative frequencies.
+This support property follows from the geometric inclusion of the tubes described in Eq.
 (12.162).
-Inversion allows to rewrite the integral (13.166) as follows:
+Inversion then allows us to rewrite the integral (13.178) as follows:
 $\int\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}\left(\frac{u^{\prime}}{2}-\frac{\left({t^{\prime}+is^{\prime}}-{\eta}\right)^{2}}{2u}\right)^{-\lambda-1}d\eta=$ (13.181)
-$=\frac{(2\pi u^{\prime})^{1/2}}{\Gamma(\lambda+1)}\int\int_{-t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{-i\omega(t^{\prime}+is^{\prime}-\eta)}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}(|\omega|u^{\prime})\theta(-\omega)}d\omega d\eta$ (13.170)
-The integral over $\eta$ contains both positive and negative frequencies (no step function is present here); the result is
+$=\frac{(2\pi u^{\prime})^{1/2}}{\Gamma(\lambda+1)}\int\int_{-t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{-i\omega(t^{\prime}+is^{\prime}-\eta)}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}(|\omega|u^{\prime})\theta(-\omega)}d\omega d\eta.$ (13.182)
+The integral over $\eta$ contains both positive and negative frequencies (no step function is present here):
 $\int_{t-u}^{{t+u}}\left(\frac{u}{2}-\frac{\left(t-{\eta}\right)^{2}}{2u}\right)^{\lambda}e^{i\omega\eta}d\eta=\sqrt{2\pi u}|\omega|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)J_{\frac{1}{2}+\lambda}(|\omega|u),\ \ \ \mathrm{Re}\,\lambda>-1.$ (13.184)
-Putting everything together we arrive at the final result:
-$W^{(2)}_{\lambda}(x_{1}(t,u),z_{2}(t^{\prime}+is^{\prime},u^{\prime}))=\frac{1}{2}{(uu^{\prime})^{1/2}}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}J_{\frac{1}{2}+\lambda}(\omega u)J_{\frac{1}{2}+\lambda}(\omega u^{\prime})d\omega.$ (13.173)
-This expression has been proven under the condition $1<\mathrm{Re}\,\lambda<0$ but is valid more generally in the sense of distributions.
-It provides a diagonalization of the two-point function which, for real $\lambda$, manifestly exhibits its positive-definiteness where it converges.
-By comparing this expression with the one obtained by inserting the Fourier representation of both the first and the second factor in Eq.
-(13.166) we get as a bonus a time-honored formula of Sonine which we do not reproduce here [21].
+Putting everything together, we arrive at the final result in the form of a Hankel representation:
+$W^{(2)}_{\lambda}\big(x_{1}(t,u),z_{2}(t^{\prime}+is^{\prime},u^{\prime})\big)$ $=$ $\frac{1}{2}\,(uu^{\prime})^{1/2}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}\,J_{\frac{1}{2}+\lambda}(\omega u)\,J_{\frac{1}{2}+\lambda}(\omega u^{\prime})\,d\omega,$ (13.185)
+which has been proven under the condition $1<\mathrm{Re}\,\lambda<0$, but is valid more generally in the sense of distributions.
+It provides a diagonalization of the two-point function which, for real $\lambda$, manifestly exhibits its positive-definiteness in the region where it converges.
+By comparing this expression with the one obtained by inserting the Fourier representation of both the first and second factors in Eq.
+(13.178), we also recover a classical formula of Sonine, which we do not reproduce here [19].
 13.1 The other cycle
-In preparation for the general case, it is instructive to integrate also over the other non-compact parabolic cycle.
-(see Figure 3).
-Even though we expect a result identical to (13.174) the way to reach it is rather different as the integration cycle is now unbounded and touches infinity.
-The delicate point is the behavior of the waves at infinity and the way to circumvent such a difficulty will give crucial indications on how to deal with the problem in general spacetime dimension $d$.
-The relevant cycle is
+In preparation for the general case, it is instructive to integrate also over the other non-compact parabolic cycle (see Figure 3).
+Even though we expect a result identical to (LABEL:24), the method to reach it is rather different, as the integration cycle is now unbounded and extends to infinity.
+The delicate point lies in the behavior of the waves at infinity, and the strategy to circumvent this difficulty will provide crucial guidance for handling the problem in general spacetime dimension $d$.
+The relevant relative cycle is
 $\gamma_{P}^{-}(x_{1})=\left\{\xi\in\gamma_{P}^{-}:\ \ x_{1}\cdot\xi(\chi)=-\frac{u}{2}+\frac{\left({{\rm x}}+{\chi}\right)^{2}}{2u}\geq 0\right\}.$ (13.187)
@@
 $\lim_{\chi\to\pm\infty}\text{Arg}\left(\frac{(t^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)=0.$ (13.189)
-This condition has to be interpreted as follows: in formula (13.176) there are two disconnected semi-infinite intervals of integration; the phase of the second factor of the integrand has to be chosen with two different determinations.
-Eq.
-(13.176) actually means
+This condition is to be interpreted as follows.
+In formula (13.188), there are two disconnected semi-infinite intervals of integration; the phase of the second factor of the integrand must be chosen with two different determinations, one for each connected part of the integration cycle.
+Equation (13.188) thus actually means
 $W^{(2)}_{\lambda}(z_{1},z_{2})=\frac{1}{4\pi}\int_{-\infty}^{-t-u}\left(\frac{(t+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{(t^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)_{-}^{-\lambda-1}d\chi$ (13.190)
@@
 $W^{(2)}_{\lambda}(z_{1},z_{2})=$ (13.196)
-$=$ $\frac{\Gamma(-\lambda)\sin(\pi\lambda)\sqrt{u^{\prime}}}{4\pi\sqrt{2\pi}}\int_{-\infty}^{0}e^{i\omega(t-t^{\prime}-is^{\prime})}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)}{}d\omega\int_{u}^{\infty}\cos(\omega\chi-\pi\lambda)\left(\frac{\eta^{2}-u^{2}}{2u}\right)^{\lambda}d\chi$ (13.186)
+$=$ $\frac{\Gamma(-\lambda)\sin(\pi\lambda)\sqrt{u^{\prime}}}{4\pi\sqrt{2\pi}}\int_{-\infty}^{0}e^{i\omega(t-t^{\prime}-is^{\prime})}{|\omega|^{\frac{1}{2}+\lambda}J_{\frac{1}{2}+\lambda}\left(|\omega|u^{\prime}\right)}{}d\omega\times$ (13.199)
+${}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\int_{u}^{\infty}\cos(\omega\chi-\pi\lambda)\left(\frac{\eta^{2}-u^{2}}{2u}\right)^{\lambda}d\chi=$
 $=$ $\frac{1}{2}{(uu^{\prime})^{1/2}}\int_{0}^{\infty}e^{-i\omega(t-t^{\prime}-is^{\prime})}J_{\frac{1}{2}+\lambda}(\omega u)J_{\frac{1}{2}+\lambda}(\omega u^{\prime})d\omega.$ (13.200)
-As anticipated, we recover Eq.
-(13.174), though by means of a significantly more involved procedure.
-In the case of general dimensions, the two separate integration cycles contribute differently.
-Therefore, it is essential to account for both contributions to obtain the complete result.
+As anticipated, we recover the same Hankel representation (LABEL:24), although by a significantly more involved procedure.
+In general spacetime dimension $d>2$, the two separate integration cycles contribute differently.
+Therefore, it is essential to account for both contributions in order to obtain the complete result.
 14 Waves and two-point functions in Poincaré coordinates: general case
-Let us now face the problem of diagonalizing the two-point function (1.1) in Poincaré coordinates in arbitrary spacetime dimension $d$.
-To this end let us choose $x_{1}=x_{1}({\rm x},u)$ real and $z_{2}=z_{2}({\rm z}^{\prime},u^{\prime})\in{\cal Z}_{+}$ with ${\rm z}^{\prime}\in T_{+}$ in the relevant complex Minkowskian brane.
-The integration cycle consists of two disconnected parts, denoted as $\gamma_{P}(x_{1})={\gamma^{+}_{P}(x_{1})\,\cup\,\gamma^{-}_{P}(x_{1})}$ where
-$\gamma^{+}_{P}(x_{1})=\left\{\xi(\eta)\in\gamma_{P}^{+}:\left(\eta^{0}-{\rm x}^{0}\right)^{2}\leq\left(\vec{\eta}-\vec{\rm x}\right)^{2}+u^{2}\right\},$ (14.188)
-$\gamma^{-}_{P}(x_{1})=\left\{\xi(\chi)\in\gamma_{P}^{-}:\left(\chi^{0}+{\rm x}^{0}\right)^{2}\geq\left(\vec{\chi}+\vec{\rm x}\right)^{2}+u^{2}\right\}.$ (14.189)
-As emphasized previously, unlike the two-dimensional case, it is essential to integrate over both disconnected cycles. The following expression is the one that requires evaluation.
+Let us now face the problem of diagonalizing the two-point function (1.1) in Poincaré coordinates in arbitrary spacetime dimension $d$. To this end, we choose
+$x_{1}=x_{1}({\rm x},u)\in AdS_{M}\quad\text{and}\quad z_{2}=z_{2}({\rm z}^{\prime},u^{\prime})\in{\cal Z}_{+},\quad{\rm z}^{\prime}\in T_{+},$
+in the relevant complex Minkowskian brane. The integration cycle consists of two disconnected parts, denoted as
+$\gamma_{P}(x_{1})=\gamma_{P}^{+}(x_{1})\,\cup\,\gamma_{P}^{-}(x_{1}),$
+where
+$\gamma_{P}^{+}(x_{1})$ $=$ $\left\{\xi(\eta)\in\gamma_{P}^{+}:(\eta^{0}-{\rm x}^{0})^{2}\leq(\vec{\eta}-\vec{{\rm x}})^{2}+u^{2}\right\},$ (14.201)
+$\gamma_{P}^{-}(x_{1})$ $=$ $\left\{\xi(\chi)\in\gamma_{P}^{-}:(\chi^{0}+{\rm x}^{0})^{2}\geq(\vec{\chi}+\vec{{\rm x}})^{2}+u^{2}\right\}.$ (14.202)
+As emphasized previously, unlike the two-dimensional case, it is essential to integrate over both disconnected cycles. The following expression is the one that requires evaluation:
 $W^{(d)}_{\lambda}(x_{1}({\rm x},u),z_{2}({\rm z}^{\prime},u^{\prime}))=$ (14.203)
@@
 $+c_{d}(\lambda)\int d\vec{\chi}\int^{\infty}_{-{\rm x}^{0}+\sqrt{\left(\vec{\chi}+\vec{\rm x}\right)^{2}+u^{2}}}\left(\frac{({\rm x}+\chi)^{2}-u^{2}}{2u}\right)^{\lambda}\left(\frac{({\rm z}^{\prime}+\chi)^{2}-{u^{\prime}}^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}d\chi^{0}$ (14.207)
-where the first line is the integral over $\gamma^{+}_{P}(x_{1})$} and the second two lines are the integral over $\gamma^{-}_{P}(x_{1})$}.
-The challenging task of evaluating the three integrals above will be rewarded by an exceptionally simple final result.
-Additionally, the calculation will yield new integral formulae involving Bessel functions.
-Let us start with the integral over $\gamma^{+}_{P}(x_{1})$}; in this case the relevant Fourier transform of the waves is as follows:
-$\widetilde{\phi^{\pm}_{\lambda}}({\rm k})=\int(x(t\pm i\epsilon,\vec{\rm x},u)\cdot\xi(\eta))^{\lambda}e^{i{\rm k}{\rm x}}\,d{\rm x}\,=e^{i{\rm k}\eta}\int\left[\frac{u}{2}-\frac{({\rm y}^{0}\pm i\epsilon)^{2}-\vec{\rm y}^{2}}{2u}\right]^{\lambda}e^{i{\rm k}{\rm y}}d{\rm y}$ (14.195)
-where ${\rm y}={\rm x}-\eta$ (and we left the dependence on $u$ of the lhs implicit). Let us chose ${\rm k}=({\rm k}^{0},\vec{0})$ and perform the radial integration first:
-$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)=e^{i{\rm k}^{0}\eta^{0}}\frac{2\pi^{\frac{d-2}{2}}}{\Gamma(\frac{d-2}{2})}\int_{-\infty}^{\infty}d{\rm y}^{0}\int_{0}^{\infty}\left[\frac{u}{2}-\frac{({\rm y}^{0}+i\epsilon)^{2}-r^{2}}{2u}\right]^{\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}r^{d-3}\,dr=$ (14.196)
+Here, the first line corresponds to the integral over $\gamma_{P}^{+}(x_{1})$, while the second and third lines correspond to the integral over $\gamma_{P}^{-}(x_{1})$.
+The challenging task of evaluating these three integrals will be rewarded by an exceptionally simple final result.
+Moreover, the calculation will yield new integral formulae involving Bessel functions.
+Let us start with the integral over $\gamma_{P}^{+}(x_{1})$. In this case, the relevant Fourier transform of the waves is given by
+$\widetilde{\phi^{\pm}_{\lambda}}({\rm k})$ $=$ $\int\big(x({\rm x}^{0}\pm i\epsilon,\vec{{\rm x}},u)\cdot\xi(\eta)\big)^{\lambda}\,e^{i{\rm k}\cdot{\rm x}}\,d{\rm x}$ (14.208)
+$=$ $e^{i{\rm k}\cdot\eta}\int\left[\frac{u}{2}-\frac{({\rm y}^{0}\pm i\epsilon)^{2}-\vec{{\rm y}}^{2}}{2u}\right]^{\lambda}e^{i{\rm k}\cdot{\rm y}}\,d{\rm y},$ (14.209)
+where ${\rm y}={\rm x}-\eta$, and we leave the dependence on $u$ of the left-hand side implicit.
+Let us choose ${\rm k}=({\rm k}^{0},{0})$ and perform the radial integration first:
+$\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)$ $=$ $e^{i{\rm k}^{0}\eta^{0}}\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{-\infty}^{\infty}d{\rm y}^{0}\int_{0}^{\infty}\left[\frac{u}{2}-\frac{({\rm y}^{0}+i\epsilon)^{2}-r^{2}}{2u}\right]^{\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}r^{d-3}\,dr$ (14.210)
 $=$ $\frac{(2u)^{-\lambda}\,e^{i{\rm k}^{0}\eta^{0}}\,\pi^{\frac{d-2}{2}}\,\Gamma\left(\frac{2-d}{2}-\lambda\right)}{\Gamma(-\lambda)}\int_{-\infty}^{\infty}\left(u^{2}-({\rm y}^{0}+i\epsilon)^{2}\right)^{\frac{d-2}{2}+\lambda}e^{i{\rm k}^{0}{\rm y}^{0}}\,d{\rm y}^{0}$ (14.211)
-valid for $\mathrm{Re}\,(2\lambda+d)<2$ and $\mathrm{Re}\,d>2$ (these restrictions will be removed at the end when the integral will be considered in the sense of distributions).
-When ${\rm k}^{0}>0$ the contour of integration may be pushed to infinity and the integral is shown to be zero.
-The geometric and physical reason behind this possibility is once more the already mentioned immersion of the Minkowskian tubes of the branes in the AdS chiral tuboids.
+which is valid for $\mathrm{Re}\,(2\lambda+d)<2$ and $\mathrm{Re}\,d>2$. These restrictions will be removed at the end when the integral will be considered in the sense of distributions.
+When ${\rm k}^{0}>0$, the contour of integration can be pushed to infinity, and the integral is then seen to vanish.
+The geometric and physical reason behind this possibility is, once again, the previously mentioned immersion of the Minkowskian tubes of the branes into the AdS chiral tuboids.
 When ${\rm k}^{0}<0$ the above contour distortion is not possible and we get
@@
 $\widetilde{\phi^{+}_{\lambda}}({\rm k}^{0},0)=\frac{4\sqrt{2}|{\rm k}^{0}|^{-\frac{1}{2}-\lambda}e^{i\eta{\rm k}^{0}}\pi^{\frac{d+1}{2}}u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)\Gamma(-\lambda)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right).$ (14.213)
-By comparing Eqs. (14.198) and (14.199) we deduce the identity
-$\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)=2^{\frac{d-4}{2}}\Gamma\left(\frac{d-2}{2}\right)|{\rm k}^{0}|^{1-\frac{d}{2}}u^{\frac{d-1}{2}+\lambda}J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u)$ (14.200)
-valid when the integral converges. This identity is special instance of a known integral formula by Watson [21] and will be used later.
-By restoring Lorentz invariance we obtain the Fourier integral representation of the wave (which we write in a way ready to be used in our calculation):
+By comparing Eqs. (14.212) and (14.213), we deduce the identity
+$\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{1}{4}+\frac{\lambda}{2}}J_{-\frac{1}{2}-\lambda}\Big(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\Big)\,dr=2^{\frac{d-4}{2}}\Gamma\Big(\frac{d-2}{2}\Big)|{\rm k}^{0}|^{\frac{2-d}{2}}u^{\frac{d-1}{2}+\lambda}J_{-\frac{d-1}{2}-\lambda}(|{\rm k}^{0}|u),$ (14.214)
+valid when the integral converges. This identity is a special case of a known integral formula by Watson [19] and will be used later.
+Restoring Lorentz invariance, we obtain the Fourier integral representation of the wave (written in a form ready to be used in our calculation):
 $(x({\rm z}^{\prime},u^{\prime})\cdot\xi(\eta))^{-\lambda-d+1}=\left(\frac{{u^{\prime}}^{2}-({\rm z}^{\prime}-\vec{\eta})^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}=$ (14.215)
 $=\frac{(2\pi)^{\frac{3-d}{2}}{u^{\prime}}^{\frac{d-1}{2}}}{\Gamma(\lambda+d-1)}\int e^{-i{\rm k}({\rm z}^{\prime}-\eta)}\theta(-{\rm k}^{0})\theta({\rm k}^{2})\left(\sqrt{{\rm k}^{2}}\right)^{\frac{d-1}{2}+\lambda}J_{\frac{d-1}{2}+\lambda}(\sqrt{{\rm k}^{2}}\,u^{\prime})d{\rm k},$ (14.216)
-where the factor $\theta(-{\rm k}^{0})$ guarantees the integral convergence as long as ${\rm z}^{\prime}\in T_{+}$.
-We are now ready to evaluate the first line (14.194) (i.e.
-the integral over the cycle $\gamma^{(+)}_{P}(x_{1})$).
-Without loss of generality we may simplify the calculation by choosing ${\rm x}=0$ i.e.
-$x_{1}=x_{1}(0,u).$ Proceeding as in the two-dimensional case (see Eq.
-(13.172)) we are led to examine the integral
-$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}dr\int_{-\sqrt{r^{2}+u^{2}}}^{\sqrt{r^{2}+u^{2}}}e^{i{\rm k}^{0}\eta^{0}}\left(\frac{u^{2}+r^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}d\eta^{0}=$ (14.204)
-$=\frac{2\sqrt{2}\pi^{\frac{d-1}{2}}|{\rm k}^{0}|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)dr$ (14.205)
-where we set $|\vec{\eta}|=r$ and integrated over the angles; we have assumed $\mathrm{Re}\,\lambda>-1$.
-Though similar to (14.199) the integral on the rhs of (14.206) seems to be unknown.
-Following the example of Eq.
-(14.199), we may evaluate it by integrating first the variable $r$ on the lhs of (14.206):
+where the factor $\theta(-{\rm k}^{0})$ guarantees convergence of the integral as long as ${\rm z}^{\prime}\in T_{+}$.
+We are now ready to evaluate the first line of Eq.
+(14.207) (i.e., the integral over the cycle $\gamma_{P}^{+}(x_{1})$).
+Without loss of generality, we may simplify the calculation by choosing ${\rm x}=0$, i.e.,
+$x_{1}=x_{1}(0,u).$
+Proceeding as in the two-dimensional case (see Eq. (13.184)), we are led to examine the integral
+$\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\,dr\int_{-\sqrt{r^{2}+u^{2}}}^{\sqrt{r^{2}+u^{2}}}e^{i{\rm k}^{0}\eta^{0}}\left(\frac{u^{2}+r^{2}-(\eta^{0})^{2}}{2u}\right)^{\lambda}d\eta^{0}=$
+$=\frac{2\sqrt{2}\pi^{\frac{d-1}{2}}|{\rm k}^{0}|^{-\lambda-\frac{1}{2}}\Gamma(\lambda+1)u^{-\lambda}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}\left(r^{2}+u^{2}\right)^{\frac{\lambda}{2}+\frac{1}{4}}J_{\lambda+\frac{1}{2}}\left(|{\rm k}^{0}|\sqrt{r^{2}+u^{2}}\right)dr.$ (14.218)
+Here we set $|\vec{\eta}|=r$ and integrate over the angular variables, assuming that $\mathrm{Re}\,\lambda>-1$.
+Although the integral on the right-hand side of (14.219) is similar to that in (14.213), it does not appear to be directly known.
+Following the procedure used in Eq.
+(14.213), we evaluate it by first integrating over the variable $r$ on the left-hand side of (14.219):
 $\frac{2\pi^{\frac{d-2}{2}}}{\Gamma\left(\frac{d-2}{2}\right)}\int_{0}^{\infty}r^{d-3}dr\int_{-\sqrt{r^{2}+u^{2}}}^{\sqrt{r^{2}+u^{2}}}e^{i{\rm k}^{0}\eta^{0}}\left(\frac{u^{2}+r^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}d\eta^{0}=$ (14.220)
@@
 $\times\theta({\rm k}^{2})\left(\cos(\pi\lambda)J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)-\sin\left(\frac{\pi d}{2}\right)J_{-\frac{d-1}{2}-\lambda}\left(u\sqrt{{\rm k}^{2}}\right)\right).$ (14.227)
-By inserting Eqs. (14.203) and (14.214) into Eq. (14.194) we get the first part of the result computed in the special case when $x_{1}=x_{1}(0,u)$:
+By substituting Eqs. (14.217) and (14.227) into Eq. (14.207), we obtain the first part of the result, computed for the special case in which $x_{1}=x_{1}(0,u)$:
 $\int\int_{-\sqrt{u^{2}+\vec{\eta}^{2}}}^{\sqrt{u^{2}+\vec{\eta}^{2}}}\left(\frac{u^{2}+\vec{\eta}^{2}-{\eta^{0}}^{2}}{2u}\right)^{\lambda}\left(\frac{{u^{\prime}}^{2}({\rm z}^{\prime}-\eta)^{2}}{2u^{\prime}}\right)^{-\lambda-d+1}d\eta^{0}d\eta=$ (14.228)
@@
 Let us now focus on the other cycle $\gamma^{(-)}_{P}(x_{1})$.
-Here the situation is more complicated because the cycle goes through infinity and, as in Eq.
-(13.179), we need the Fourier transforms of the waves whose phase is vanishing at infinity.
-The analogous of Eq.
+In this case the situation is more complicated because the cycle passes through infinity and, as in Eq.
+(13.191), we need the Fourier transforms of the waves whose phase vanishes at infinity.
+The analogue of Eq.
 (13.189) is here
@@
 $\int_{\gamma_{P}(x_{1})}(x_{1}\cdot\xi)^{\lambda}(\xi\cdot z_{2})^{-\lambda-d+1}d\mu_{\gamma}(\xi)=$ (14.247)
-$=$ $-\frac{4\pi\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{\Gamma(d+\lambda-1)}\int e^{-i{\rm k}({\rm x}-x^{\prime})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}(u\sqrt{k^{2}})J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{k^{2}}\right)$ (14.235)
+$=$ $-\frac{4\pi\cos\left(\frac{\pi d}{2}\right)\Gamma(\lambda+1)\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{\Gamma(d+\lambda-1)}\int e^{-i{\rm k}({\rm x}-{\rm x}^{\prime})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}(u\sqrt{k^{2}})J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{k^{2}}\right)$ (14.248)
 Finally we restore the correct normalization of the two-point function by multiplying both sides of the above formula by $c_{d}(\lambda)$ and get
@@
 $u^{d}{\partial_{u}}(u^{2-d}\partial_{u}\left(u^{\frac{d-1}{2}}J_{\frac{d-1}{2}+\lambda}\left(mu\right)\right)=u^{\frac{d-1}{2}}\left(\lambda(d+\lambda-1)-m^{2}u^{2}\right)J_{\frac{d-1}{2}+\lambda}(mu).$ (14.255)
-The remarkable formula (14.240) shows that the maximally analytic AdS two-point function can be represented as a Källén-Lehmann superposition of maximally analytic two-point functions defined in a Minkowski space with one dimension less, explicitly linking the analytic and geometric structures of AdS and Minkowski spacetimes.
+The remarkable Hankel formula (14.253) shows that the maximally analytic AdS two-point function can be represented as a Källén-Lehmann superposition of maximally analytic two-point functions defined in a Minkowski space with one dimension less, explicitly linking the analytic and geometric structures of AdS and Minkowski spacetimes.
 Eq.
-(14.240) was anticipated in [23, 24] by employing Hilbert space methods suitable for generic warped manifolds; the value of the derivation in [23, 24] was also to clarify the Breitenlohner and Friedmann phenomenon [9] in terms of the self-adjointness of a one-dimensional transverse Schrodinger operator; however, a crucial part of the proof was based on a known Hankel transform of the Legendre function $Q$.
-The present derivation is fully self-contained and specifically tailored to the AdS geometry.
+(14.253) was anticipated in [27] and rediscovered by [29, 30].
+In [25, 26] it was derived by employing Hilbert space methods suitable for generic warped manifolds; the value of the derivation in [25, 26] was also to clarify the Breitenlohner and Friedmann phenomenon [9] in terms of the self-adjointness of a one-dimensional transverse Schrodinger operator; however, a crucial part of the proof was based on a known Hankel transform of the Legendre function $Q$ [17].
+The present derivation is fully self-contained and specifically tailored to the geometry of the complex AdS manifold and the AdS spectral condition.
 Eq.
@@
 15 Green functions, Schwinger functions and Wick rotations
-Eq. (14.240) may be used to promptly build a new explicit integral of the Schwinger function (8.80) extended to non coincident Euclidean points of $EAdS_{d}$:
+Eq. (14.253) may be used to promptly build a new explicit Hankel representation of the Schwinger function (8.91) extended to non coincident Euclidean points of $EAdS_{d}$:
 $S^{(d)}_{\lambda}(z_{\scriptscriptstyle{E}},z^{\prime}_{\scriptscriptstyle{E}})=\frac{1}{2}\left(uu^{\prime}\right)^{\frac{d-1}{2}}\int S^{E_{d-1}}_{m}({\rm x}_{\scriptscriptstyle{E}},{\rm x}^{\prime}_{\scriptscriptstyle{E}})J_{\frac{d-1}{2}+\lambda}\left(mu\right)J_{\frac{d-1}{2}+\lambda}\left(mu^{\prime}\right)dm^{2},$ (15.262)
@@
 There exists a commonly used integral representation of the AdS Schwinger functions called “split representation” [34, 35, 36, 37]; it may be obtained by first representing the Legendre function of the second kind $Q$ as a Mehler-Fock transform of the corresponding Legendre function of the first kind $P$ [18, 38] and then expanding $P$ in plane waves on the Lobachevsky manifold [14, 39].
-The advantage of the representation (15.249) is that it has admits a counterpart on the reals that may be obtained by Wick rotation while the split representation cannot be Wick-rotated.
+The advantage of the representation (15.262) is that it has admits a counterpart on the reals that may be obtained by Wick rotation while the split representation cannot be Wick-rotated (see appendix B).
 Here is the corresponding representation of the Feynman propagator of the Klein-Gordon operator $\Box+m^{2}_{\lambda}$ in the real manifold $AdS_{M}$:
@@
 $\widetilde{\cal C}_{\pm}$ are given by the same formulae where the real variable $t$ has been unfolded.
-Acknowledgments.
-I want to express my gratitude to Sergio Cacciatori for numerous suggestions.
-I would also like to thank the Institut des Hautes Etudes Scientifiques (Bures sur Yvette) and the Physics Department of Fudan University for their warm hospitality and support while writing this paper.
+Appendix B Remarks on the representations of the Euclidean propagator
+In section 8, we showed that certain specific Legendre functions of the second kind, multiplied by $(z^{2}-1)^{-\frac{d-2}{4}}$, possess analyticity properties ideally suited to the geometry of the anti-de Sitter manifold and its universal covering.
+In particular, the maximal analyticity of the two-point function (8.77) in the covering cut-plane $\widetilde{\Delta}_{1}$ arises from AdS invariance, local commutativity, and the positivity of the energy spectrum, while the location of the cut reflects causal structure.
+By contrast, as recalled in section 8, certain Legendre functions of the first kind, multiplied by the same factor, exhibit analyticity properties naturally adapted to the de Sitter geometry and are analytic in the cut-plane $\Delta$.
+The above analytic structures may be used to compare the status of the so-called ”split representation” of the AdS Euclidean propagator with its Hankel form (15.262).
+To this end, it is necessary to recall some properties of the plane-wave expansion [13, 14] of the (Bunch-Davies) de Sitter Wightman function, keeping the details to a strict minimum.
+The complex de Sitter universe may be represented as the manifold
+$dS_{d}^{(c)}=\{z\in M_{d+1}^{(c)}:\langle z,z\rangle=-1\},$ (B.290)
+(we set the radius $R=1$) immersed in a $(d+1)$-dimensional complex Minkowski spacetime.
+The scalar product [^5] We use the notation $\langle z_{1},z_{2}\rangle$ for the dS scalar product.
+This is distinct from the AdS scalar product (2.6): $z_{1}\cdot z_{2}=z_{1}^{0}z_{2}^{0}-z_{1}^{1}z_{2}^{1}-\dots-z_{1}^{d-1}z_{2}^{d-1}+z_{1}^{d}z_{2}^{d}.$ The real manifolds $dS_{d}$ and $AdS_{d}$ are real submanifolds of the same complex sphere.
+of two events is understood in the ambient Minkowski spacetime sense:
+$\zeta=\langle z_{1},z_{2}\rangle=z_{1}^{0}z_{2}^{0}-z_{1}^{1}z_{2}^{1}-\dots-z_{1}^{d}z_{2}^{d}.$ (B.291)
+As complex manifolds, $dS^{(c)}_{d}$ and $AdS_{d}^{(c)}$ coincide and are equivalent to the complex $d$-dimensional sphere $S_{d}^{(c)}$:
+$dS_{d}^{(c)}\simeq AdS_{d}^{(c)}\simeq S_{d}^{(c)}.$
+The real de Sitter universe is identified with the one-sheeted hyperboloid
+$dS_{d}=\{x\in M_{d+1}:\langle x,x\rangle=-1\}.$ (B.292)
+Two real events $x_{1}$ and $x_{2}$ are timelike separated if and only if
+$(x_{1}-x_{2})^{2}=-2-2\,x_{1}\cdot x_{2}>0.$
+The forward lightcone
+$C_{+}=\partial V_{+}=\{\xi\in M_{d+1}:\xi^{2}=\langle\xi,\xi\rangle=0,\ \xi^{0}>0\},$ (B.293)
+which describes the boundary at timelike infinity of the real de Sitter manifold, may be interpreted as the space of momentum directions of free particles [31].
+It is called “the absolute” in Gel’fand’s classic book [42], and has been curiously renamed the “celestial sphere” in more recent literature.
+The complex de Sitter space-time contains de Sitter-invariant tubular domains of analyticity [14], whose topology is much simpler than that of the AdS chiral tubes introduced in section 4.
+They can be described simply as the intersections of $dS_{d}^{(c)}$ with the tubes $T_{\pm}$ of the ambient complex Minkowski space $M_{d+1}^{(c)}$:
+${\cal T}_{\pm}=dS_{d}^{(c)}\cap T_{\pm}=\{x+iy\in dS_{d}^{(c)}:y\in\pm V_{+}\}.$ (B.294)
+The tubes $\mathcal{T}_{-}$ and $\mathcal{T}_{+}$ are bounded by two Riemannian (rather than Euclidean) manifolds, $\mathcal{H}_{-}$ and $\mathcal{H}_{+}$, whose points are purely imaginary:
+$\mathcal{H}_{\pm}=\{x+iy\in dS_{d}^{(c)}:x=0,\;y^{2}=\langle y,y\rangle=1,\;\mathrm{sign}\,y^{0}=\pm\}.$ (B.295)
+No analogous structure appears in the Minkowski or AdS tubes, which lack such boundaries away from the real boundary.
+Both $\mathcal{H}_{-}$ and $\mathcal{H}_{+}$ are equivalent to the $d$-dimensional Lobachevsky space $H_{d}$, i.e.
+the Euclidean AdS manifold.
+The simple topology of the tubes is reflected in an equally simple definition of global plane waves, in contrast with the complexity of the corresponding AdS holomorphic waves described in the present paper.
+Given any complex number $\lambda$, plane waves in the dS universe are constructed as follows [13, 14]:
+${\cal T}_{\pm}\times C_{+}\ni(z,\xi)\mapsto\psi^{\pm}_{\lambda}(z,\xi)=\langle\xi\,,z\rangle^{\lambda}.$ (B.296)
+They are univalued and holomorphic in the tubular domains $\cal T$$_{\pm}$ and solve there the Klein-Gordon equation
+$(\Box_{dS}+\mu_{\lambda}^{2})\,\psi^{\pm}_{\lambda}=0$ (B.297)
+with complex squared mass
+$\mu_{\lambda}^{2}=\lambda(1-d-\lambda)=\frac{(d-1)^{2}}{4}+\kappa^{2},\qquad\lambda=-\frac{d-1}{2}+i\kappa.$ (B.298)
+A plane-wave expansion of the dS two-point function in terms of the above plane waves has been available for a long time [13, 14].
+For any pair of events $z_{1},z_{2}$ in the domain ${\cal T}_{-}\times{\cal T}_{+}$, the canonically normalized (so-called) Bunch-Davies de Sitter two-point function is a holomorphic superposition of plane waves, precisely as in (1.2) and (1.3):
+$W^{(dS)}_{\kappa}(z_{1},z_{2})$ $=\frac{\Gamma\Big(\frac{d-1}{2}+i\kappa\Big)\Gamma\Big(\frac{d-1}{2}-i\kappa\Big)e^{\pi\kappa}}{2^{d+1}\pi^{d}}\int_{\gamma}\langle\xi\,,z_{1}\rangle^{-\frac{d-1}{2}-i\kappa}\langle\xi\,,z_{2}\rangle^{-\frac{d-1}{2}+i\kappa}\,d\mu(\xi),$ (B.299)
+where $d\mu_{\gamma}(\xi)$ is the restriction of the volume form of the cone $C_{+}$ to a complete (d1)(d-1)-dimensional cycle [^6] For example one may chose the spherical section of the cone canonically oriented: $\gamma_{0}=C_{+}\cap\{\xi\ :\ \xi^{0}=1\}=\{\xi\in C_{+}:{\xi^{1}}^{2}+\ldots+{\xi^{d}}^{2}=1\}.$ (B.300) With this choice $d\mu(\xi)$ coincides with the rotation invariant measure normalized usual: $\omega_{d}=\int_{\gamma_{0}}d\mu(\xi)=\frac{2\pi^{\frac{d}{2}}}{\Gamma\left(\frac{d}{2}\right)}.$ (B.301) $\gamma$.
+The homogeneity degree (1d)(1-d) of the integrand implies that it is a closed differential form and therefore the result depends only on the invariant scalar product $\zeta$.
+A straightforward calculation [13, 14] shows that
+$W^{(dS)}_{\kappa}(z_{1},z_{2})$ $=\frac{\Gamma\Big(\frac{d-1}{2}+i\kappa\Big)\Gamma\Big(\frac{d-1}{2}-i\kappa\Big)}{2(2\pi)^{d/2}}(\zeta-1)^{-\frac{d-2}{4}}(\zeta+1)^{-\frac{d-2}{4}}P_{-\frac{1}{2}+i\kappa}^{-\frac{d-2}{2}}(\zeta).$ (B.302)
+As opposed to the AdS case, the two-point function (B.302) is invariant under the involution $\kappa\to-\kappa$.
+It explicitly provides the analytic continuation of (B.299) to the cut plane $\Delta$, which is the projection of all pairs of complex events of $dS_{d}^{(c)}$, except for pairs of real events which are timelike separated, whose image forms the causal cut.
+In particular, $\Delta$ contains the images of all pairs of non-coincident de Sitter Euclidean points.
+Eq. (8.88) — which is nothing but an identity among Legendre functions — may now be rewritten in a more suggestive way in the present context:
+$W^{(d)}_{-\nu}(z_{1}\cdot z_{2})-W^{(d)}_{\nu}(z_{1}\cdot z_{2})=\sin(\pi\nu)\,W^{(dS)}_{i\nu}(z_{1}\cdot z_{2}),$ (B.303)
+valid for pairs of events $z_{1},z_{2}\in AdS_{d}^{(c)}$ such that their AdS scalar product $z_{1}\cdot z_{2}$ belongs to the cut plane $\Delta_{2}$, here regarded as the fundamental sheet $\widetilde{\Delta}_{1}$.
+B.1 Constructing the split representation of the Euclidean propagator
+We now present the construction of the ”split representation” of the AdS Euclidean propagator (8.91).
+This is done in two steps.
+In the first step we expand the propagator (8.91), which is essentially a Legendre function of the second kind, as a superposition of Legendre functions of the first kind.
+This idea was first introduced as a useful technical tool in [43]. In the two-dimensional case such an expansion has been known for a very long time [44, Eq. (7.213)]:
+$Q_{-\frac{1}{2}+\nu}(z)=\int_{0}^{\infty}\kappa\,\mbox{th\,}(\pi\kappa)P_{-\frac{1}{2}+i\kappa}(z)\,(\kappa^{2}+\nu^{2})^{-1}\ d\kappa,\ \ \ \mathrm{Re}\,\nu>0,\ \ z\in\Delta_{2}\;.$ (B.304)
+The above identity holds, in particular, for real $z>1$.
+However, it is clear that a problem arises on the cut [1,1][-1,1].
+Indeed, the integrand is holomorphic in the cut-plane $\Delta$, whereas the Legendre function on the left-hand side is defined only for $z\in\Delta_{2}\subset\Delta$.
+Since the branch cut of $Q$ extends further, the integral representation becomes ill-defined when $z$ lies on the cut.
+To find the extension of the above formula to general dimension $d$ we may take the Riemann-Liouville transform of both sides of (B.304) [41]; the sought extension immediately follows:
+$e^{-i\pi\frac{d-2}{2}}Q_{-\frac{1}{2}+\nu}^{\frac{d-2}{2}}(z)$ $=\int_{0}^{\infty}\frac{\kappa\,\mbox{sh\,}(\pi\kappa)\Gamma\left(\frac{d-1}{2}-i\kappa\right)\Gamma\left(\frac{d-1}{2}+i\kappa\right)}{\pi(\kappa^{2}+\nu^{2})}{}P^{-\frac{d-2}{2}}_{-\frac{1}{2}+i\kappa}(z)d\kappa.$ (B.305)
+In our context, by using eqs. (8.80) and (B.302), the above identity may be rewritten as follows
+$W^{(d)}_{\nu}(z_{1}\cdot z_{2})=\frac{2}{\pi}\int_{0}^{\infty}\frac{\kappa\,\mbox{sh\,}(\pi\kappa)}{(\kappa^{2}+\nu^{2})}W^{(dS)}_{\kappa}(z_{1}\cdot z_{2})d\kappa,\ \ z_{1},z_{2}\in AdS_{d}^{(c)},\ z_{1}\cdot z_{2}\in\Delta_{2}.$ (B.306)
+However this is not the only possible choice to deploy eq.
+(B.305).
+Given any two complex AdS events $z_{1}$ and $z_{2}$ such that $z_{1}\cdot z_{2}\in\Delta_{2}$, there exist infinitely many pairs of complex de Sitter events $\tilde{z}_{1}$ and $\tilde{z}_{2}$ such that $\tilde{z}_{1}\in{\cal T}_{-}$  , $\tilde{z}_{2}\in{\cal T}_{+}$ and
+$(z_{1}\cdot z_{2})_{AdS}=\langle\tilde{z}_{1},\tilde{z}_{2}\rangle_{dS}.$ (B.307)
+With all the above specifications the following identity holds:
+$W^{(d)}_{\nu}(z_{1}\cdot z_{2})=\int_{0}^{\infty}d\kappa\frac{e^{\pi\kappa}\Gamma\Big(\frac{d-1}{2}+i\kappa\Big)\Gamma\Big(\frac{d-1}{2}-i\kappa\Big)}{(2\pi)^{d}(\kappa^{2}+\nu^{2})\Gamma(i\kappa)\Gamma(-i\kappa)}\int_{\gamma}\langle\xi\,,\tilde{z}_{1}\rangle^{-\frac{d-1}{2}-i\kappa}\langle\xi\,,\tilde{z}_{2}\rangle^{-\frac{d-1}{2}+i\kappa}\,d\mu(\xi).$ (B.308)
+Suppose in particular that the AdS events are Euclidean: $z_{1},z_{2}\in EAdS_{d}$.
+We may introduce two possible global maps from the Euclidean AdS manifold to the each of the inner boundaries of the de Sitter tubes as follows:
+$EAdS_{d}\ni\left\{\begin{array}[]{l}y_{1}^{\scriptscriptstyle{E}}=(iy_{1}^{0},\vec{y}_{1},y_{1}^{d})\to\tilde{y_{1}}=(-iy_{1}^{d},-i\vec{y}_{1},-iy_{1}^{0})\in{\cal H}_{-}\ \ (y^{d}_{1}>0),\cr\cr y_{2}^{\scriptscriptstyle{E}}=(iy_{2}^{0},\vec{y}_{2},y_{2}^{d})\to\tilde{y_{2}}=(iy_{2}^{d},i\vec{y}_{2},iy_{2}^{0})\in{\cal H}_{+}\ \ \ \ \ \ \ \ \ (y^{d}_{2}>0)\end{array}\right$.
+(B.309)
+It follows that
+$z_{1}^{\scriptscriptstyle{E}}\cdot z_{1}^{\scriptscriptstyle{E}}=y_{1}^{d}y_{2}^{d}-\vec{y}_{1}\cdot\vec{y}_{2}-y_{1}^{0}y_{2}^{0}=\langle\tilde{z_{1}},\tilde{z_{2}}\rangle=\mbox{ch\,}v.$ (B.310)
+Inserting (B.310) into Eq. (B.306) we finally arrive at the ”split representation” [34, 35, 36, 37] of the AdS Schwinger propagator:
+$S^{(d)}_{\lambda}(y_{1}^{\scriptscriptstyle{E}},y_{2}^{\scriptscriptstyle{E}})=W^{(d)}_{\nu(\lambda)}(y_{1}^{\scriptscriptstyle{E}}\cdot y_{2}^{\scriptscriptstyle{E}})=\frac{e^{-i\pi\frac{d-2}{2}}}{(2\pi)^{\frac{d}{2}}}(\mbox{sh\,}(v))^{-\frac{d-2}{2}}Q^{\frac{d-2}{2}}_{-\frac{1}{2}+\nu}(\mbox{ch\,}v)=$ (B.311)
+$=\int_{0}^{\infty}d\kappa\frac{e^{\pi\kappa}\Gamma\Big(\frac{d-1}{2}+i\kappa\Big)\Gamma\Big(\frac{d-1}{2}-i\kappa\Big)}{(2\pi)^{d}(\kappa^{2}+\nu^{2})\Gamma(i\kappa)\Gamma(-i\kappa)}\int_{\gamma}\langle\xi\,,\tilde{y}_{1}\rangle^{-\frac{d-1}{2}-i\kappa}\langle\xi\,,\tilde{y}_{2}\rangle^{-\frac{d-1}{2}+i\kappa}\,d\mu(\xi)$ (B.312)
+which has proven to be useful in computing diagrams on the Euclidean anti de Sitter space.
+However, it appears from the above construction that a straightforward Wick rotation to the real anti de Sitter manifold is out of the question.
+Acknowledgments
+I would like to express my gratitude to Sergio Cacciatori for his numerous suggestions.
+I also thank Charlotte Sleight and Massimo Taronna for several insightful discussions.
+I thank an anonymous referee for his observations, which have helped to improve the manuscript.
+My deep appreciation goes to the Institut des Hautes Études Scientifiques (Bures-sur-Yvette) and the Physics Department of Fudan University for their warm hospitality and generous support during the preparation of this paper.